### v4

# Uncertainty Quantification for LLMs through Minimum Bayes Risk: Bridging Confidence and Consistency

Abstract
Uncertainty quantification (UQ) methods for Large Language Models (LLMs) encompass a variety of approaches, with two major types being particularly prominent: information-based, which focus on model confidence expressed as token probabilities, and consistency-based, which assess the semantic relationship between multiple outputs generated using repeated sampling. Several recent methods have combined these two approaches to boost UQ performance. However, they sometimes fail to outperform much simpler baseline methods. Our work discusses the fundamental approach to constructing uncertainty measures that directly links uncertainty with the minimum Bayes risks achieved by LLM decoding. Building on these findings, we propose a novel approach to integrating model confidence with output consistency, resulting in a family of efficient and robust UQ methods. We evaluate our approach across various tasks such as question answering, abstractive summarization, and machine translation, demonstrating sizable improvements over state-of-the-art UQ approaches.
1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling advances in information retrieval Zhu et al. (2023), question answering Kwiatkowski et al. (2019), machine translation Kocmi and Federmann (2023), and a broad range of other NLP applications. As these models become an integral part of our everyday life, ensuring the reliability of their outputs is crucial, especially in high-stakes scenarios where errors can have serious consequences. One way to address this challenge is through uncertainty quantification (UQ), which estimates the reliability of model outputs and enables abstention from generations that are likely to be incorrect.
UQ for LLMs is a rapidly advancing research area, with new UQ methods emerging each year. Most novel techniques are based on two fundamental approaches: (a) information-theoretic analysis or (b) assessment of output consistency.
Information-theoretic methods quantify the confidence of a model by analyzing the probability distributions it induces for predictions (Malinin and Gales, 2021; Fomicheva et al., 2020). A key limitation of these methods is that they cannot account for semantic variability across multiple possible responses for the same input. Specifically, the model may generate answers with the same meaning but with very different assigned probabilities; see Figure 1. LLMs are trained to predict the next token in a sequence based on patterns observed in vast amounts of data, resulting in varying probabilities for semantically equivalent output sequences.
In contrast, consistency-based methods directly analyze the semantic relationship between the sampled outputs (Fomicheva et al., 2020; Lin et al., 2024), capturing uncertainty as objective variability of meaning among the sampled outputs.
Information-theoretic and consistency-based methods have complementary strengths. For this reason, recent state-of-the-art methods aimed to unify these approaches (Kuhn et al., 2023; Duan et al., 2024). We follow this direction and propose a way of quantifying risk as a combination of these basic measures. This results in a family of efficient and robust UQ techniques. Our approach combines the strengths of both information-based and consistency-based methods, providing a more comprehensive and accurate assessment of uncertainty.
Recently, it has been argued Wang and Holmes (2025); Daheim et al. (2025) that UQ in NLP tasks can be viewed through the lens of the Minimum Bayes Risk (MBR) framework. Following this, we formulate our methods as particular forms of risk functions under the MBR approach.
Our main contributions can be summarized as follows:
We propose a new way of quantifying risk that combines information-theoretic and consistency-based measures, and derive a family of Confidence and Consistency-based Approaches (CoCoA) to uncertainty quantification in LLMs.
We show that these measures can be approximated by a learned function trained on unlabeled held-out set with a negligible loss of performance, eliminating the need for costly repeated sampling from the LLM.
We evaluate our approaches across a variety of NLP tasks, including question answering, summarization, and machine translation. Our experiments demonstrate sizable improvements in the reliability and robustness of UQ compared to state-of-the-art methods.
2 Background
2.1 Language Model Decoding
LLMs define a probabilistic output distribution for the given input sequence . The standard ways to obtain a particular output given include various variants of sampling and greedy decoding, where
An alternative approach to decoding is Minimum Bayes Risk (MBR) decoding Kumar and Byrne (2004):
where is the set of candidate sequences and is a risk function:
with being a pairwise loss function. The standard choice in the MBR literature is to take , where the so-called utility function represents some notion of similarity between generated sequences and . Below we discuss how various decoding strategies and corresponding risk functions lead to well-grounded definitions of uncertainty.
2.2 From Risk to Uncertainty
Generally, an uncertainty function is a mapping that quantifies the level of uncertainty associated with the output of a model , conditioned on the input sequence , which we denote as .
Bayesian uncertainty measures are well known to be strongly connected with minimum risks of various kinds Xu and Raginsky (2022); Kotelevskii et al. (2025). Recently, this connection was revisited in the context of natural language generation Wang and Holmes (2025); Daheim et al. (2025), leading to the development of new MBR-based UQ methods. More precisely, one can consider the smallest achievable risk within or, equivalently, to maximum expected utility:
where is given by (2). Interestingly, MBR-based uncertainty (4) is applicable even to consisting of a single sequence .
Below, we discuss how various existing uncertainty measures can be seen as particular approximations of minimum Bayes risks for various choices of utility functions. We also derive new uncertainty measures based on the MBR framework.
Single Sequence Information-Based Methods. Information-based methods estimate the uncertainty of the generated sequence by aggregating the uncertainty scores of individual tokens. Within the framework of MBR decoding, consider . In this case, Bayes risk corresponds to the expected zero-one loss of decoding:
This leads to one of the simplest information-based uncertainty measures, Maximum Sequence Probability (MSP):
Although widely used as an uncertainty measure across various applications, including natural language generation, its connection to statistical risk has only recently been explored in the UQ literature Kotelevskii et al. (2022, 2025); Aichberger et al. (2024).
Several other measures fall into this category, including Perplexity and Mean Token Entropy (Fomicheva et al., 2020); see Appendix D.1 for details. While using only a single sample makes them computationally efficient, these techniques face three major challenges:
1.
LLMs provide us only the probability of a specific answer, even though the same meaning can often be conveyed in multiple ways. To obtain a proper probability for the meaning of an answer, we need to marginalize over its various possible rephrasings. However, this is not feasible if we generate only a single sample.
2.
Reliability of such methods depends heavily on the underlying LLM being well-calibrated, which is a quality that is hard to define and harder to ensure.
3.
These methods are point estimates that do not provide information about the shape of the output distribution.
Semantic Consistency-Based Methods. The aforementioned issues lead to the development of consistency-based methods based on repetitive sampling from the LLM. Consider that we have sampled a set of outputs , where . Consistency-based UQ methods rely on the diversity of answers sampled from the LLM. The idea is that if the model outputs similar answers for the same prompt over and over again, it is confident in its predictions; otherwise, it is uncertain. These techniques do not require a probability distribution estimated by LLM and can be applied in the black-box setting, when only the generated tokens are available. This case is quite common when LLMs are deployed as a service and are accessible through a limited API.
Formally, consistency-based method start from defining some similarity function between arbitrary generations and . The value indicates the complete equivalence between and , and indicates that there is no similarity. Similarity could be computed in various ways. For instance, the Lexical Similarity method Fomicheva et al. (2020) relies on surface-form similarity, measuring the degree of word- or phrase-level overlap between generated texts. More advanced techniques propose various methods for taking into account the semantic similarity of the generated answers by hard or soft clustering (Lin et al., 2024).
Finally, given samples from the model, standard consistency-based methods compute a similarity matrix , where . Then, various statistics of are computed in order to estimate uncertainty (Lin et al., 2024), see details in Appendix D.2. Alternatively, one can focus on a particular output and compute uncertainty measures based on its similarities with other samples ; see Lin et al. (2024) for more details.
Interestingly, it was recently shown Wang and Holmes (2025); Daheim et al. (2025) that the latter approach has direct relation to MBR decoding. Building on the MBR framework, one can define the utility function as and derive the following consistency-based uncertainty score:
Then the Monte Carlo approximation of the minimum Bayes risk is given by:
where . In our ablation study, we show that such an uncertainty measure reliably outperforms consistency-based measures that aggregate the pairwise similarities of all the samples (see Appendix C.2).
A key strength of consistency-based techniques is that, by generating multiple samples and analyzing their semantic similarity, they can estimate empirical probabilities over meanings rather than individual answers. The main drawback is that they discard the useful information that comes from the probability distribution represented by the LLM, including estimates of probabilities of specific answers.
3 CoCoA: Bridging Confidence and Consistency for Better Uncertainty Quantification
Both information-based and (semantic) consistency-based methods provide grounded and useful uncertainty quantification measures. There exist approaches that bridge the gap between information-based and consistency-based methods that show great promise but lack the fundamental base; see discussion in Section 4. In what follows, we present a family of Confidence and Consistency-based Approaches (CoCoA) for UQ, offering a new way to merge information- and consistency-based measures for uncertainty quantification in LLMs via the unified MBR-based framework.
If we consider only the consistency measure, like the one in (8), we omit the information that is contained in the model confidence. Thus, we want the resulting uncertainty measure to explicitly consider both semantic consistency and model-based uncertainty. Let us consider the risk of the form:
where
is any utility function used in the standard MBR decoding (usually representing semantic similarity);
is the model-based uncertainty measure for the output sequence . For example, we can use (similarly to (6)) or .
Then, the corresponding Bayes risk is and for an output sequence , the resulting uncertainty measure becomes:
Finally, given a set of samples we obtain an empirical uncertainty estimate:
where .
The resulting uncertainty measure integrates both global (semantic) and local (model-specific) uncertainty signals. It ensures that uncertainty is amplified for sequences that are both intrinsically uncertain (high ) and semantically inconsistent with the other samples (high ), while keeping it low for the opposite scenario; see Figure 2.
CoCoA Light. Computing the consistency-based uncertainty measure within requires multiple samples, leading to significant computational overhead. To address this, we propose to approximate the behavior of with a learned function. Our method closely matches the original sampling-based measure in quality while requiring only greedy output generation during inference, eliminating the need for additional sampling. Notably, this approximation can be learned without access to ground-truth labels, relying solely on the input data and the associated uncertainty values.
Thus, the CoCoA uncertainty measure that incorporates an approximation of the consistency uncertainty can be defined as:
We name this approach CoCoA Light. Below, we describe how this approximation can be obtained.
Assume access to a held-out set of input sequences . We emphasize that this held-out set is not labeled. For each sequence, we extract the embeddings of the corresponding greedy model generations . The corresponding targets are the consistency uncertainty scores , computed via Monte Carlo estimation of the minimum Bayes risk; see equation (8).
A lightweight auxiliary model, , is trained in a supervised fashion to map model embeddings to uncertainty scores. During inference, the model generates a greedy output for a test input , from which the embeddings are obtained and passed to the auxiliary model to predict the uncertainty score:
We note that a similar approach (Kossen et al., 2024) has been previously proposed to approximate semantic entropy using hidden states from the model.
4 Related Work
Information-based methods. These methods are one of the most commonly used. They quantify uncertainty by analyzing the probability distributions of tokens within a given output. The simplest of these methods, Maximum Sequence Probability (MSP), measures the probability of the most likely output sequence given a specific input. Perplexity is another common uncertainty measure: it gauges how well the model predicts each token in a sequence and is formally defined as the exponential of the mean negative log-likelihood of those tokens (Fomicheva et al., 2020). The Mean Token Entropy method evaluates token-level predictions across the entire sequence by computing the average entropy of the token probability distributions at each position (Fomicheva et al., 2020). While these methods provide useful uncertainty estimates, they do not address the broader uncertainty inherent in generative tasks, as they rely on a single sequence and fail to capture the diversity of possible outputs that the model could generate for the same input.
Consistency-Based Methods. These methods estimate uncertainty by generating multiple sequences from the model’s output distribution and analyzing the variability among the sampled outputs. These methods are particularly relevant to NLP tasks, as they account for scenarios in which multiple plausible outputs may exist for a given input. Among the simplest consistency-based methods are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian Lin et al. (2024), which quantify uncertainty by how many distinct “meanings” the model produces. While effective at capturing global variation, they do not yield uncertainty scores for individual responses. To overcome this, the diagonal of the Degree Matrix was proposed to assess how similar each output is to the rest, enabling per-response uncertainty estimation Lin et al. (2024). A related approach, Lexical Similarity, operates by calculating the average similarity of words or phrases across every pair of responses in the sample Fomicheva et al. (2020).
Information-Based Methods with Repeated Sampling. More recent methods attempt to reconcile output consistency with information-based signals—i.e., they not only look at how varied outputs are, but also how probable each output is under the model’s own generation process. For instance, Semantic Entropy (Kuhn et al., 2023) groups outputs into semantically homogeneous clusters—capturing distinct “meanings” that might just differ in phrasing—and calculates entropy across these clusters. SentenceSAR refines this idea by defining “relevance” as the sum of pairwise similarities between a given sentence and others weighted by each sentence’s model-assigned generation probability (Duan et al., 2024). SAR combines the sentence-level relevance of SentenceSAR with token-level probability adjustments, attempting a more fine-grained balance of semantic and token-level uncertainty (Duan et al., 2024). Semantic Density, Qiu and Miikkulainen (2024) evaluates uncertainty by evaluating how densely a model’s generated response is situated within the semantic space of all possible outputs, with lower density indicating higher uncertainty. One limitation of sampling-based approaches is their computational cost: generating multiple outputs from the model can be expensive. Recent work has sought to address this by learning to predict semantic uncertainty directly from hidden states, allowing for the approximation of semantic entropy without repeated sampling Kossen et al. (2024). Finally, BSDetector (Chen and Mueller, 2024) proposes an additive composition of self-reported confidence with observed consistency, however it relies on questionable ability of the model to assess it’s own confidence in text form. Another limitation is that it requires selection of a trade-off coefficient between confidence and consistency parts.
While these probability-weighted measures can provide deeper insights, they also sometimes underperform in practice (Vashurin et al., 2025), especially when the model’s probability estimates are unreliable, or when important nuances of semantic diversity are lost in the weighting process. Proper balancing raw output consistency with generation likelihood remains an open problem.
Minimum Bayes Risk for LLMs. Recent work has proposed incorporating uncertainty into Minimum Bayes Risk decoding by introducing a posterior over model parameters, enabling uncertainty-aware generation Daheim et al. (2025). Complementing this, a Bayesian framework for subjective uncertainty quantification was developed yielding improved calibration in generation tasks (Wang and Holmes, 2025).
5 Experiments
5.1 Experimental Setup
To evaluate the effectiveness of the proposed method, we extended the LM-Polygraph library (Vashurin et al., 2025; Fadeeva et al., 2023) by implementing our approach within its framework. Since the library already includes tools for calculating other uncertainty scores, it provided a convenient and efficient environment for setting up and running experiments. The primary objective of our experiments is to evaluate whether our method offers improved performance in key tasks such as question answering (QA), summarization (SUM), and machine translation (MT), compared to existing baselines.
Datasets. For QA, we selected diverse datasets to capture a variety of challenges: TriviaQA (Joshi et al., 2017), an open-domain factual QA dataset; CoQA (Reddy et al., 2019), a conversational QA benchmark requiring multi-turn contextual understanding; MMLU (Hendrycks et al., 2021), a multi-task dataset spanning 57 topics to test broad knowledge; and GSM8k (Cobbe et al., 2021), which focuses on grade-school math problems requiring logical reasoning. For translation, we evaluated our method on WMT14 French-English (Bojar et al., 2014) and WMT19 German-English (Barrault et al., 2019). Finally, for summarization, we used XSUM (Narayan et al., 2018), a dataset of complex documents paired with concise abstractive summaries.
For all datasets, we follow (Vashurin et al., 2025) for selecting the subsets, for prompt formatting, and for the number and sourcing of few-shot examples.
Models. We evaluated our method using the base versions of three open-weight language models: LLaMA 3.1 8B Touvron et al. (2023), Mistral 7B Jiang et al. (2023), and Falcon 3 7B Team (2024). The open-source nature of these models provides full access to their token probabilities, which are essential for implementing our UQ method. For all models, we consider base versions, without instruction tuning.
Similarity Function. To measure the similarity, we use the RoBERTa-large cross-encoder model, fine-tuned on the Semantic Textual Similarity benchmark dataset (Liu et al., 2019; Reimers and Gurevych, 2019; Cer et al., 2017). This model is widely regarded as one of the most reliable and commonly used approaches for evaluating sentence similarity. The cross-encoder processes two sequences jointly and directly outputs a similarity score ranging from 0 to 1, providing a nuanced measure.
Appendix C.1 contains comparative experiments with cross-encoder and other choices of the similarity function, substantiating this choice.
Baselines. We compare the performance of the proposed method against a diverse set of baselines and state-of-the-art UQ scores, including confidence-based, consistency-based, and hybrid approaches. Information-based methods include Maximum Sequence Probability (MSP), Perplexity (PPL), Mean Token Entropy (MTE), Monte Carlo Sequence Entropy (MCSE), and Monte Carlo Normalized Sequence Entropy (MCNSE). Consistency-based methods include the Degree Matrix (DegMat) and the Sum of Eigenvalues of the Graph Laplacian (EigValLaplacian). Finally, hybrid methods include Semantic Entropy and SAR. All formulations for these baselines can be found in Appendix D. Lastly, we evaluate performance of (Consistency) and (Consitency Light) as an uncertainty measure.
Evaluation measure. As our evaluation measure, we choose the Prediction Rejection Ratio (PRR), which measures the effectiveness of the uncertainty scores for identifying high-quality predictions (Malinin and Gales, 2021). PRR operates by progressively rejecting predictions with uncertainty scores above a threshold and observing how the average quality of the remaining predictions changes (Figure 3). The metric is calculated as the ratio of two areas: the area between the Prediction Rejection (PR) curves for the evaluated uncertainty score and a random baseline, and the area between the oracle (the ideal uncertainty score that perfectly ranks instances by quality) and the random baseline. Formally, PRR is defined as follows:
Higher PRR values indicate better alignment of uncertainty scores with prediction quality, approaching the performance of an oracle. To ensure practical applicability, we compute PRR only up to a rejection threshold of 50%, preventing cases where excessive rejection artificially inflates quality measures.
For some of the datasets we also report AUROC results, see Appendix G.
Quality Measures. PRR requires an appropriate quality measure for each specific task to effectively evaluate the model output. For question-answering tasks, we use Accuracy to directly evaluate whether the generated answers match the ground truth in short-form QA tasks (e.g., MMLU), and we use the AlignScore between the correct answer and generated sequence for assessing the performance for long-form QA tasks (Zha et al., 2023). For summarization tasks, we use AlignScore to measure the alignment between the output summary and the input document. It serves as a quality indicator by evaluating the relevance and the overlap between the generated content and the source text. For translation tasks, we use COMET, as it captures both semantic adequacy and fluency, ensuring that translations are accurate and linguistically appropriate (Rei et al., 2020).
To further improve comprehensiveness of reported results, we considered alternative choices of quality measures, and report corresponding PRR scores in Appendix F.
Generation Setup. We discuss the generation parameters, the decoding strategy, and sample selection procedure in depth in Appendix A. In short, we report evaluation results in two distinct setups: greedy decoding and stochastic sampling with a focus on the most probable sequence among the generated outputs (best-sample). These two setups offer the highest-quality outputs and are the most reasonable generation approaches in practice.
CoCoA Light Details. The multilayer perceptron is used as an auxiliary model. The features used are embeddings from the middle layers of the base LLM, which are the most informative according to recent work Chen et al. (2024). In particular, for Llama 3.1 8B and Mistral 7B models, embeddings are taken from the 16th layer, and for Falcon 3 7B from the 14th layer.
| Metric | Llama8b-Base | Mistral7b-Base | Falcon7b-Base | ||||||
| QA | NMT | SUM | QA | NMT | SUM | QA | NMT | SUM | |
| MCSE | 0.310 | 0.323 | 0.033 | 0.389 | 0.304 | 0.007 | 0.414 | 0.317 | 0.159 |
| MCNSE | 0.309 | 0.393 | 0.022 | 0.384 | 0.410 | 0.009 | 0.405 | 0.422 | 0.108 |
| Semantic Entropy | 0.356 | 0.343 | 0.033 | 0.423 | 0.327 | 0.008 | 0.439 | 0.348 | 0.164 |
| DegMat | 0.406 | 0.302 | 0.081 | 0.423 | 0.305 | 0.137 | 0.483 | 0.353 | 0.201 |
| EigValLaplacian | 0.375 | 0.238 | 0.079 | 0.391 | 0.267 | 0.132 | 0.459 | 0.312 | 0.201 |
| SAR | 0.414 | 0.455 | 0.077 | 0.462 | 0.435 | 0.094 | 0.481 | 0.458 | 0.144 |
| Consistency | 0.437 | 0.421 | 0.024 | 0.471 | 0.392 | 0.051 | 0.494 | 0.416 | 0.226 |
| Consistency Light | 0.390 | 0.458 | -0.022 | 0.444 | 0.387 | -0.006 | 0.427 | 0.476 | 0.232 |
| MSP | 0.409 | 0.399 | 0.328 | 0.475 | 0.383 | 0.287 | 0.475 | 0.356 | 0.201 |
| 0.451 | 0.519 | 0.378 | 0.509 | 0.497 | 0.330 | 0.511 | 0.505 | 0.257 | |
| 0.449 | 0.502 | 0.358 | 0.503 | 0.480 | 0.309 | 0.492 | 0.514 | 0.242 | |
| PPL | 0.381 | 0.386 | 0.369 | 0.424 | 0.427 | 0.204 | 0.456 | 0.450 | 0.155 |
| 0.454 | 0.481 | 0.387 | 0.494 | 0.472 | 0.286 | 0.523 | 0.508 | 0.229 | |
| 0.445 | 0.487 | 0.382 | 0.479 | 0.480 | 0.260 | 0.512 | 0.528 | 0.234 | |
| MTE | 0.353 | 0.382 | 0.357 | 0.417 | 0.438 | 0.182 | 0.456 | 0.473 | 0.152 |
| 0.447 | 0.478 | 0.380 | 0.492 | 0.469 | 0.288 | 0.527 | 0.508 | 0.228 | |
| 0.428 | 0.494 | 0.372 | 0.475 | 0.482 | 0.254 | 0.507 | 0.533 | 0.234 | |
5.2 Results
Main results. Table 1 shows the PRR scores under the greedy generation setup (See Appendix B for Best Sample and MBR Sample). We report aggregated PRR for each type of task – question answering, neural machine translation (NMT), and summarization (SUM) – by averaging the results across all relevant datasets (e.g., TriviaQA, MMLU, CoQA, GSM8k for QA). This aggregated score provides a concise measure of the performance for each model for each task. Detailed results for each dataset separately can be found in Appendix E.
We can see that our CoCoA methods are the best across all tasks and models. They outperform existing consistency-based and hybrid state-of-the-art approaches, like Semantic Entropy and SAR. In addition, the proposed CoCoA approach consistently surpasses the baseline UQ measures: for example, outperforms standard Perplexity, illustrating the advantage of combining token-level confidence with semantic consistency. This pattern holds for other information-based metrics as well, demonstrating that using the consistency between multiple sampled outputs reliably enhances uncertainty quantification.
Ablation study. As a part of our ablation study, we investigate the impact of different similarity measures (see Appendix C.1). We find that for some tasks, the similarity score computed by the Cross-encoder does not yield optimal performance. For example, for question-answering tasks on CoQA and TriviaQA, NLI-derived similarity performs better than the Cross-encoder similarity and outperforms the pure consistency-based uncertainty discussed above.
The next section of our ablation study focuses on alternative forms of combining model confidence and consistency ; see Appendix C.2. First, we consider an additive form of combining them:
The results show that this additive formulation does not perform as well as the multiplicative one. The additive form tends to underemphasize the interaction between the two components, which is critical for capturing the nuanced relationships between confidence and consistency.
We also consider an alternative formulation of the consistency term , as the average of the full pairwise dissimilarity. In this formulation, represents the average inconsistency across all samples rather than focusing solely on the dissimilarity of the evaluated sequence with the other samples. Our experiments demonstrate that this formulation is not very strong. By distributing the consistency computation across all samples, it loses focus on the specific sequence being evaluated.
Lastly, in Appendix C.2, we also consider alternative formulations of the information-based metric that do not rely on logarithmic transformations. While we primarily use logarithms due to their numerical stability, we explore an alternative approach by converting these values back to probabilities and analyzing their impact on uncertainty estimation. Our findings indicate that both formulations exhibit consistent performance and yield similar results. This suggests that while logarithmic transformations enhance numerical stability, the choice between log-based and probability-based formulations does not affect much the overall performance.
6 Limitations
While our proposed CoCoA approach demonstrates robust empirical performance, several important considerations remain.
Task and Domain Dependency. CoCoA relies on an information-based confidence score and a semantic similarity function, whose effectiveness can vary across models, tasks, and domains. In open-ended tasks like creative generation, producing diverse outputs is expected; on the other hand, tasks requiring precise reasoning can be sensitive to subtle errors that generic similarity metrics may miss. Adapting these components to specific domains remains an important direction for future work.
Limited Sample Size. CoCoA estimates consistency by sampling multiple outputs. Generating many samples is computationally expensive, so our experiments—like most sampling-based methods—use relatively small sets. While even a few samples can yield meaningful estimates, they may miss the full diversity of model outputs for complex prompts.
Quality Metric. Finally, the CoCoA’s performance assessment depends on quality metrics (e.g., COMET for machine translation, and Accuracy for QA) that may not capture every nuance of textual outputs. Further refining or extending quality metrics to account for deeper reasoning, factual faithfulness, and stylistic appropriateness would better align uncertainty scores with real-world perceptions of model correctness.
7 Conclusion
We presented CoCoA, a unified approach that integrates Confidence and Consistency for uncertainty quantification in LLMs. By combining confidence scores with semantic similarity between multiple sampled outputs, CoCoA offers a more holistic view of uncertainty than either approach alone. In extensive evaluations on question answering, summarization, and translation, our approach outperformed existing baselines and state-of-the-art UQ methods. Moreover, CoCoA’s flexible design allows easy adaptation to a variety of tasks and settings.
Moving forward, several directions are open for further exploration. These include incorporating more adaptive sampling strategies that efficiently capture the model output space, refining semantic similarity functions for domain-specific tasks, and improving calibration techniques to strengthen the confidence metrics of the model.
Appendix A Decoding Strategy and Sample Selection
Modern LLMs are capable of producing output using a wide range of decoding strategies, and it is not readily apparent which one to use as a foundation for UQ experiments. On top of that, when sampling multiple outputs stochastically, one has to decide which sample to select for comparison with the target sequence and UQ purposes.
To facilitate the choice of decoding and sample selection strategies for our experiments, we conducted an evaluation of model performance with different approaches to both. Table 2 shows average values of corresponding quality metrics for all combinations of models and datasets. We considered 4 approaches for the selection of output that subsequently is used to calculate the quality of generation:
Greedy decoding produces single output by selecting top-1 candidate token at each generation step, thus no further selection of sample is needed.
Random sample corresponds to the case where random output is selected among the number of samples produced by repeatedly prompting the model with the same question. In practice we use first generated sample, highlighting model performance when stochastic decoding is done only once.
Most Probable sample selects the output with highest model-assigned probability among several sampled outputs.
MBR sample selects the output with highest average consistency with respect to other sampled outputs.
We note that selecting a random sample from the model outputs incurs a significant drop in the quality of results on several datasets, most prominently on GSM8k. Based on these observations, we evaluate the efficacy of UQ in three setups: greedy decoding, stochastic sampling with a focus on the most probable sample and MBR decoding.
Regardless of the way the main model response was obtained, responses that were used to quantify consistency in (8) were generated via repeated stochastic sampling. In all the experiments, where stochastic sampling is involved, it was performed with temperature , top-k equal to 50, and top-p equal to .
| Sample | Metric | Greedy | Random Sample | Most Probable Sample | MBR Sample |
| Falcon7b-Base | |||||
| CoQA | Align Score | 0.793 | 0.706 | 0.783 | 0.793 |
| GSM8k | Accuracy | 0.776 | 0.313 | 0.205 | 0.419 |
| MMLU | Accuracy | 0.715 | 0.638 | 0.715 | 0.699 |
| TriviaQA | Align Score | 0.557 | 0.473 | 0.568 | 0.559 |
| WMT14FrEn | Comet | 0.867 | 0.833 | 0.857 | 0.855 |
| WMT19DeEn | Comet | 0.846 | 0.807 | 0.826 | 0.828 |
| XSUM | Align Score | 0.842 | 0.734 | 0.782 | 0.801 |
| Llama8b-Base | |||||
| CoQA | Align Score | 0.756 | 0.671 | 0.743 | 0.766 |
| GSM8k | Accuracy | 0.548 | 0.234 | 0.261 | 0.346 |
| MMLU | Accuracy | 0.570 | 0.368 | 0.577 | 0.469 |
| TriviaQA | Align Score | 0.686 | 0.625 | 0.687 | 0.699 |
| WMT14FrEn | Comet | 0.863 | 0.819 | 0.852 | 0.844 |
| WMT19DeEn | Comet | 0.870 | 0.816 | 0.854 | 0.845 |
| XSUM | Align Score | 0.848 | 0.608 | 0.825 | 0.703 |
| Mistral7b-Base | |||||
| CoQA | Align Score | 0.793 | 0.695 | 0.777 | 0.790 |
| GSM8k | Accuracy | 0.382 | 0.169 | 0.190 | 0.259 |
| MMLU | Accuracy | 0.632 | 0.552 | 0.632 | 0.612 |
| TriviaQA | Align Score | 0.743 | 0.655 | 0.750 | 0.739 |
| WMT14FrEn | Comet | 0.863 | 0.812 | 0.830 | 0.841 |
| WMT19DeEn | Comet | 0.864 | 0.805 | 0.836 | 0.837 |
| XSUM | Align Score | 0.803 | 0.578 | 0.775 | 0.665 |
Appendix B Results Summary for Most Probable Sample and MBR Sample strategies
Tables 3 and 4 report the PRR scores under the Most Probable Sample and MBR Sample generation setups, as discussed in Section 5.2. We observe that CoCoA-family methods, and their Supervised versions, consistently improve base uncertainty estimators.
| Metric | Llama8b-Base | Mistral7b-Base | Falcon7b-Base | ||||||
| QA | NMT | SUM | QA | NMT | SUM | QA | NMT | SUM | |
| MCSE | 0.356 | 0.380 | 0.192 | 0.453 | 0.406 | 0.162 | 0.460 | 0.409 | 0.128 |
| MCNSE | 0.380 | 0.429 | 0.186 | 0.466 | 0.489 | 0.196 | 0.530 | 0.424 | 0.153 |
| Semantic Entropy | 0.396 | 0.411 | 0.194 | 0.482 | 0.438 | 0.164 | 0.479 | 0.440 | 0.134 |
| DegMat | 0.422 | 0.342 | 0.191 | 0.465 | 0.425 | 0.205 | 0.543 | 0.386 | 0.177 |
| EigValLaplacian | 0.388 | 0.274 | 0.190 | 0.426 | 0.366 | 0.197 | 0.498 | 0.336 | 0.174 |
| SAR | 0.478 | 0.506 | 0.159 | 0.542 | 0.576 | 0.175 | 0.590 | 0.488 | 0.193 |
| Consistency | 0.535 | 0.536 | 0.030 | 0.572 | 0.689 | 0.071 | 0.626 | 0.571 | 0.282 |
| MSP | 0.395 | 0.376 | 0.464 | 0.444 | 0.252 | 0.330 | 0.343 | 0.381 | 0.099 |
| 0.484 | 0.607 | 0.484 | 0.526 | 0.721 | 0.366 | 0.529 | 0.631 | 0.210 | |
| PPL | 0.532 | 0.563 | 0.458 | 0.587 | 0.686 | 0.365 | 0.627 | 0.589 | 0.275 |
| 0.571 | 0.617 | 0.450 | 0.613 | 0.745 | 0.372 | 0.647 | 0.648 | 0.310 | |
| MTE | 0.476 | 0.469 | 0.449 | 0.559 | 0.637 | 0.350 | 0.602 | 0.492 | 0.186 |
| 0.547 | 0.579 | 0.451 | 0.600 | 0.720 | 0.373 | 0.641 | 0.614 | 0.289 | |
| Metric | Llama8b-Base | Mistral7b-Base | Falcon7b-Base | ||||||
| QA | NMT | SUM | QA | NMT | SUM | QA | NMT | SUM | |
| MCSE | 0.402 | 0.297 | 0.158 | 0.474 | 0.330 | 0.237 | 0.437 | 0.346 | 0.076 |
| MCNSE | 0.392 | 0.389 | 0.124 | 0.441 | 0.435 | 0.192 | 0.446 | 0.440 | 0.078 |
| Semantic Entropy | 0.443 | 0.316 | 0.159 | 0.507 | 0.352 | 0.236 | 0.458 | 0.373 | 0.077 |
| DegMat | 0.478 | 0.290 | 0.047 | 0.471 | 0.322 | 0.124 | 0.499 | 0.371 | -0.018 |
| EigValLaplacian | 0.434 | 0.224 | 0.037 | 0.437 | 0.279 | 0.115 | 0.460 | 0.326 | -0.022 |
| SAR | 0.509 | 0.456 | 0.187 | 0.532 | 0.448 | 0.266 | 0.531 | 0.478 | 0.083 |
| Consistency | 0.504 | 0.365 | 0.235 | 0.518 | 0.367 | 0.284 | 0.538 | 0.410 | 0.099 |
| MSP | 0.415 | 0.404 | 0.266 | 0.447 | 0.433 | 0.312 | 0.405 | 0.376 | 0.170 |
| 0.527 | 0.496 | 0.293 | 0.553 | 0.515 | 0.348 | 0.500 | 0.526 | 0.154 | |
| PPL | 0.408 | 0.421 | 0.164 | 0.432 | 0.452 | 0.273 | 0.445 | 0.438 | 0.100 |
| 0.532 | 0.504 | 0.215 | 0.554 | 0.514 | 0.328 | 0.552 | 0.537 | 0.109 | |
| MTE | 0.467 | 0.461 | 0.148 | 0.515 | 0.496 | 0.219 | 0.490 | 0.521 | 0.069 |
| 0.542 | 0.479 | 0.201 | 0.574 | 0.486 | 0.290 | 0.562 | 0.527 | 0.081 | |
Appendix C Ablation
C.1 Choice of Similarity Function
For sample consistency estimation, one could come up with a variety of similarity functions . We perform a comparison of the effectiveness of CoCoA-family methods using several such functions. We consider the following functions:
Tables 5, 6 and 7 report these results. There exists a considerable variation of relative effectiveness of proposed methods with various similarity function choices, depending on a task at hand. We opt to report all results in other sections with CrossEncoder-based similarity as it by itself provides a significant improvement over baselines, and for consistency and ease of comparison reasons. However, when applying these methods to a particular task, we encourage users to select appropriate underlying similarity function for best results.
| Method | Dataset | ||||||
| XSUM | WMT14FrEn | WMT19DeEn | CoQA | TriviaQA | MMLU | GSM8k | |
| Mistral7b-Base | |||||||
| AlignScore | 0.334 | 0.293 | 0.445 | 0.354 | 0.655 | 0.466 | 0.550 |
| RougeL | 0.289 | 0.358 | 0.546 | 0.369 | 0.649 | 0.466 | 0.536 |
| NLI | 0.308 | 0.313 | 0.477 | 0.403 | 0.677 | 0.470 | 0.568 |
| CrossEncoder | 0.327 | 0.397 | 0.595 | 0.381 | 0.671 | 0.466 | 0.505 |
| AlignScore | 0.307 | 0.308 | 0.489 | 0.373 | 0.666 | 0.466 | 0.536 |
| RougeL | 0.226 | 0.369 | 0.531 | 0.352 | 0.653 | 0.466 | 0.466 |
| NLI | 0.233 | 0.316 | 0.501 | 0.376 | 0.682 | 0.470 | 0.480 |
| CrossEncoder | 0.281 | 0.371 | 0.565 | 0.365 | 0.674 | 0.466 | 0.465 |
| AlignScore | 0.302 | 0.299 | 0.477 | 0.366 | 0.664 | 0.450 | 0.555 |
| RougeL | 0.212 | 0.377 | 0.528 | 0.345 | 0.652 | 0.449 | 0.497 |
| NLI | 0.219 | 0.313 | 0.488 | 0.362 | 0.681 | 0.453 | 0.490 |
| CrossEncoder | 0.282 | 0.368 | 0.560 | 0.351 | 0.673 | 0.448 | 0.486 |
| Llama8b-Base | |||||||
| AlignScore | 0.367 | 0.331 | 0.452 | 0.308 | 0.596 | 0.484 | 0.401 |
| RougeL | 0.336 | 0.393 | 0.545 | 0.321 | 0.563 | 0.474 | 0.375 |
| NLI | 0.344 | 0.352 | 0.467 | 0.364 | 0.606 | 0.478 | 0.419 |
| CrossEncoder | 0.375 | 0.454 | 0.583 | 0.350 | 0.598 | 0.480 | 0.367 |
| AlignScore | 0.422 | 0.346 | 0.450 | 0.337 | 0.596 | 0.453 | 0.446 |
| RougeL | 0.370 | 0.408 | 0.486 | 0.319 | 0.552 | 0.441 | 0.418 |
| NLI | 0.374 | 0.354 | 0.438 | 0.348 | 0.600 | 0.446 | 0.409 |
| CrossEncoder | 0.380 | 0.444 | 0.514 | 0.339 | 0.593 | 0.447 | 0.429 |
| AlignScore | 0.419 | 0.340 | 0.438 | 0.339 | 0.605 | 0.411 | 0.459 |
| RougeL | 0.362 | 0.417 | 0.481 | 0.319 | 0.560 | 0.390 | 0.440 |
| NLI | 0.366 | 0.342 | 0.428 | 0.340 | 0.612 | 0.396 | 0.420 |
| CrossEncoder | 0.374 | 0.441 | 0.511 | 0.337 | 0.601 | 0.394 | 0.444 |
| Falcon7b-Base | |||||||
| AlignScore | 0.278 | 0.306 | 0.475 | 0.361 | 0.677 | 0.528 | 0.470 |
| RougeL | 0.205 | 0.394 | 0.499 | 0.378 | 0.678 | 0.527 | 0.417 |
| NLI | 0.236 | 0.361 | 0.511 | 0.407 | 0.684 | 0.532 | 0.532 |
| CrossEncoder | 0.253 | 0.436 | 0.577 | 0.396 | 0.685 | 0.529 | 0.428 |
| AlignScore | 0.252 | 0.340 | 0.523 | 0.410 | 0.678 | 0.528 | 0.521 |
| RougeL | 0.170 | 0.409 | 0.537 | 0.389 | 0.668 | 0.527 | 0.439 |
| NLI | 0.193 | 0.364 | 0.531 | 0.408 | 0.680 | 0.532 | 0.499 |
| CrossEncoder | 0.226 | 0.437 | 0.579 | 0.405 | 0.677 | 0.529 | 0.474 |
| AlignScore | 0.253 | 0.337 | 0.519 | 0.403 | 0.683 | 0.515 | 0.554 |
| RougeL | 0.170 | 0.426 | 0.540 | 0.382 | 0.673 | 0.514 | 0.472 |
| NLI | 0.190 | 0.364 | 0.525 | 0.398 | 0.687 | 0.521 | 0.514 |
| CrossEncoder | 0.223 | 0.438 | 0.575 | 0.395 | 0.685 | 0.517 | 0.505 |
| Method | Dataset | ||||||
| XSUM | WMT14FrEn | WMT19DeEn | CoQA | TriviaQA | MMLU | GSM8k | |
| Mistral7b-Base | |||||||
| AlignScore | 0.393 | 0.448 | 0.491 | 0.399 | 0.626 | 0.467 | 0.476 |
| RougeL | 0.344 | 0.602 | 0.597 | 0.420 | 0.622 | 0.466 | 0.538 |
| NLI | 0.340 | 0.615 | 0.604 | 0.445 | 0.651 | 0.470 | 0.456 |
| CrossEncoder | 0.366 | 0.712 | 0.730 | 0.430 | 0.644 | 0.466 | 0.562 |
| AlignScore | 0.474 | 0.619 | 0.657 | 0.408 | 0.638 | 0.467 | 0.910 |
| RougeL | 0.362 | 0.710 | 0.717 | 0.391 | 0.627 | 0.466 | 0.950 |
| NLI | 0.370 | 0.677 | 0.684 | 0.414 | 0.657 | 0.470 | 0.941 |
| CrossEncoder | 0.372 | 0.735 | 0.755 | 0.402 | 0.648 | 0.466 | 0.937 |
| AlignScore | 0.492 | 0.547 | 0.590 | 0.383 | 0.633 | 0.449 | 0.914 |
| RougeL | 0.355 | 0.695 | 0.684 | 0.366 | 0.624 | 0.448 | 0.959 |
| NLI | 0.364 | 0.656 | 0.658 | 0.387 | 0.656 | 0.453 | 0.918 |
| CrossEncoder | 0.373 | 0.708 | 0.732 | 0.373 | 0.645 | 0.447 | 0.935 |
| Llama8b-Base | |||||||
| AlignScore | 0.520 | 0.332 | 0.491 | 0.354 | 0.587 | 0.457 | 0.401 |
| RougeL | 0.471 | 0.470 | 0.588 | 0.362 | 0.551 | 0.446 | 0.499 |
| NLI | 0.466 | 0.442 | 0.577 | 0.386 | 0.597 | 0.446 | 0.470 |
| CrossEncoder | 0.484 | 0.529 | 0.685 | 0.384 | 0.587 | 0.452 | 0.513 |
| AlignScore | 0.546 | 0.406 | 0.561 | 0.376 | 0.577 | 0.429 | 0.875 |
| RougeL | 0.452 | 0.518 | 0.639 | 0.352 | 0.532 | 0.417 | 0.931 |
| NLI | 0.458 | 0.466 | 0.597 | 0.365 | 0.583 | 0.418 | 0.912 |
| CrossEncoder | 0.450 | 0.544 | 0.689 | 0.364 | 0.573 | 0.422 | 0.925 |
| AlignScore | 0.561 | 0.325 | 0.497 | 0.365 | 0.589 | 0.380 | 0.821 |
| RougeL | 0.448 | 0.496 | 0.598 | 0.336 | 0.539 | 0.361 | 0.921 |
| NLI | 0.449 | 0.446 | 0.565 | 0.344 | 0.598 | 0.359 | 0.881 |
| CrossEncoder | 0.451 | 0.520 | 0.638 | 0.346 | 0.582 | 0.363 | 0.900 |
| Falcon7b-Base | |||||||
| AlignScore | 0.181 | 0.378 | 0.473 | 0.410 | 0.654 | 0.528 | 0.239 |
| RougeL | 0.122 | 0.531 | 0.581 | 0.420 | 0.655 | 0.528 | 0.426 |
| NLI | 0.120 | 0.496 | 0.607 | 0.437 | 0.658 | 0.533 | 0.458 |
| CrossEncoder | 0.210 | 0.564 | 0.698 | 0.428 | 0.659 | 0.530 | 0.498 |
| AlignScore | 0.384 | 0.454 | 0.586 | 0.440 | 0.648 | 0.528 | 0.994 |
| RougeL | 0.280 | 0.565 | 0.668 | 0.410 | 0.637 | 0.528 | 1.000 |
| NLI | 0.283 | 0.515 | 0.671 | 0.424 | 0.647 | 0.533 | 0.998 |
| CrossEncoder | 0.310 | 0.579 | 0.717 | 0.415 | 0.644 | 0.530 | 1.000 |
| AlignScore | 0.292 | 0.386 | 0.498 | 0.435 | 0.648 | 0.515 | 0.972 |
| RougeL | 0.222 | 0.545 | 0.607 | 0.400 | 0.633 | 0.515 | 0.998 |
| NLI | 0.201 | 0.498 | 0.636 | 0.415 | 0.645 | 0.521 | 0.987 |
| CrossEncoder | 0.289 | 0.551 | 0.678 | 0.402 | 0.646 | 0.517 | 0.998 |
| Method | Dataset | ||||||
| XSUM | WMT14FrEn | WMT19DeEn | CoQA | TriviaQA | MMLU | GSM8k | |
| Mistral7b-Base | |||||||
| AlignScore | 0.287 | 0.326 | 0.460 | 0.312 | 0.575 | 0.515 | 0.747 |
| RougeL | 0.259 | 0.420 | 0.562 | 0.304 | 0.502 | 0.515 | 0.750 |
| NLI | 0.278 | 0.375 | 0.506 | 0.350 | 0.589 | 0.512 | 0.766 |
| CrossEncoder | 0.295 | 0.441 | 0.589 | 0.336 | 0.597 | 0.514 | 0.766 |
| AlignScore | 0.283 | 0.333 | 0.482 | 0.330 | 0.596 | 0.515 | 0.738 |
| RougeL | 0.219 | 0.427 | 0.561 | 0.300 | 0.517 | 0.515 | 0.744 |
| NLI | 0.219 | 0.390 | 0.508 | 0.334 | 0.602 | 0.512 | 0.723 |
| CrossEncoder | 0.258 | 0.451 | 0.577 | 0.322 | 0.612 | 0.514 | 0.768 |
| AlignScore | 0.269 | 0.314 | 0.447 | 0.357 | 0.663 | 0.497 | 0.746 |
| RougeL | 0.181 | 0.446 | 0.571 | 0.327 | 0.622 | 0.497 | 0.786 |
| NLI | 0.182 | 0.370 | 0.482 | 0.348 | 0.686 | 0.499 | 0.704 |
| CrossEncoder | 0.236 | 0.419 | 0.553 | 0.337 | 0.680 | 0.494 | 0.783 |
| Llama8b-Base | |||||||
| AlignScore | 0.221 | 0.309 | 0.462 | 0.261 | 0.467 | 0.622 | 0.717 |
| RougeL | 0.195 | 0.401 | 0.542 | 0.250 | 0.361 | 0.631 | 0.721 |
| NLI | 0.222 | 0.311 | 0.463 | 0.311 | 0.480 | 0.580 | 0.769 |
| CrossEncoder | 0.225 | 0.403 | 0.590 | 0.304 | 0.487 | 0.586 | 0.731 |
| AlignScore | 0.225 | 0.325 | 0.484 | 0.304 | 0.476 | 0.603 | 0.747 |
| RougeL | 0.169 | 0.416 | 0.537 | 0.274 | 0.370 | 0.612 | 0.758 |
| NLI | 0.180 | 0.334 | 0.460 | 0.319 | 0.488 | 0.558 | 0.761 |
| CrossEncoder | 0.198 | 0.428 | 0.580 | 0.320 | 0.496 | 0.562 | 0.751 |
| AlignScore | 0.209 | 0.293 | 0.460 | 0.320 | 0.563 | 0.556 | 0.749 |
| RougeL | 0.148 | 0.436 | 0.559 | 0.314 | 0.497 | 0.587 | 0.785 |
| NLI | 0.156 | 0.298 | 0.445 | 0.340 | 0.592 | 0.460 | 0.740 |
| CrossEncoder | 0.189 | 0.398 | 0.561 | 0.345 | 0.587 | 0.472 | 0.765 |
| Falcon7b-Base | |||||||
| AlignScore | 0.279 | 0.307 | 0.473 | 0.312 | 0.594 | 0.574 | 0.460 |
| RougeL | 0.202 | 0.424 | 0.487 | 0.307 | 0.566 | 0.574 | 0.447 |
| NLI | 0.231 | 0.362 | 0.508 | 0.355 | 0.618 | 0.577 | 0.531 |
| CrossEncoder | 0.248 | 0.461 | 0.590 | 0.336 | 0.617 | 0.575 | 0.472 |
| AlignScore | 0.278 | 0.320 | 0.534 | 0.350 | 0.607 | 0.574 | 0.617 |
| RougeL | 0.190 | 0.428 | 0.556 | 0.333 | 0.571 | 0.574 | 0.632 |
| NLI | 0.216 | 0.376 | 0.549 | 0.368 | 0.623 | 0.578 | 0.661 |
| CrossEncoder | 0.242 | 0.462 | 0.613 | 0.345 | 0.624 | 0.575 | 0.662 |
| AlignScore | 0.277 | 0.302 | 0.522 | 0.382 | 0.657 | 0.556 | 0.598 |
| RougeL | 0.175 | 0.472 | 0.573 | 0.364 | 0.642 | 0.556 | 0.617 |
| NLI | 0.196 | 0.359 | 0.530 | 0.400 | 0.671 | 0.562 | 0.639 |
| CrossEncoder | 0.237 | 0.454 | 0.600 | 0.373 | 0.666 | 0.557 | 0.650 |
C.2 Different Ways of Combining Confidence and Consistency
We justify the particular form of equation (11) by considering alternative ways to combine sample-focused confidence with consistency estimation. Results are presented in Tables 8, 9 and 10. In particular, we investigate the performance of the additive approach (AdditiveCoCoA):
and the same multiplicative combination, replacing sample-focused dissimilarity from (8) with the average of the full pairwise dissimilarity matrix (24):
In addition, we present results where we compute using probability-based scores instead of log-likelihood, see equation (6). We refer to this variant as . While the results are generally similar, the log-likelihood-based formulation offers greater numerical stability. Therefore, we adopt it as the default in the main paper.
It is evident that on average the multiplicative form proposed in equation (10) with both confidence and consistency terms focused on a single sample is the better performing variant.
| Method | Dataset | ||||||
| XSUM | WMT14FrEn | WMT19DeEn | CoQA | TriviaQA | MMLU | GSM8k | |
| Mistral7b-Base | |||||||
| 0.290 | 0.319 | 0.459 | 0.351 | 0.654 | 0.471 | 0.472 | |
| 0.319 | 0.385 | 0.590 | 0.357 | 0.668 | 0.467 | 0.505 | |
| 0.059 | 0.302 | 0.520 | 0.390 | 0.671 | 0.461 | 0.435 | |
| 0.330 | 0.396 | 0.598 | 0.383 | 0.670 | 0.466 | 0.517 | |
| 0.262 | 0.392 | 0.564 | 0.369 | 0.671 | 0.464 | 0.494 | |
| 0.277 | 0.373 | 0.551 | 0.334 | 0.672 | 0.467 | 0.435 | |
| 0.297 | 0.369 | 0.566 | 0.373 | 0.674 | 0.464 | 0.475 | |
| 0.286 | 0.375 | 0.568 | 0.369 | 0.674 | 0.466 | 0.467 | |
| -0.279 | -0.058 | -0.072 | 0.098 | 0.312 | 0.079 | 0.187 | |
| 0.274 | 0.368 | 0.543 | 0.309 | 0.668 | 0.442 | 0.456 | |
| 0.288 | 0.374 | 0.564 | 0.355 | 0.673 | 0.447 | 0.491 | |
| Llama8b-Base | |||||||
| 0.330 | 0.345 | 0.462 | 0.301 | 0.566 | 0.502 | 0.326 | |
| 0.358 | 0.434 | 0.564 | 0.333 | 0.589 | 0.488 | 0.354 | |
| 0.031 | 0.405 | 0.471 | 0.371 | 0.612 | 0.461 | 0.368 | |
| 0.378 | 0.456 | 0.582 | 0.349 | 0.597 | 0.485 | 0.372 | |
| 0.368 | 0.431 | 0.504 | 0.336 | 0.595 | 0.455 | 0.437 | |
| 0.389 | 0.420 | 0.487 | 0.314 | 0.580 | 0.450 | 0.399 | |
| 0.381 | 0.445 | 0.513 | 0.345 | 0.599 | 0.446 | 0.438 | |
| 0.387 | 0.448 | 0.514 | 0.338 | 0.593 | 0.452 | 0.433 | |
| -0.331 | -0.042 | -0.122 | 0.089 | 0.321 | -0.122 | 0.117 | |
| 0.383 | 0.410 | 0.481 | 0.308 | 0.588 | 0.363 | 0.414 | |
| 0.380 | 0.446 | 0.511 | 0.337 | 0.601 | 0.402 | 0.447 | |
| Falcon7b-Base | |||||||
| 0.203 | 0.318 | 0.409 | 0.350 | 0.674 | 0.533 | 0.379 | |
| 0.225 | 0.423 | 0.571 | 0.388 | 0.678 | 0.533 | 0.404 | |
| 0.226 | 0.367 | 0.515 | 0.416 | 0.680 | 0.526 | 0.426 | |
| 0.257 | 0.433 | 0.578 | 0.396 | 0.684 | 0.529 | 0.436 | |
| 0.222 | 0.433 | 0.580 | 0.413 | 0.677 | 0.525 | 0.489 | |
| 0.204 | 0.425 | 0.565 | 0.393 | 0.669 | 0.533 | 0.437 | |
| 0.235 | 0.433 | 0.576 | 0.410 | 0.680 | 0.528 | 0.482 | |
| 0.229 | 0.436 | 0.580 | 0.406 | 0.677 | 0.529 | 0.478 | |
| 0.001 | -0.103 | -0.106 | 0.114 | 0.041 | 0.138 | 0.221 | |
| 0.201 | 0.425 | 0.557 | 0.377 | 0.675 | 0.519 | 0.470 | |
| 0.228 | 0.439 | 0.577 | 0.395 | 0.685 | 0.517 | 0.510 | |
| Method | Dataset | ||||||
| XSUM | WMT14FrEn | WMT19DeEn | CoQA | TriviaQA | MMLU | GSM8k | |
| Mistral7b-Base | |||||||
| 0.333 | 0.239 | 0.310 | 0.406 | 0.631 | 0.472 | 0.311 | |
| 0.354 | 0.543 | 0.565 | 0.412 | 0.643 | 0.468 | 0.428 | |
| 0.076 | 0.684 | 0.721 | 0.428 | 0.643 | 0.464 | 0.846 | |
| 0.366 | 0.712 | 0.730 | 0.430 | 0.644 | 0.466 | 0.562 | |
| 0.368 | 0.737 | 0.751 | 0.406 | 0.644 | 0.465 | 0.939 | |
| 0.383 | 0.714 | 0.723 | 0.379 | 0.649 | 0.468 | 0.933 | |
| 0.369 | 0.738 | 0.756 | 0.401 | 0.649 | 0.467 | 0.935 | |
| 0.372 | 0.735 | 0.755 | 0.402 | 0.648 | 0.466 | 0.937 | |
| 0.368 | 0.723 | 0.702 | 0.332 | 0.643 | 0.452 | 0.942 | |
| 0.380 | 0.661 | 0.653 | 0.331 | 0.643 | 0.442 | 0.929 | |
| 0.373 | 0.708 | 0.732 | 0.373 | 0.645 | 0.447 | 0.935 | |
| Llama8b-Base | |||||||
| 0.466 | 0.349 | 0.425 | 0.333 | 0.555 | 0.473 | 0.285 | |
| 0.476 | 0.462 | 0.619 | 0.363 | 0.574 | 0.464 | 0.379 | |
| 0.035 | 0.491 | 0.617 | 0.398 | 0.598 | 0.433 | 0.795 | |
| 0.484 | 0.529 | 0.685 | 0.384 | 0.587 | 0.452 | 0.513 | |
| 0.454 | 0.536 | 0.673 | 0.358 | 0.575 | 0.425 | 0.923 | |
| 0.459 | 0.525 | 0.649 | 0.343 | 0.556 | 0.430 | 0.914 | |
| 0.438 | 0.547 | 0.689 | 0.364 | 0.574 | 0.419 | 0.923 | |
| 0.450 | 0.544 | 0.689 | 0.364 | 0.573 | 0.422 | 0.925 | |
| 0.457 | 0.496 | 0.579 | 0.304 | 0.561 | 0.361 | 0.901 | |
| 0.455 | 0.464 | 0.577 | 0.313 | 0.563 | 0.341 | 0.878 | |
| 0.451 | 0.520 | 0.638 | 0.346 | 0.582 | 0.363 | 0.900 | |
| Falcon7b-Base | |||||||
| 0.100 | 0.397 | 0.394 | 0.393 | 0.649 | 0.534 | -0.156 | |
| 0.144 | 0.531 | 0.607 | 0.416 | 0.654 | 0.533 | 0.189 | |
| 0.282 | 0.522 | 0.670 | 0.434 | 0.658 | 0.529 | 0.978 | |
| 0.210 | 0.564 | 0.698 | 0.428 | 0.659 | 0.530 | 0.498 | |
| 0.297 | 0.582 | 0.706 | 0.417 | 0.643 | 0.526 | 1.000 | |
| 0.297 | 0.560 | 0.670 | 0.405 | 0.641 | 0.533 | 1.000 | |
| 0.311 | 0.587 | 0.718 | 0.414 | 0.648 | 0.531 | 1.000 | |
| 0.310 | 0.579 | 0.717 | 0.415 | 0.644 | 0.530 | 1.000 | |
| 0.253 | 0.554 | 0.634 | 0.383 | 0.630 | 0.523 | 0.997 | |
| 0.237 | 0.502 | 0.554 | 0.383 | 0.636 | 0.519 | 0.989 | |
| 0.289 | 0.551 | 0.678 | 0.402 | 0.646 | 0.517 | 0.998 | |
| Method | Dataset | ||||||
| XSUM | WMT14FrEn | WMT19DeEn | CoQA | TriviaQA | MMLU | GSM8k | |
| Mistral7b-Base | |||||||
| 0.259 | 0.382 | 0.488 | 0.231 | 0.388 | 0.512 | 0.711 | |
| 0.287 | 0.447 | 0.591 | 0.318 | 0.570 | 0.516 | 0.748 | |
| 0.258 | 0.374 | 0.473 | 0.149 | 0.099 | -0.265 | 0.710 | |
| 0.295 | 0.441 | 0.589 | 0.336 | 0.597 | 0.514 | 0.766 | |
| 0.234 | 0.430 | 0.560 | 0.255 | 0.517 | 0.507 | 0.721 | |
| 0.251 | 0.458 | 0.571 | 0.305 | 0.587 | 0.516 | 0.745 | |
| 0.094 | -0.123 | -0.143 | -0.202 | -0.404 | -0.423 | 0.027 | |
| 0.258 | 0.451 | 0.577 | 0.322 | 0.612 | 0.514 | 0.768 | |
| 0.201 | 0.468 | 0.585 | 0.296 | 0.669 | 0.501 | 0.760 | |
| 0.232 | 0.426 | 0.544 | 0.318 | 0.673 | 0.496 | 0.772 | |
| 0.236 | 0.419 | 0.553 | 0.337 | 0.680 | 0.494 | 0.783 | |
| Llama8b-Base | |||||||
| 0.197 | 0.345 | 0.465 | 0.177 | 0.287 | 0.571 | 0.676 | |
| 0.214 | 0.409 | 0.589 | 0.289 | 0.449 | 0.589 | 0.724 | |
| 0.196 | 0.337 | 0.449 | 0.102 | 0.053 | 0.434 | 0.673 | |
| 0.225 | 0.403 | 0.590 | 0.304 | 0.487 | 0.586 | 0.731 | |
| 0.182 | 0.406 | 0.500 | 0.229 | 0.401 | 0.540 | 0.715 | |
| 0.186 | 0.433 | 0.563 | 0.304 | 0.461 | 0.565 | 0.746 | |
| 0.101 | -0.097 | -0.152 | -0.175 | -0.367 | 0.099 | 0.053 | |
| 0.198 | 0.428 | 0.580 | 0.320 | 0.496 | 0.562 | 0.751 | |
| 0.160 | 0.437 | 0.554 | 0.302 | 0.560 | 0.477 | 0.748 | |
| 0.174 | 0.400 | 0.543 | 0.334 | 0.568 | 0.475 | 0.762 | |
| 0.189 | 0.398 | 0.561 | 0.345 | 0.587 | 0.472 | 0.765 | |
| Falcon7b-Base | |||||||
| 0.188 | 0.356 | 0.401 | 0.228 | 0.488 | 0.578 | 0.368 | |
| 0.237 | 0.468 | 0.586 | 0.325 | 0.593 | 0.576 | 0.453 | |
| 0.184 | 0.339 | 0.378 | 0.144 | 0.350 | -0.345 | 0.364 | |
| 0.248 | 0.461 | 0.590 | 0.336 | 0.617 | 0.575 | 0.472 | |
| 0.206 | 0.404 | 0.565 | 0.271 | 0.537 | 0.572 | 0.625 | |
| 0.231 | 0.464 | 0.606 | 0.337 | 0.598 | 0.577 | 0.650 | |
| 0.014 | -0.167 | -0.265 | -0.187 | -0.237 | -0.416 | -0.206 | |
| 0.242 | 0.462 | 0.613 | 0.345 | 0.624 | 0.575 | 0.662 | |
| 0.191 | 0.488 | 0.612 | 0.353 | 0.644 | 0.564 | 0.617 | |
| 0.219 | 0.452 | 0.591 | 0.368 | 0.654 | 0.556 | 0.629 | |
| 0.237 | 0.454 | 0.600 | 0.373 | 0.666 | 0.557 | 0.650 | |
Appendix D Detailed Description of Uncertainty Quantification Methods
In this section, we provide a detailed description of the uncertainty quantification methods used in this study.
D.1 Information-Based Methods
Information-based methods are commonly used to estimate uncertainty by analyzing the probability distributions of tokens within a given output. These methods examine different levels of model generation, such as the model’s confidence in producing a specific sequence, its ability to predict individual tokens at each generation step, and the variability in the token-level predictions across the sequence.
Maximum Sequence Probability (MSP) is one of the simplest and most direct methods for estimating uncertainty. It measures the probability of the most likely output sequence given a specific input. Thus, uncertainty is quantified by calculating the probability of the sequence with the highest likelihood, under the assumption that the model is most confident in this output. In its simplest form it is given by (6). An equivalent (in terms of ordering of predictions for different inputs) but more numerically stable formulation is logarithmic. It is defined as:
Perplexity (PPL) is another widely used metric for estimating uncertainty in language models [Fomicheva et al., 2020]. It measures the model’s confidence by evaluating the average likelihood of generating the sequence tokens:
Mean Token Entropy takes a broader view of uncertainty by considering the token-level predictions across the entire sequence [Fomicheva et al., 2020]. Instead of evaluating the model’s confidence in a single output or individual token predictions, Mean Token Entropy calculates the average entropy of the token probability distributions for each token in the sequence:
where is an entropy of the token distribution .
The TokenSAR method, introduced in Duan et al. [2024], generalizes length-normalized log probability by computing a weighted average of the negative log probabilities of generated tokens, where weights are based on token relevance to the overall text. Using a similarity function and token relevance function , the uncertainty estimate is calculated as:
where
This measure is central for computing SAR uncertainty measure.
D.2 Consistency-Based Methods
Consistency-based methods assess the uncertainty of a language model by evaluating the semantic consistency of its predictions across multiple outputs for the same prompt. The core idea is that semantically similar outputs indicate higher confidence, while diverse or conflicting outputs suggest greater uncertainty. Since language models can express the same meaning in different surface forms, these methods construct a semantic similarity matrix , where each entry represents the degree of similarity between pairs of responses. By clustering responses into groups with equivalent meanings, these methods provide a semantic measure of the model’s consistency.
The work Lin et al. [2024] offers two similarity measures to evaluate the similarity of sequences. The first is the Jaccard similarity, which treats sequences as sets of words and calculates the proportion of shared words to the total number of unique words in both sequences: .
Natural Language Inference (NLI) provides another method for computing similarity between sequences. We use the DeBERTa-large NLI model He et al. [2021], following Kuhn et al. [2023]. For each pair of sequences, an NLI model predicts two probabilities: , indicating entailment, and , indicating contradiction. Similarity is then defined as either or .
Among the simplest consistency-based approaches are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian [Lin et al., 2024]. Number of Semantic Sets estimates how many distinct “meanings” the model produces by clustering its outputs with an NLI model. The number of semantic sets is initially equal to the total number of generated answers, . Two sentences are grouped into the same cluster if the following conditions are satisfied: . This computation is performed for all pairs of answers, and the final number of distinct clusters is denoted by .
Sum of Eigenvalues of the Graph Laplacian examines global diversity: it constructs a similarity matrix among the sampled outputs and computes a continuous uncertainty score from the eigenvalues of the Laplacian of that similarity graph. The work Lin et al. [2024] proposes computing an averaged similarity matrix as . The Laplacian for the matrix is defined as , where is a (diagonal) degree matrix with elements . Consequently, the following formula is derived:
Both Number of Semantic Sets and Sum of Eigenvalues of the Graph Laplacian effectively capture overall variation in generated text but cannot produce an individual uncertainty score for each output. To address this, the work Lin et al. [2024] proposes to use the diagonal Degree Matrix which represents the total similarity of each answer with all others. The corrected trace of provides an average pairwise distance between answers, and uncertainty is computed as:
D.3 Information-Based Methods with Repeated Sampling
The natural idea is to somehow benefit from having multiple samples from the model while using important information contained in the output probabilities estimated by an LLM. Below, we examine several approaches that have sought to achieve this.
Averaging uncertainties.
We can compute the entropy on the sequence level , where the expectation is taken over the sequences randomly generated from the distribution . Unfortunately, while for token level, we have an exact way of computing the entropy, for the sequence level, we need to adhere to some approximations. In practice, we can use Monte-Carlo integration, i.e. generate several sequences via random sampling and compute Monte Carlo Sequence Entropy [Kuhn et al., 2023]:
We can replace with its length-normalized version leading to a more reliable uncertainty measure in some cases.
While simple averaging represents a natural way to aggregate uncertainties, it has certain issues related to the nature of LLMs. First of all, in the vast majority of applications, an LLM-based system should produce a single output for an input query. When we consider or other similar measure, we essentially perform averaging of uncertainties of different sequences, thus somewhat assessing the uncertainty related to the entire generative distribution for the input , but not for a particular generated sequence . This averaged uncertainty might not be adequate for this particular sequence and, remarkably, often performs worse than the uncertainty , which is related solely to the output . Moreover, although intuitive, this naïve aggregation method assumes that all outputs contribute equally to the final uncertainty estimate, regardless of their semantic relationships. This can lead to inconsistencies when semantically equivalent outputs have varying uncertainty scores or when outputs with low similarity are treated as equally important.
Semantically weighted averaging. Semantic Entropy Kuhn et al. [2023] addresses the issue of generated sequences with similar meanings but differing probabilities according to the model, which can heavily influence the resulting entropy value (25). The method clusters generated sequences into semantically homogeneous groups (where ) using a bi-directional entailment algorithm. Probabilities of sequences are averaged within each cluster. The entropy estimate is then defined as:
where represents the aggregated probability for cluster .
SentenceSAR Duan et al. [2024] enhances the probability of sentences that are more relevant. It uses a sentence relevance measure to evaluate the relevance of with respect to . SentenceSAR is calculated as:
where is a temperature parameter used to control the scale of shifting to relevance, and
The combination of SentenceSAR and TokenSAR results in a unified method called SAR Duan et al. [2024]. In this approach, the generative probability in the SentenceSAR formula is replaced with the token-shifted probability , creating a comprehensive measure that integrates both sentence- and token-level adjustments.
The aggregation approaches like Semantic Entropy Kuhn et al. [2023] or SAR Duan et al. [2024] can be unified into a semantically-aware Generalized Monte Carlo uncertainty estimate, defined as
Here, the inner summation aggregates sequence probabilities weighted by their semantic similarity to the -th output, and the outer summation averages these contributions across all samples. The function provides an additional layer of flexibility, transforming the reweighted uncertainty scores, making the method a generalized framework for uncertainty quantification.
Unfortunately, methods that fall under GMCU, while offering benefits, also inherit the aforementioned issues from both categories of methods. In particular, the outer summation in (29), similarly to the case of simple Monte Carlo averaging, often fails to outperform the uncertainty of a single generated sequence .
Appendix E Detailed Experimental Results
Tables 3 and 1 presented PRR scores averaged over datasets corresponding to each task. Here we present expanded results for each dataset, see Tables 11, 12 and 13.
| Method | Dataset | ||||||
| XSUM | WMT14FrEn | WMT19DeEn | CoQA | TriviaQA | MMLU | GSM8k | |
| Mistral7b-Base | |||||||
| MCSE | 0.007 | 0.257 | 0.350 | 0.247 | 0.496 | 0.337 | 0.475 |
| MCNSE | 0.009 | 0.342 | 0.478 | 0.238 | 0.540 | 0.356 | 0.401 |
| Semantic Entropy | 0.008 | 0.271 | 0.382 | 0.271 | 0.562 | 0.387 | 0.472 |
| DegMat | 0.137 | 0.229 | 0.382 | 0.336 | 0.646 | 0.410 | 0.299 |
| EigValLaplacian | 0.132 | 0.207 | 0.328 | 0.301 | 0.624 | 0.398 | 0.241 |
| SAR | 0.094 | 0.353 | 0.517 | 0.313 | 0.644 | 0.419 | 0.471 |
| Consistency Light | -0.006 | 0.317 | 0.457 | 0.306 | 0.597 | 0.432 | 0.441 |
| Consistency | 0.051 | 0.285 | 0.500 | 0.379 | 0.647 | 0.423 | 0.435 |
| MSP | 0.287 | 0.315 | 0.451 | 0.326 | 0.628 | 0.474 | 0.471 |
| 0.330 | 0.396 | 0.598 | 0.383 | 0.670 | 0.466 | 0.517 | |
| Light | 0.309 | 0.380 | 0.580 | 0.351 | 0.674 | 0.480 | 0.506 |
| PPL | 0.204 | 0.365 | 0.489 | 0.281 | 0.632 | 0.474 | 0.311 |
| 0.286 | 0.375 | 0.568 | 0.369 | 0.674 | 0.466 | 0.467 | |
| Light | 0.260 | 0.404 | 0.557 | 0.326 | 0.684 | 0.480 | 0.428 |
| MTE | 0.182 | 0.392 | 0.484 | 0.243 | 0.619 | 0.456 | 0.350 |
| 0.288 | 0.374 | 0.564 | 0.355 | 0.673 | 0.447 | 0.491 | |
| Light | 0.254 | 0.415 | 0.550 | 0.303 | 0.672 | 0.465 | 0.458 |
| Llama8b-Base | |||||||
| MCSE | 0.033 | 0.293 | 0.354 | 0.237 | 0.482 | 0.171 | 0.351 |
| MCNSE | 0.022 | 0.370 | 0.415 | 0.219 | 0.501 | 0.170 | 0.344 |
| Semantic Entropy | 0.033 | 0.297 | 0.389 | 0.272 | 0.549 | 0.229 | 0.375 |
| DegMat | 0.081 | 0.250 | 0.355 | 0.353 | 0.622 | 0.342 | 0.309 |
| EigValLaplacian | 0.079 | 0.198 | 0.278 | 0.332 | 0.604 | 0.292 | 0.273 |
| SAR | 0.077 | 0.427 | 0.483 | 0.311 | 0.595 | 0.352 | 0.398 |
| Consistency Light | -0.022 | 0.475 | 0.441 | 0.353 | 0.552 | 0.279 | 0.378 |
| Consistency | 0.024 | 0.389 | 0.453 | 0.375 | 0.614 | 0.391 | 0.368 |
| MSP | 0.328 | 0.342 | 0.456 | 0.277 | 0.526 | 0.508 | 0.324 |
| 0.378 | 0.456 | 0.582 | 0.349 | 0.597 | 0.485 | 0.372 | |
| Light | 0.358 | 0.448 | 0.556 | 0.340 | 0.598 | 0.504 | 0.353 |
| PPL | 0.369 | 0.351 | 0.422 | 0.253 | 0.507 | 0.461 | 0.303 |
| 0.387 | 0.448 | 0.514 | 0.338 | 0.593 | 0.452 | 0.433 | |
| Light | 0.382 | 0.483 | 0.491 | 0.319 | 0.597 | 0.467 | 0.398 |
| MTE | 0.357 | 0.357 | 0.408 | 0.239 | 0.497 | 0.349 | 0.326 |
| 0.380 | 0.446 | 0.511 | 0.337 | 0.601 | 0.401 | 0.447 | |
| Light | 0.372 | 0.501 | 0.487 | 0.316 | 0.599 | 0.387 | 0.412 |
| Falcon7b-Base | |||||||
| MCSE | 0.159 | 0.297 | 0.337 | 0.258 | 0.549 | 0.420 | 0.427 |
| MCNSE | 0.108 | 0.371 | 0.474 | 0.293 | 0.586 | 0.442 | 0.299 |
| Semantic Entropy | 0.164 | 0.307 | 0.389 | 0.294 | 0.581 | 0.463 | 0.418 |
| DegMat | 0.201 | 0.274 | 0.431 | 0.407 | 0.651 | 0.480 | 0.395 |
| EigValLaplacian | 0.201 | 0.229 | 0.394 | 0.381 | 0.645 | 0.454 | 0.358 |
| SAR | 0.144 | 0.398 | 0.517 | 0.381 | 0.649 | 0.508 | 0.387 |
| Consistency Light | 0.232 | 0.483 | 0.468 | 0.327 | 0.657 | 0.362 | 0.363 |
| Consistency | 0.226 | 0.337 | 0.496 | 0.408 | 0.656 | 0.485 | 0.426 |
| MSP | 0.201 | 0.312 | 0.400 | 0.321 | 0.662 | 0.539 | 0.377 |
| 0.257 | 0.433 | 0.578 | 0.396 | 0.684 | 0.529 | 0.436 | |
| Light | 0.242 | 0.463 | 0.566 | 0.349 | 0.691 | 0.522 | 0.405 |
| PPL | 0.155 | 0.375 | 0.525 | 0.316 | 0.644 | 0.539 | 0.326 |
| 0.229 | 0.436 | 0.580 | 0.406 | 0.677 | 0.529 | 0.478 | |
| Light | 0.234 | 0.497 | 0.558 | 0.364 | 0.694 | 0.522 | 0.468 |
| MTE | 0.152 | 0.409 | 0.537 | 0.291 | 0.633 | 0.533 | 0.367 |
| 0.228 | 0.439 | 0.577 | 0.395 | 0.685 | 0.517 | 0.510 | |
| Light | 0.234 | 0.518 | 0.549 | 0.345 | 0.693 | 0.491 | 0.500 |
| Method | Dataset | ||||||
| XSUM | WMT14FrEn | WMT19DeEn | CoQA | TriviaQA | MMLU | GSM8k | |
| Mistral7b-Base | |||||||
| MCSE | 0.162 | 0.406 | 0.407 | 0.289 | 0.492 | 0.339 | 0.693 |
| MCNSE | 0.196 | 0.471 | 0.507 | 0.277 | 0.529 | 0.358 | 0.700 |
| Semantic Entropy | 0.164 | 0.434 | 0.442 | 0.312 | 0.554 | 0.389 | 0.675 |
| DegMat | 0.205 | 0.439 | 0.410 | 0.376 | 0.618 | 0.410 | 0.454 |
| EigValLaplacian | 0.197 | 0.388 | 0.344 | 0.342 | 0.600 | 0.399 | 0.361 |
| SAR | 0.175 | 0.563 | 0.590 | 0.347 | 0.620 | 0.421 | 0.780 |
| Consistency | 0.071 | 0.670 | 0.708 | 0.405 | 0.614 | 0.423 | 0.846 |
| MSP | 0.330 | 0.212 | 0.291 | 0.388 | 0.607 | 0.476 | 0.307 |
| 0.366 | 0.712 | 0.730 | 0.430 | 0.644 | 0.466 | 0.562 | |
| PPL | 0.365 | 0.695 | 0.676 | 0.327 | 0.615 | 0.476 | 0.931 |
| 0.372 | 0.735 | 0.755 | 0.402 | 0.648 | 0.466 | 0.937 | |
| MTE | 0.350 | 0.668 | 0.606 | 0.254 | 0.594 | 0.457 | 0.932 |
| 0.373 | 0.708 | 0.732 | 0.373 | 0.645 | 0.447 | 0.935 | |
| Llama8b-Base | |||||||
| MCSE | 0.192 | 0.366 | 0.395 | 0.258 | 0.465 | 0.158 | 0.545 |
| MCNSE | 0.186 | 0.377 | 0.480 | 0.239 | 0.484 | 0.165 | 0.631 |
| Semantic Entropy | 0.194 | 0.371 | 0.451 | 0.286 | 0.528 | 0.213 | 0.557 |
| DegMat | 0.191 | 0.274 | 0.409 | 0.366 | 0.606 | 0.320 | 0.396 |
| EigValLaplacian | 0.190 | 0.216 | 0.333 | 0.339 | 0.587 | 0.274 | 0.351 |
| SAR | 0.159 | 0.441 | 0.571 | 0.327 | 0.578 | 0.340 | 0.667 |
| Consistency | 0.030 | 0.473 | 0.598 | 0.394 | 0.600 | 0.353 | 0.793 |
| MSP | 0.464 | 0.339 | 0.413 | 0.304 | 0.514 | 0.483 | 0.280 |
| 0.484 | 0.529 | 0.685 | 0.384 | 0.587 | 0.452 | 0.513 | |
| PPL | 0.458 | 0.504 | 0.622 | 0.293 | 0.483 | 0.441 | 0.911 |
| 0.450 | 0.544 | 0.689 | 0.364 | 0.573 | 0.422 | 0.924 | |
| MTE | 0.449 | 0.437 | 0.501 | 0.238 | 0.458 | 0.326 | 0.883 |
| 0.451 | 0.520 | 0.638 | 0.345 | 0.582 | 0.363 | 0.899 | |
| Falcon7b-Base | |||||||
| MCSE | 0.128 | 0.399 | 0.419 | 0.285 | 0.535 | 0.421 | 0.598 |
| MCNSE | 0.153 | 0.395 | 0.452 | 0.318 | 0.588 | 0.443 | 0.771 |
| Semantic Entropy | 0.134 | 0.420 | 0.460 | 0.319 | 0.566 | 0.463 | 0.567 |
| DegMat | 0.177 | 0.350 | 0.422 | 0.422 | 0.637 | 0.480 | 0.633 |
| EigValLaplacian | 0.174 | 0.289 | 0.382 | 0.393 | 0.622 | 0.454 | 0.522 |
| SAR | 0.193 | 0.455 | 0.521 | 0.385 | 0.642 | 0.509 | 0.826 |
| Consistency | 0.282 | 0.491 | 0.651 | 0.416 | 0.627 | 0.484 | 0.979 |
| MSP | 0.099 | 0.385 | 0.378 | 0.369 | 0.638 | 0.540 | -0.175 |
| 0.210 | 0.564 | 0.698 | 0.428 | 0.659 | 0.530 | 0.498 | |
| PPL | 0.275 | 0.541 | 0.637 | 0.353 | 0.614 | 0.540 | 1.000 |
| 0.310 | 0.579 | 0.717 | 0.415 | 0.644 | 0.530 | 1.000 | |
| MTE | 0.186 | 0.475 | 0.510 | 0.317 | 0.573 | 0.534 | 0.984 |
| 0.289 | 0.551 | 0.678 | 0.402 | 0.646 | 0.517 | 0.998 | |
| Method | Dataset | ||||||
| XSUM | WMT14FrEn | WMT19DeEn | CoQA | TriviaQA | MMLU | GSM8k | |
| Mistral7b-Base | |||||||
| MCSE | 0.123 | 0.297 | 0.363 | 0.245 | 0.515 | 0.385 | 0.751 |
| MCNSE | 0.105 | 0.385 | 0.485 | 0.256 | 0.552 | 0.406 | 0.548 |
| Semantic Entropy | 0.127 | 0.310 | 0.395 | 0.272 | 0.580 | 0.439 | 0.735 |
| DegMat | 0.244 | 0.280 | 0.365 | 0.358 | 0.673 | 0.461 | 0.392 |
| EigValLaplacian | 0.243 | 0.238 | 0.319 | 0.334 | 0.654 | 0.443 | 0.318 |
| SAR | 0.204 | 0.392 | 0.503 | 0.334 | 0.658 | 0.478 | 0.656 |
| Consistency | 0.211 | 0.297 | 0.437 | 0.355 | 0.659 | 0.458 | 0.603 |
| MSP | 0.259 | 0.381 | 0.486 | 0.216 | 0.346 | 0.514 | 0.711 |
| 0.295 | 0.441 | 0.589 | 0.336 | 0.597 | 0.514 | 0.766 | |
| PPL | 0.205 | 0.391 | 0.513 | 0.186 | 0.358 | 0.514 | 0.672 |
| 0.258 | 0.451 | 0.577 | 0.322 | 0.612 | 0.514 | 0.768 | |
| MTE | 0.156 | 0.443 | 0.550 | 0.234 | 0.598 | 0.508 | 0.717 |
| 0.236 | 0.419 | 0.553 | 0.337 | 0.680 | 0.494 | 0.783 | |
| Llama8b-Base | |||||||
| MCSE | 0.089 | 0.240 | 0.354 | 0.220 | 0.444 | 0.297 | 0.647 |
| MCNSE | 0.055 | 0.334 | 0.443 | 0.247 | 0.470 | 0.306 | 0.545 |
| Semantic Entropy | 0.089 | 0.250 | 0.382 | 0.245 | 0.514 | 0.361 | 0.650 |
| DegMat | 0.217 | 0.218 | 0.361 | 0.368 | 0.611 | 0.391 | 0.539 |
| EigValLaplacian | 0.217 | 0.167 | 0.280 | 0.347 | 0.592 | 0.305 | 0.492 |
| SAR | 0.125 | 0.393 | 0.520 | 0.345 | 0.578 | 0.472 | 0.642 |
| Consistency | 0.146 | 0.290 | 0.440 | 0.387 | 0.603 | 0.428 | 0.599 |
| MSP | 0.197 | 0.344 | 0.463 | 0.161 | 0.247 | 0.577 | 0.675 |
| 0.225 | 0.403 | 0.590 | 0.304 | 0.487 | 0.586 | 0.731 | |
| PPL | 0.168 | 0.382 | 0.460 | 0.161 | 0.253 | 0.549 | 0.668 |
| 0.198 | 0.428 | 0.580 | 0.320 | 0.496 | 0.562 | 0.751 | |
| MTE | 0.130 | 0.412 | 0.509 | 0.245 | 0.466 | 0.459 | 0.699 |
| 0.189 | 0.398 | 0.561 | 0.345 | 0.587 | 0.472 | 0.765 | |
| Falcon7b-Base | |||||||
| MCSE | 0.188 | 0.331 | 0.362 | 0.274 | 0.560 | 0.455 | 0.460 |
| MCNSE | 0.126 | 0.387 | 0.494 | 0.306 | 0.605 | 0.479 | 0.396 |
| Semantic Entropy | 0.192 | 0.332 | 0.414 | 0.304 | 0.594 | 0.497 | 0.438 |
| DegMat | 0.246 | 0.271 | 0.470 | 0.392 | 0.672 | 0.516 | 0.415 |
| EigValLaplacian | 0.238 | 0.220 | 0.432 | 0.365 | 0.660 | 0.476 | 0.339 |
| SAR | 0.184 | 0.412 | 0.543 | 0.370 | 0.664 | 0.553 | 0.536 |
| Consistency | 0.235 | 0.323 | 0.498 | 0.376 | 0.665 | 0.523 | 0.588 |
| MSP | 0.187 | 0.355 | 0.398 | 0.213 | 0.458 | 0.583 | 0.367 |
| 0.248 | 0.461 | 0.590 | 0.336 | 0.617 | 0.575 | 0.472 | |
| PPL | 0.177 | 0.371 | 0.504 | 0.201 | 0.439 | 0.583 | 0.556 |
| 0.242 | 0.462 | 0.613 | 0.345 | 0.624 | 0.575 | 0.662 | |
| MTE | 0.159 | 0.462 | 0.580 | 0.290 | 0.588 | 0.565 | 0.517 |
| 0.237 | 0.454 | 0.600 | 0.373 | 0.666 | 0.557 | 0.650 | |
Appendix F Alternative Performance Metrics for PRR
The choice of PRR as a UQ quality metric of choice is dictated by its ability to handle both continuous performance metrics, like Comet and AlignScore without the need for selecting arbitrary thresholds, as well as relative robustness to class imbalance in case of binary performance metric. However, PRR scores are calculated for a particular choice of underlying performance metric. For a truly comprehensive evaluation, we perform the same evaluation as in our main experimental run, but with alternative choice of performance metrics.
Tables 14 and 15 report PRRs for these metrics on all models for NMT and QA tasks. MetricX Juraska et al. [2024] was used for NMT and GPT-as-a-judge (gpt-4o-2024-08-06) was used to score QA datasets. The following is the prompt used to facilitate GPT QA scoring:
You are a text evaluator. The model was asked the following question:
{question}
The ’Generated’ text is a model’s response. The ’Target’ is the correct answer.
If the generated answer correctly answers the question based on the target, return 1.
If it is wrong, return 0.
Respond ONLY with a single digit: 1 or 0.
Generated: {model output}
Target: {target sequence}
| NMT (MetricX) | QA (GPT-4o) | |||||
| Method | WMT14FrEn | WMT19DeEn | CoQA/Gpt | TriviaQA/Gpt | MMLU/Gpt | GSM8k/Gpt |
| Mistral7b-Base | ||||||
| MCSE | 0.127 | 0.212 | 0.226 | 0.478 | 0.338 | 0.453 |
| MCNSE | 0.267 | 0.417 | 0.246 | 0.535 | 0.357 | 0.388 |
| Semantic Entropy | 0.158 | 0.261 | 0.254 | 0.554 | 0.387 | 0.454 |
| CEDegMat | 0.270 | 0.402 | 0.336 | 0.659 | 0.402 | 0.356 |
| SAR | 0.302 | 0.451 | 0.333 | 0.660 | 0.419 | 0.456 |
| Semantic Density | 0.291 | 0.366 | 0.056 | 0.688 | 0.335 | 0.195 |
| MSP | 0.173 | 0.287 | 0.287 | 0.640 | 0.473 | 0.485 |
| 0.314 | 0.462 | 0.367 | 0.681 | 0.465 | 0.537 | |
| PPL | 0.316 | 0.448 | 0.220 | 0.645 | 0.473 | 0.345 |
| 0.361 | 0.512 | 0.331 | 0.687 | 0.465 | 0.499 | |
| MTE | 0.350 | 0.455 | 0.190 | 0.631 | 0.456 | 0.387 |
| 0.366 | 0.512 | 0.323 | 0.684 | 0.447 | 0.524 | |
| Llama8b-Base | ||||||
| MCSE | 0.145 | 0.181 | 0.200 | 0.486 | 0.158 | 0.332 |
| MCNSE | 0.319 | 0.361 | 0.183 | 0.530 | 0.170 | 0.324 |
| Semantic Entropy | 0.157 | 0.235 | 0.239 | 0.552 | 0.203 | 0.361 |
| CEDegMat | 0.287 | 0.367 | 0.350 | 0.621 | 0.299 | 0.300 |
| SAR | 0.354 | 0.420 | 0.302 | 0.610 | 0.319 | 0.382 |
| Semantic Density | 0.252 | 0.328 | 0.069 | 0.650 | 0.335 | 0.299 |
| MSP | 0.176 | 0.274 | 0.228 | 0.564 | 0.469 | 0.313 |
| 0.329 | 0.451 | 0.332 | 0.625 | 0.449 | 0.365 | |
| PPL | 0.327 | 0.388 | 0.217 | 0.549 | 0.462 | 0.309 |
| 0.395 | 0.469 | 0.329 | 0.623 | 0.443 | 0.445 | |
| MTE | 0.353 | 0.388 | 0.214 | 0.543 | 0.358 | 0.329 |
| 0.394 | 0.472 | 0.333 | 0.634 | 0.393 | 0.455 | |
| Falcon7b-Base | ||||||
| MCSE | 0.168 | 0.218 | 0.228 | 0.587 | 0.420 | 0.443 |
| MCNSE | 0.304 | 0.399 | 0.265 | 0.640 | 0.442 | 0.323 |
| Semantic Entropy | 0.201 | 0.276 | 0.262 | 0.623 | 0.463 | 0.448 |
| CEDegMat | 0.287 | 0.428 | 0.380 | 0.712 | 0.476 | 0.347 |
| SAR | 0.339 | 0.457 | 0.367 | 0.715 | 0.508 | 0.410 |
| Semantic Density | 0.323 | 0.422 | 0.076 | 0.746 | 0.396 | 0.349 |
| MSP | 0.176 | 0.275 | 0.279 | 0.737 | 0.539 | 0.391 |
| 0.335 | 0.483 | 0.364 | 0.763 | 0.529 | 0.459 | |
| PPL | 0.322 | 0.454 | 0.245 | 0.722 | 0.539 | 0.341 |
| 0.394 | 0.531 | 0.345 | 0.754 | 0.529 | 0.513 | |
| MTE | 0.370 | 0.455 | 0.228 | 0.707 | 0.533 | 0.385 |
| 0.403 | 0.529 | 0.341 | 0.762 | 0.516 | 0.551 | |
| NMT (MetricX) | QA (GPT-4o) | |||||
| Method | WMT14FrEn | WMT19DeEn | CoQA/Gpt | TriviaQA/Gpt | MMLU/Gpt | GSM8k/Gpt |
| Mistral7b-Base | ||||||
| MCSE | 0.305 | 0.324 | 0.244 | 0.464 | 0.339 | 0.392 |
| MCNSE | 0.360 | 0.466 | 0.252 | 0.516 | 0.358 | 0.188 |
| Semantic Entropy | 0.336 | 0.366 | 0.272 | 0.536 | 0.389 | 0.420 |
| CEDegMat | 0.468 | 0.485 | 0.339 | 0.623 | 0.404 | 0.239 |
| SAR | 0.455 | 0.527 | 0.336 | 0.628 | 0.421 | 0.298 |
| Semantic Density | 0.565 | 0.554 | 0.043 | 0.646 | 0.306 | 0.421 |
| MSP | 0.169 | 0.252 | 0.311 | 0.608 | 0.474 | 0.460 |
| 0.607 | 0.623 | 0.376 | 0.647 | 0.466 | 0.550 | |
| PPL | 0.510 | 0.552 | 0.265 | 0.619 | 0.474 | 0.096 |
| 0.620 | 0.667 | 0.346 | 0.655 | 0.466 | 0.091 | |
| MTE | 0.487 | 0.499 | 0.211 | 0.604 | 0.457 | 0.167 |
| 0.581 | 0.638 | 0.324 | 0.653 | 0.447 | 0.151 | |
| Llama8b-Base | ||||||
| MCSE | 0.284 | 0.271 | 0.191 | 0.451 | 0.155 | 0.302 |
| MCNSE | 0.354 | 0.419 | 0.197 | 0.493 | 0.171 | 0.251 |
| Semantic Entropy | 0.299 | 0.333 | 0.230 | 0.515 | 0.201 | 0.341 |
| CEDegMat | 0.343 | 0.451 | 0.369 | 0.598 | 0.301 | 0.241 |
| SAR | 0.412 | 0.497 | 0.315 | 0.584 | 0.323 | 0.311 |
| Semantic Density | 0.403 | 0.452 | 0.065 | 0.638 | 0.286 | 0.396 |
| MSP | 0.247 | 0.289 | 0.225 | 0.523 | 0.467 | 0.332 |
| 0.473 | 0.568 | 0.346 | 0.597 | 0.440 | 0.418 | |
| PPL | 0.437 | 0.517 | 0.238 | 0.498 | 0.459 | 0.185 |
| 0.510 | 0.614 | 0.335 | 0.588 | 0.433 | 0.199 | |
| MTE | 0.387 | 0.439 | 0.201 | 0.485 | 0.345 | 0.226 |
| 0.495 | 0.589 | 0.330 | 0.604 | 0.372 | 0.237 | |
| Falcon7b-Base | ||||||
| MCSE | 0.269 | 0.341 | 0.261 | 0.553 | 0.421 | 0.159 |
| MCNSE | 0.325 | 0.403 | 0.312 | 0.620 | 0.443 | 0.064 |
| Semantic Entropy | 0.308 | 0.397 | 0.296 | 0.588 | 0.463 | 0.172 |
| CEDegMat | 0.381 | 0.478 | 0.398 | 0.686 | 0.477 | 0.083 |
| SAR | 0.389 | 0.483 | 0.398 | 0.685 | 0.509 | 0.152 |
| Semantic Density | 0.434 | 0.518 | 0.047 | 0.661 | 0.377 | 0.299 |
| MSP | 0.269 | 0.311 | 0.340 | 0.683 | 0.540 | 0.104 |
| 0.465 | 0.621 | 0.407 | 0.705 | 0.530 | 0.300 | |
| PPL | 0.430 | 0.553 | 0.308 | 0.663 | 0.540 | 0.156 |
| 0.526 | 0.655 | 0.376 | 0.691 | 0.530 | 0.128 | |
| MTE | 0.418 | 0.452 | 0.275 | 0.615 | 0.535 | 0.054 |
| 0.512 | 0.620 | 0.370 | 0.692 | 0.517 | 0.050 | |
Appendix G AUROC
We strongly believe that due to the considerations presented in Appendix F, PRR is a superior metric to AUROC when comparing relative performance of UQ methods. However, we acknowledge that AUROC is widely used in modern UQ literature, so we opt to include AUROC results on QA datasets. The results are presented in Tables 16 and 17. All results here were obtained with GPT-as-a-judge correctness scoring.
| CoQA/Gpt | TriviaQA/Gpt | MMLU/Gpt | GSM8k/Gpt | |
| Mistral7b-Base | ||||
| MCSE | 0.642 | 0.784 | 0.737 | 0.716 |
| MCNSE | 0.653 | 0.806 | 0.744 | 0.684 |
| Semantic Entropy | 0.653 | 0.817 | 0.756 | 0.716 |
| CEDegMat | 0.690 | 0.859 | 0.759 | 0.682 |
| SAR | 0.694 | 0.862 | 0.770 | 0.717 |
| Semantic Density | 0.531 | 0.871 | 0.713 | 0.576 |
| MSP | 0.674 | 0.849 | 0.794 | 0.684 |
| 0.708 | 0.867 | 0.790 | 0.729 | |
| PPL | 0.636 | 0.855 | 0.794 | 0.628 |
| 0.690 | 0.871 | 0.790 | 0.714 | |
| MTE | 0.620 | 0.849 | 0.787 | 0.640 |
| 0.687 | 0.870 | 0.783 | 0.727 | |
| Llama8b-Base | ||||
| MCSE | 0.635 | 0.779 | 0.620 | 0.705 |
| MCNSE | 0.630 | 0.799 | 0.629 | 0.689 |
| Semantic Entropy | 0.655 | 0.815 | 0.652 | 0.713 |
| CEDegMat | 0.708 | 0.845 | 0.703 | 0.691 |
| SAR | 0.694 | 0.842 | 0.715 | 0.724 |
| Semantic Density | 0.540 | 0.855 | 0.680 | 0.655 |
| MSP | 0.647 | 0.816 | 0.787 | 0.669 |
| 0.700 | 0.847 | 0.776 | 0.718 | |
| PPL | 0.640 | 0.815 | 0.778 | 0.661 |
| 0.699 | 0.848 | 0.770 | 0.745 | |
| MTE | 0.633 | 0.817 | 0.726 | 0.667 |
| 0.702 | 0.854 | 0.746 | 0.753 | |
| Falcon7b-Base | ||||
| MCSE | 0.648 | 0.807 | 0.770 | 0.766 |
| MCNSE | 0.660 | 0.831 | 0.779 | 0.709 |
| Semantic Entropy | 0.669 | 0.830 | 0.788 | 0.761 |
| CEDegMat | 0.718 | 0.870 | 0.784 | 0.722 |
| SAR | 0.713 | 0.870 | 0.805 | 0.757 |
| Semantic Density | 0.544 | 0.881 | 0.746 | 0.716 |
| MSP | 0.681 | 0.866 | 0.820 | 0.731 |
| 0.719 | 0.884 | 0.815 | 0.774 | |
| PPL | 0.645 | 0.860 | 0.820 | 0.700 |
| 0.705 | 0.882 | 0.815 | 0.800 | |
| MTE | 0.630 | 0.854 | 0.818 | 0.721 |
| 0.704 | 0.886 | 0.810 | 0.814 | |
| CoQA/Gpt | TriviaQA/Gpt | MMLU/Gpt | GSM8k/Gpt | |
| Mistral7b-Base | ||||
| MCSE | 0.650 | 0.777 | 0.738 | 0.676 |
| MCNSE | 0.656 | 0.798 | 0.745 | 0.602 |
| Semantic Entropy | 0.661 | 0.808 | 0.757 | 0.692 |
| CEDegMat | 0.693 | 0.845 | 0.760 | 0.625 |
| SAR | 0.695 | 0.849 | 0.771 | 0.643 |
| Semantic Density | 0.523 | 0.854 | 0.696 | 0.633 |
| MSP | 0.684 | 0.836 | 0.795 | 0.663 |
| 0.713 | 0.853 | 0.790 | 0.720 | |
| PPL | 0.657 | 0.843 | 0.795 | 0.592 |
| 0.697 | 0.858 | 0.790 | 0.615 | |
| MTE | 0.627 | 0.838 | 0.788 | 0.606 |
| 0.689 | 0.858 | 0.783 | 0.632 | |
| Llama8b-Base | ||||
| MCSE | 0.632 | 0.763 | 0.617 | 0.653 |
| MCNSE | 0.633 | 0.783 | 0.629 | 0.629 |
| Semantic Entropy | 0.653 | 0.800 | 0.650 | 0.669 |
| CEDegMat | 0.716 | 0.836 | 0.703 | 0.633 |
| SAR | 0.698 | 0.831 | 0.715 | 0.659 |
| Semantic Density | 0.534 | 0.849 | 0.655 | 0.666 |
| MSP | 0.647 | 0.801 | 0.787 | 0.627 |
| 0.706 | 0.836 | 0.771 | 0.686 | |
| PPL | 0.648 | 0.796 | 0.776 | 0.635 |
| 0.702 | 0.835 | 0.765 | 0.658 | |
| MTE | 0.624 | 0.794 | 0.719 | 0.644 |
| 0.701 | 0.842 | 0.736 | 0.673 | |
| Falcon7b-Base | ||||
| MCSE | 0.662 | 0.797 | 0.770 | 0.589 |
| MCNSE | 0.682 | 0.826 | 0.779 | 0.571 |
| Semantic Entropy | 0.684 | 0.821 | 0.788 | 0.593 |
| CEDegMat | 0.726 | 0.863 | 0.785 | 0.580 |
| SAR | 0.727 | 0.862 | 0.806 | 0.611 |
| Semantic Density | 0.525 | 0.856 | 0.734 | 0.684 |
| MSP | 0.706 | 0.850 | 0.820 | 0.555 |
| 0.738 | 0.869 | 0.815 | 0.667 | |
| PPL | 0.677 | 0.844 | 0.820 | 0.657 |
| 0.719 | 0.866 | 0.815 | 0.657 | |
| MTE | 0.654 | 0.825 | 0.818 | 0.603 |
| 0.717 | 0.867 | 0.810 | 0.628 | |
Appendix H Computational Budget
Total available computational resources used to produce results in this paper amounted to 12 compute nodes each having 4xA100 40Gb GPUs. Total computational budget spent to produce results was around 400 GPU-days. Each individual combination of model and dataset amounted roughly to 16 GPU-days on average to obtain all model outputs and hidden states needed for computing the results.