### v5

# CoCoA: A Minimum Bayes Risk Framework Bridging Confidence and Consistency for Uncertainty Quantification in LLMs

Abstract
Uncertainty quantification for Large Language Models (LLMs) encompasses a diverse range of approaches, with two major families being particularly prominent: (i) information-based, which estimate model confidence from token-level probabilities, and (ii) consistency-based, which assess the semantic agreement among multiple outputs generated using repeated sampling. While several recent methods have sought to combine these two paradigms to improve uncertainty quantification performance, they often fail to consistently outperform simpler baselines. In this work, we revisit the foundations of uncertainty estimation through the lens of Minimum Bayes Risk decoding, establishing a direct link between uncertainty and the optimal decision-making process of LLMs. Building on these findings, we propose CoCoA, a unified framework that integrates model confidence with output consistency, yielding a family of efficient and robust uncertainty quantification methods. We evaluate CoCoA across diverse tasks, including question answering, abstractive text summarization, and machine translation, and demonstrate sizable improvements over state-of-the-art uncertainty quantification approaches.
1 Introduction
Figure 1: Example of inconsistent probabilities assigned to semantically identical answers by an LLM, demonstrating the limitation of relying solely on sequence-level information.
Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling advances in information retrieval (zhu2023large), question answering (kwiatkowski-etal-2019-natural), machine translation (kocmi-federmann-2023-large), and a broad range of other NLP applications. As these models become an integral part of our everyday life, ensuring the reliability of their outputs is crucial, especially in high-stakes scenarios where errors can have serious consequences. One way to address this challenge is through uncertainty quantification (UQ), which focuses on estimating the confidence of model predictions.
Figure 2: Illustration of our method: the LLM generates a response, evaluates the similarity to alternatives, computes the confidence, and finally combines the confidence with the similarity measure. High similarity to alternatives reduces the uncertainty, while low similarity keeps it high.
UQ for LLMs is a rapidly advancing research area, with new UQ methods emerging each year. Most novel techniques are based on two fundamental approaches: (a) information-theoretic analysis or (b) assessment of output consistency.
Information-theoretic methods quantify the confidence of a model by analyzing the probability distributions it induces for predictions (malinin2020uncertainty; fomicheva-etal-2020-unsupervised). A key limitation of these methods is that they cannot account for semantic variability across multiple possible responses for the same input. Specifically, the model may generate answers with the same meaning but with very different assigned probabilities; see Figure 1. LLMs are trained to predict the next token in a sequence based on patterns observed in vast amounts of data, resulting in varying probabilities for semantically equivalent output sequences.
In contrast, consistency-based methods directly analyze the semantic relationship between the sampled outputs (fomicheva-etal-2020-unsupervised; lin2023generating), capturing uncertainty as objective variability of meaning among the sampled outputs.
Information-theoretic and consistency-based methods have complementary strengths. For this reason, recent state-of-the-art methods aimed to unify these approaches (kuhn2023semantic; duan-etal-2024-shifting). We follow this direction and propose a way of quantifying risk as a combination of these basic measures. This results in a family of efficient and robust UQ techniques. Our approach, illustrated in Figure 2, combines the strengths of both information-based and consistency-based methods, providing a more comprehensive and accurate assessment of uncertainty.
Recently, it has been argued (wangsubjective; daheimuncertainty) that UQ in NLP tasks can be viewed through the lens of the Minimum Bayes Risk (MBR) framework. Following this, we formulate our methods as particular forms of risk functions under the MBR approach.
Our main contributions can be summarized as follows:
  • We propose a new way of quantifying risk that combines information-theoretic and consistency-based measures, and derive a family of Confidence and Consistency-based Approaches (CoCoA) to uncertainty quantification in LLMs [^1] Our code is available publicly at https://github.com/stat-ml/llm_uncertainty_cocoa.
  • We show that the consistency component can be approximated by a learned function trained on unlabeled held-out set with a negligible loss of performance, eliminating the need for costly repeated sampling from the LLM. We call this variation of our method CoCoA Light.
  • We evaluate our approaches across a variety of NLP tasks, including question answering, summarization, and machine translation. Our experiments demonstrate sizable improvements in the reliability and the robustness of UQ compared to state-of-the-art methods.
2 Background
2.1 Language Model Decoding
LLMs define a probabilistic output distribution $p(\mathbf{y}\mid\mathbf{x})$ for a given input sequence $\mathbf{x}$. The standard ways to obtain a particular output $\mathbf{y}_{*}$ given $p(\mathbf{y}\mid\mathbf{x})$ include various variants of sampling $\mathbf{y}_{*}\sim p(\mathbf{y}\mid\mathbf{x})$ and greedy decoding, where
$\mathbf{y}_{*}=\arg\max_{\mathbf{y}}p(\mathbf{y}\mid\mathbf{x}).$ (1)
An alternative approach is to use Minimum Bayes Risk (MBR) decoding (kumar2004minimum):
$\mathbf{y}_{*}=\arg\min_{\mathbf{y}\in\mathcal{Y}}R(\mathbf{y}\mid\mathbf{x}),$ (2)
where $\mathcal{Y}$ is the set of candidate sequences and $R(\mathbf{y}\mid\mathbf{x})$ is a risk function:
$R(\mathbf{y}\mid\mathbf{x})=\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\,\,r(\mathbf{y},\mathbf{y}^{\prime})$ (3)
with $r(\mathbf{y},\mathbf{y}^{\prime})$ being a pairwise loss function.
The standard choice in the MBR literature is to take $r(\mathbf{y},\mathbf{y}^{\prime})\propto-s(\mathbf{y},\mathbf{y}^{\prime})$, where the so-called utility function $s(\mathbf{y},\mathbf{y}^{\prime})$ represents some notion of similarity between generated sequences $\mathbf{y}$ and $\mathbf{y}^{\prime}$. Below we discuss how various decoding strategies and corresponding risk functions lead to well-grounded definitions of uncertainty.
2.2 From Risk to Uncertainty
Generally, an uncertainty function $U$ is a mapping that quantifies the level of uncertainty associated with the output of a model $\mathbf{y}$, conditioned on the input sequence $\mathbf{x}$, which we denote as $U(\mathbf{y}\mid\mathbf{x})$.
Bayesian uncertainty measures are well known to be strongly connected with minimum risks of various kinds (xu2022minimum; kotelevskii2025risk). Recently, this connection was revisited in the context of natural language generation (wangsubjective; daheimuncertainty), leading to the development of new MBR-based UQ methods. More precisely, one can consider the smallest achievable risk within $\mathcal{Y}$ or, equivalently, to maximum expected utility:
$U_{\text{MBR}}(\mathbf{y}_{*}\mid\mathbf{x})=\min_{\mathbf{y}\in\mathcal{Y}}R(\mathbf{y}\mid\mathbf{x})=R(\mathbf{y}_{*}\mid\mathbf{x}),$ (4)
where $\mathbf{y}_{*}$ is given by (2). Interestingly, MBR-based uncertainty (4) is applicable even to $\mathcal{Y}$ consisting of a single sequence $\mathbf{y}_{*}$.
Below, we discuss how various existing uncertainty measures can be seen as particular approximations of minimum Bayes risks for various choices of utility functions. We also derive new uncertainty measures based on the MBR framework.
Single Sequence Information-Based Methods. Information-based methods estimate the uncertainty of the generated sequence by aggregating the uncertainty scores of individual tokens. Within the framework of MBR decoding, consider $r(\mathbf{y},\mathbf{y}^{\prime})=\mathbf{1}\{\mathbf{y}^{\prime}\neq\mathbf{y}\}$. In this case, the Bayes risk corresponds to the expected zero-one loss of decoding:
$R_{0/1}(\mathbf{y}\mid\mathbf{x})=\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\mathbf{1}\{\mathbf{y}^{\prime}\neq\mathbf{y}\}=1-p(\mathbf{y}\mid\mathbf{x}).$ (5)
This leads to one of the simplest information-based uncertainty measures, Sequence Probability (SP):
$U_{\text{SP}}(\mathbf{y}_{*}\mid\mathbf{x})=1-p\bigl(\mathbf{y}_{*}\mid\mathbf{x}\bigr).$ (6)
Although widely used as an uncertainty measure across various applications, including natural language generation, its connection to statistical risk has only recently been explored in the UQ literature (kotelevskii2022nonparametric; kotelevskii2025risk; aichberger2024rethinking).
Several other measures fall into this category, including Perplexity and Mean Token Entropy (fomicheva-etal-2020-unsupervised); see Appendix D.1 for details. While using only a single sample makes them computationally efficient, these techniques face three major challenges:
  1. 1.
    LLMs only give the probability of a specific answer, even though the same meaning can often be conveyed in multiple ways. To obtain a proper probability for the meaning of an answer, we need to marginalize over its various possible rephrasings. However, this is not feasible if we generate only a single sample.
  2. 2.
    The reliability of such methods depends heavily on the underlying LLM being well-calibrated, which is a quality that is hard to define and harder to ensure.
  3. 3.
    These methods are point estimates that do not provide information about the shape of the output distribution.
Semantic Consistency-Based Methods. The aforementioned issues lead to the development of consistency-based methods based on repetitive sampling from the LLM. Consider that we have sampled a set of outputs $\bigl\{\mathbf{y}^{(i)}\bigr\}_{i=1}^{M}$, where $\mathbf{y}^{(i)}\sim p(\mathbf{y}\mid\mathbf{x})$}. Consistency-based UQ methods rely on the diversity of the answers $\mathbf{y}^{(i)}$ sampled from the LLM. The idea is that if the model outputs similar answers for the same prompt over and over again, it is confident in its predictions; otherwise, it is uncertain. These techniques do not require a probability distribution estimated by an LLM and can be applied in a black-box setting, where only the generated tokens are available. This case is quite common when LLMs are deployed as a service and are accessible through a limited API.
Formally, consistency-based methods start from defining some similarity function $s\bigl(\mathbf{y},\mathbf{y}^{\prime}\bigr)\in[0,1]$ between arbitrary LLM generations $\mathbf{y}$ and $\mathbf{y}^{\prime}$. The value $s\bigl(\mathbf{y},\mathbf{y}^{\prime}\bigr)=1$ indicates the complete equivalence between $\mathbf{y}$ and $\mathbf{y}^{\prime}$, and $s\bigl(\mathbf{y},\mathbf{y}^{\prime}\bigr)=0$ indicates that there is no similarity. The similarity could be computed in various ways. For instance, the Lexical Similarity method (fomicheva-etal-2020-unsupervised) relies on surface-form similarity, measuring the degree of word-level or phrase-level overlap between the generated texts. More advanced techniques propose various methods for taking into account the semantic similarity between the generated answers by means of hard or soft clustering (lin2023generating).
Finally, given $M$ samples from the model, standard consistency-based methods compute a similarity matrix $S$, where $s_{ij}=s\bigl(\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr)$. Then, various statistics of $S$ are computed in order to estimate the uncertainty (lin2023generating); see Appendix D.2 for more detail. Alternatively, one can focus on a particular output $\mathbf{y}_{*}$ and compute uncertainty measures based on its similarity $s_{*i}=s\bigl(\mathbf{y}_{*},\mathbf{y}^{(i)}\bigr)$ with other samples $\mathbf{y}^{(i)}$; see (lin2023generating) for more detail.
Interestingly, it was recently shown (wangsubjective; daheimuncertainty) that the latter approach has direct relation to MBR decoding. Building on the MBR framework, one can define the utility function as $r(\mathbf{y},\mathbf{y}^{\prime})=1-s(\mathbf{y},\mathbf{y}^{\prime})$ and derive the following consistency-based uncertainty score:
$U_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})=\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\bigl[1-s(\mathbf{y}_{*},\mathbf{y}^{\prime})\bigr].$ (7)
Then the Monte Carlo approximation of the minimum Bayes risk is given by
$\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})=\frac{1}{M}\sum\nolimits_{i=1}^{M}(1-s_{*i}),$ (8)
where $s_{*i}=s\bigl(\mathbf{y}_{*},\mathbf{y}^{(i)}\bigr)$.
In our ablation study below, we show that such an uncertainty measure reliably outperforms consistency-based measures that aggregate the pairwise similarities of all samples (see Appendix C.2).
A key strength of consistency-based techniques is that, by generating multiple samples and analyzing their semantic similarity, they can estimate empirical probabilities over meanings rather than over individual answers. Their main drawback is that they discard the useful information that comes from the probability distribution represented by the LLM, including estimates of the probabilities of the specific answers.
3 CoCoA: Bridging Confidence and Consistency for Better Uncertainty Quantification
Both information-based and (semantic) consistency-based methods provide grounded and useful uncertainty quantification measures. There exist approaches that bridge the gap between information-based and consistency-based methods that show great promise but lack the fundamental base; see discussion in Section 4. In what follows, we present a family of Confidence and Consistency-based Approaches (CoCoA) for UQ, offering a new way to merge information- and consistency-based measures for uncertainty quantification in LLMs via the unifying MBR-based framework.
If we only consider the consistency measure, e.g., as given in (8), we miss the information that is contained in the model confidence. Thus, we want the resulting uncertainty measure to explicitly consider both the semantic consistency and the model-based uncertainty. Let us consider a risk of the following form:
$r(\mathbf{y},\mathbf{y}^{\prime}\mid\mathbf{x})=u(\mathbf{y}\mid\mathbf{x})\cdot\bigl(1-s(\mathbf{y},\mathbf{y}^{\prime})\bigr),$ (9)
where
  • $s(\mathbf{y},\mathbf{y}^{\prime})\in[0,1]$ is any utility function used in the standard MBR decoding (usually representing semantic similarity);
  • $u(\mathbf{y}\mid\mathbf{x})\geq 0$ is the model-based uncertainty measure for the output sequence $\mathbf{y}$. For example, we can use $u(\mathbf{y}\mid\mathbf{x})=1-p(\mathbf{y}\mid\mathbf{x})$ (similarly to (6)) or $u(\mathbf{y}\mid\mathbf{x})=-\log p(\mathbf{y}\mid\mathbf{x})$.
Then, the corresponding Bayes risk is $R_{\text{CoCoA}}(\mathbf{y}\mid\mathbf{x})=u(\mathbf{y}\mid\mathbf{x})\cdot\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\,\bigl(1-s(\mathbf{y},\mathbf{y}^{\prime})\bigr)$ and for an output sequence $\mathbf{y}_{*}$, the resulting uncertainty measure becomes
$U_{\text{CoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\,\bigl(1-s(\mathbf{y}_{*},\mathbf{y}^{\prime})\bigr).$ (10)
Finally, given a set of samples $\bigl\{\mathbf{y}^{(i)}\bigr\}_{i=1}^{M}$ we obtain an empirical uncertainty estimate:
$\widehat{U}_{\text{CoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot\frac{1}{M}\sum\nolimits_{i=1}^{M}(1-s_{*i})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x}),$ (11)
where $s_{*i}=s\bigl(\mathbf{y}_{*},\mathbf{y}^{(i)}\bigr)$.
The resulting uncertainty measure integrates both global (semantic) and local (model-specific) uncertainty signals. It ensures that uncertainty is amplified for sequences $\mathbf{y}_{*}$ that are both intrinsically uncertain (high $u(\mathbf{y}_{*}\mid\mathbf{x})$) and semantically inconsistent with respect to the other samples (high $\frac{1}{M}\sum\nolimits_{i=1}^{M}(1-s_{*i})$), while keeping it low for the opposite scenario; see Figure 2 for an example.
CoCoA Light. Computing the consistency-based uncertainty measure $\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$ within $\widehat{U}_{\text{CoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})$ requires multiple samples, whoch poses significant computational overhead. To address this, we propose to approximate the behavior of $\widehat{U}_{\text{cons}}(\mathbf{x})$ with a learned function. Our method closely matches the original sampling-based measure in quality while requiring only greedy output generation during inference, eliminating the need for additional sampling. Notably, this approximation can be learned without access to ground-truth labels, relying solely on the input data and the associated uncertainty values.
Thus, the CoCoA uncertainty measure that incorporates an approximation of the consistency uncertainty $\widehat{U}^{L}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})\approx\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$ can be defined as follows:
$\widehat{U}^{L}_{\text{CoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot\widehat{U}^{L}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x}),$ (12)
We name this approach CoCoA Light. Below, we describe how this approximation can be obtained.
Assume access to a held-out set of input sequences $\mathbf{x}_{j},j=1,\dots,n$. We emphasize that this held-out set is not labeled. For each sequence, we extract the embeddings $\{e(\mathbf{y}_{j}\mid\mathbf{x}_{j})\}_{j=1}^{n}$ of the corresponding greedy model generations $\mathbf{y}_{j}$. The corresponding targets are the consistency uncertainty scores ${\widehat{U}_{\text{cons}}(\mathbf{y}_{j}\mid\mathbf{x}_{j})},j=1,\dots,n$, computed via Monte Carlo estimation of the minimum Bayes risk; see equation (8).
A lightweight auxiliary model, $\widehat{g}\bigl(e(\mathbf{y}_{*}\mid\mathbf{x})\bigr)$, is trained in a supervised fashion to map model embeddings to uncertainty scores. During inference, the model generates a greedy output $\mathbf{y}_{\text{test}}$ for a test input $\mathbf{x}_{\text{test}}$, from which the embeddings $e(\mathbf{y}_{\text{test}}\mid\mathbf{x}_{\text{test}})$ are obtained and passed to the auxiliary model to predict the uncertainty score:
$\widehat{U}^{L}_{\text{cons}}(\mathbf{y}_{\text{test}}\mid\mathbf{x}_{\text{test}})=\widehat{g}\bigl(e(\mathbf{y}_{\text{test}}\mid\mathbf{x}_{\text{test}})\bigr).$ (13)
We note that a similar approach (kossen2024semantic) has been previously proposed to approximate semantic entropy using the hidden states from the model.
4 Related Work
In this section, we review existing approaches to uncertainty quantification, as well as prior work on Minimum Bayes Risk in application to natural language generation.
Information-based methods.
These methods are one of the most commonly used. They quantify uncertainty by analyzing the probability distributions of the tokens within a given output. The simplest of these methods, Sequence Probability (SP), measures the probability of the sequence given a specific input. Perplexity is another common uncertainty measure: it gauges how well the model predicts each token in a sequence and is formally defined as the exponential of the mean negative log-likelihood of those tokens (fomicheva-etal-2020-unsupervised). The Mean Token Entropy method evaluates the token-level predictions across the entire sequence by computing the average entropy of the token probability distributions at each position (fomicheva-etal-2020-unsupervised). While these methods provide useful uncertainty estimates, they do not address the broader uncertainty inherent in the generative tasks, as they rely on a single sequence and fail to capture the diversity of the possible outputs that the model could generate for the same input.
Consistency-Based Methods.
These methods estimate uncertainty by generating multiple sequences from the model’s output distribution and analyzing the variability among the sampled outputs. These methods are particularly relevant to NLP tasks, as they account for scenarios in which multiple plausible outputs may exist for a given input. Among the simplest consistency-based methods are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian (lin2023generating), which quantify uncertainty by how many distinct “meanings” the model produces. While effective at capturing global variation, they do not yield uncertainty scores for individual responses. To overcome this, the diagonal of the Degree Matrix was proposed to assess how similar each output is to the rest, enabling per-response uncertainty quantification (lin2023generating). A related approach, Lexical Similarity, operates by calculating the average similarity of words or phrases across every pair of responses in the sample (fomicheva-etal-2020-unsupervised).
Information-Based Methods with Repeated Sampling.
More recent methods attempt to reconcile output consistency with information-based signals—i.e., they not only look at how varied outputs are, but also how probable each output is under the model’s own generation process. For instance, Semantic Entropy (kuhn2023semantic) groups outputs into semantically homogeneous clusters—capturing distinct “meanings” that might just differ in phrasing—and calculates entropy across these clusters. SentenceSAR refines this idea by defining “relevance” as the sum of pairwise similarities between a given sentence and others weighted by each sentence’s model-assigned generation probability (duan-etal-2024-shifting). SAR combines the sentence-level relevance of SentenceSAR with token-level probability adjustments, attempting a more fine-grained balance of semantic and token-level uncertainty (duan-etal-2024-shifting). Semantic Density (qiu2024semantic) evaluates uncertainty by evaluating how densely a model’s generated response is situated within the semantic space of all possible outputs, with lower density indicating higher uncertainty. One limitation of sampling-based approaches is their computational cost: generating multiple outputs from the model can be expensive. Recent work has sought to address this by learning to predict semantic uncertainty directly from hidden states, allowing for the approximation of semantic entropy without repeated sampling (kossen2024semantic). Finally, BSDetector (chen-mueller-2024-quantifying) proposes an additive composition of self-reported confidence with observed consistency, however, it relies on ability of the model to assess its own confidence in text form. Another limitation is that it requires selection of a trade-off coefficient between confidence and consistency parts.
While these probability-weighted measures can provide deeper insights, they also sometimes underperform in practice (vashurin2024benchmarkinguncertaintyquantificationmethods), especially when the model’s probability estimates are unreliable, or when important nuances of semantic diversity are lost in the weighting process. Proper balancing raw output consistency with generation likelihood remains an open problem.
Minimum Bayes Risk for LLMs.
Minimum Bayes Risk (MBR) decoding, originally used in machine translation (Kumar2004MinimumBD), has recently been applied to LLMs by incorporating a posterior over model parameters to enable uncertainty-aware generation (daheimuncertainty). Complementing this, a Bayesian framework for subjective uncertainty quantification was developed yielding improved calibration in generation tasks (wangsubjective).
5 Experiments
In this section, we present the experimental setup, the results, and the ablations.
5.1 Experimental Setup
To evaluate the effectiveness of our proposed framework, we extended the LM-Polygraph library (vashurin2024benchmarkinguncertaintyquantificationmethods; fadeeva-etal-2023-lm). Since it already includes tools for calculating other uncertainty scores, it provided a convenient and efficient environment for setting up and running experiments. The primary objective of our experiments is to evaluate whether our method offers improved performance in key tasks such as question answering (QA), text summarization (SUM), and machine translation (MT), compared to existing baselines.
Datasets. For QA, we selected diverse datasets to capture a variety of challenges: TriviaQA (joshi-etal-2017-triviaqa), an open-domain factual QA dataset; CoQA (coqa), a conversational QA benchmark requiring multi-turn contextual understanding; MMLU (mmlu), a multi-task dataset spanning 57 topics to test broad knowledge; and GSM8k (gsm8k), which focuses on grade-school math problems requiring logical reasoning. For translation, we evaluated our method on WMT14 French-English (wmt14) and WMT19 German-English (wmt19translate). Finally, for summarization, we used XSUM (xsum), a dataset of complex documents paired with concise abstractive summaries. For all datasets, we follow (vashurin2024benchmarkinguncertaintyquantificationmethods) in subset selection, prompt formatting, and few-shot example sourcing.
Figure 3: Prediction-Rejection Ratio (PRR) Curve illustrating the quality of the non-rejected predictions as a function of the rejection rate. Oracle represents the optimal rejection strategy, Random is a random rejection, and UQ is rejection based on the evaluated uncertainty quantification method.
Models.
We evaluated our method on the base versions of three open-weight language models: LLaMA 3.1 8B (touvron2023llama), Mistral 7B (mistral), and Falcon 3 7B (Falcon3). The open-weight nature of these models enables direct access to token probabilities, which is crucial for implementing our UQ method. All experiments were conducted using the base (non–instruction-tuned) variants. We additionally report results on the larger Gemma 3 12B-Base model (gemmateam2025gemma3technicalreport), with detailed analysis provided in Appendix J.
Similarity Function.
To measure the similarity between two generations, we use the RoBERTa-large cross-encoder model, fine-tuned on the Semantic Textual Similarity benchmark dataset (liu2019robertarobustlyoptimizedbert; DBLP:journals/corr/abs-1908-10084; huggingface:dataset:stsb_multi_mt). This model is widely regarded as one of the most reliable and commonly used approaches for evaluating sentence similarity. The cross-encoder processes two sequences jointly and directly outputs a similarity score ranging from 0 to 1, providing a nuanced measure. Appendix C.1 contains comparative experiments with cross-encoder and other choices of the similarity function, substantiating this choice.
Baselines.
We compare the performance of the proposed method against a diverse set of baselines and state-of-the-art UQ scores, including confidence-based, consistency-based, and hybrid approaches. For information-based approaches, we evaluate Sequence Probability (SP), Perplexity (PPL), Mean Token Entropy (MTE), Monte Carlo Sequence Entropy (MCSE), and Monte Carlo Normalized Sequence Entropy (MCNSE). In the consistency-based category, we consider the Degree Matrix (DegMat) and the Sum of Eigenvalues of the Graph Laplacian (EigValLaplacian). Finally, we include hybrid methods Semantic Entropy and SAR, as well as verbalized confidence method P(true) (kadavath2022language). All formulations for these baselines can be found in Appendix D. Finally, we evaluate the performance of $\widehat{U}_{\text{cons}}$ (Consistency) and $\widehat{U}_{\text{cons}}^{L}$ (Consistency Light) as an uncertainty measure.
Evaluation Measure.
As our evaluation measure, we chose the Prediction Rejection Ratio (PRR), which measures the effectiveness of the uncertainty scores for identifying high-quality predictions (malinin2020uncertainty). PRR operates by progressively rejecting predictions with uncertainty scores above a threshold $a$ and observing how the average quality of the remaining predictions changes (see Figure 3). It is calculated as the ratio of two areas: the area between the Prediction Rejection (PR) curves for the evaluated uncertainty score and a random baseline, and the area between the oracle (the ideal uncertainty score that perfectly ranks instances by quality) and the random baseline. Formally, PRR is defined as follows:
$PRR=\frac{\text{AUC}_{\text{unc}}-\text{AUC}_{\text{rnd}}}{\text{AUC}_{\text{oracle}}-\text{AUC}_{\text{rnd}}}.$ (14)
Higher PRR values indicate better alignment of uncertainty scores with prediction quality, approaching the performance of an oracle. To ensure practical applicability, we compute PRR only up to a rejection threshold of 50%, preventing cases where excessive rejection artificially inflates the quality measures.
For the QA datasets, we further report AUROC (see Appendix G).
Quality Measures.
PRR requires an appropriate quality measure for each specific task to effectively evaluate the model output. For question–answering tasks, we use Accuracy to directly evaluate whether the generated answers match the ground truth in short-form QA tasks (e.g., MMLU), and we use the AlignScore between the correct answer and generated sequence for assessing the performance for long-form QA tasks (zha2023alignscore). For summarization tasks, we use AlignScore to measure the alignment between the output summary and the input document. It serves as a quality indicator by evaluating the relevance and the overlap between the generated content and the source text. For translation tasks, we use COMET, as it captures both semantic adequacy and fluency, ensuring that translations are accurate and linguistically appropriate (rei-etal-2020-comet).
To further improve the comprehensiveness of the reported results, we considered alternative choices of quality measures, and report the corresponding PRR scores in Appendix F.
Generation Setup.
We discuss the generation parameters, the decoding strategy, and the sample selection procedure in depth in Appendix A. In short, we report the evaluation results in two distinct setups: greedy decoding and stochastic sampling with focus on the most probable sequence among the generated outputs (best-sample). These two setups offer the highest-quality outputs and are the most reasonable generation approaches in practice.
CoCoA Light Details.
We use a multilayer perceptron network as an auxiliary model. The features used are embeddings from the middle layers of the base LLM, which are the most informative according to recent work (chen2024inside). In particular, for Llama 3.1 8B and Mistral 7B models, we take embeddings from the 16th layer, and for Falcon 3 7B from the 14th layer. More details on the training are available in Appendix I.
Metric Llama8b-Base Mistral7b-Base Falcon7b-Base
QA NMT SUM QA NMT SUM QA NMT SUM
MCSE 0.310 0.323 0.033 0.389 0.304 0.007 0.414 0.317 0.159
MCNSE 0.309 0.393 0.022 0.384 0.410 0.009 0.405 0.422 0.108
Semantic Entropy 0.356 0.343 0.033 0.423 0.327 0.008 0.439 0.348 0.164
DegMat 0.406 0.302 0.081 0.423 0.305 0.137 0.483 0.353 0.201
EigValLaplacian 0.375 0.238 0.079 0.391 0.267 0.132 0.459 0.312 0.201
SAR 0.414 0.455 0.077 0.462 0.435 0.094 0.481 0.458 0.144
P(True) -0.064 0.042 0.058 -0.029 0.075 0.179 0.118 0.155 -0.159
Consistency 0.437 0.421 0.024 0.471 0.392 0.051 0.494 0.416 0.226
Consistency Light 0.390 0.458 -0.022 0.444 0.387 -0.006 0.427 0.476 0.232
SP 0.409 0.399 0.328 0.475 0.383 0.287 0.475 0.356 0.201
$\text{CoCoA}_{SP}$ 0.451 $\uparrow$ 0.519 $\uparrow$ 0.378 $\uparrow$ 0.509 $\uparrow$ 0.497 $\uparrow$ 0.330 $\uparrow$ 0.511 $\uparrow$ 0.505 $\uparrow$ 0.257 $\uparrow$
$\text{CoCoA}_{SP}\,\,\text{Light}$ 0.449 $\uparrow$ 0.502 $\uparrow$ 0.358 $\uparrow$ 0.503 $\uparrow$ 0.480 $\uparrow$ 0.309 $\uparrow$ 0.492 $\uparrow$ 0.514 $\uparrow$ 0.242 $\uparrow$
PPL 0.381 0.386 0.369 0.424 0.427 0.204 0.456 0.450 0.155
$\text{CoCoA}_{PPL}$ 0.454 $\uparrow$ 0.481 $\uparrow$ 0.387 $\uparrow$ 0.494 $\uparrow$ 0.472 $\uparrow$ 0.286 $\uparrow$ 0.523 $\uparrow$ 0.508 $\uparrow$ 0.229 $\uparrow$
$\text{CoCoA}_{PPL}\,\,\text{Light}$ 0.445 $\uparrow$ 0.487 $\uparrow$ 0.382 $\uparrow$ 0.479 $\uparrow$ 0.480 $\uparrow$ 0.260 $\uparrow$ 0.512 $\uparrow$ 0.528 $\uparrow$ 0.234 $\uparrow$
MTE 0.353 0.382 0.357 0.417 0.438 0.182 0.456 0.473 0.152
$\text{CoCoA}_{MTE}$ 0.447 $\uparrow$ 0.478 $\uparrow$ 0.380 $\uparrow$ 0.492 $\uparrow$ 0.469 $\uparrow$ 0.288 $\uparrow$ 0.527 $\uparrow$ 0.508 $\uparrow$ 0.228 $\uparrow$
$\text{CoCoA}_{MTE}\,\,\text{Light}$ 0.428 $\uparrow$ 0.494 $\uparrow$ 0.372 $\uparrow$ 0.475 $\uparrow$ 0.482 $\uparrow$ 0.254 $\uparrow$ 0.507 $\uparrow$ 0.533 $\uparrow$ 0.234 $\uparrow$
Table 1: Results for Evaluated Sequence – Greedy Sample: Mean PRR across datasets for each task. The best-performing method is shown in bold, and the second-best is underscored. The arrows indicate improvement in CoCoA over the base version.
5.2 Results
Table 1 shows the PRR scores under the greedy generation setup (See Appendix B for Best Sample and MBR Sample). We report aggregated PRR for each type of task – question answering, neural machine translation (NMT), and summarization (SUM) – by averaging the results across all relevant datasets (e.g., TriviaQA, MMLU, CoQA, GSM8k for QA). This aggregated score provides a concise measure of the performance for each model for each task. Detailed results for each dataset separately can be found in Appendix E.
We can see that our CoCoA methods are the best across all tasks and models. They outperform existing consistency-based and hybrid state-of-the-art approaches, like Semantic Entropy and SAR. In addition, the proposed CoCoA approach consistently surpasses the baseline UQ measures: for example, $\texttt{CoCoA}_{PPL}$ outperforms standard Perplexity, illustrating the advantage of combining token-level confidence with semantic consistency. This pattern holds for other information-based metrics as well, demonstrating that using the consistency between multiple sampled outputs reliably enhances uncertainty quantification.
5.3 Ablations
Similarity Function.
We investigate the impact of different similarity measures (see Appendix C.1). On average, the cross-encoder provides strong performance. However, for summarization tasks, AlignScore yields better results, while in long-form QA, similarity based on Natural Language Inference (NLI) sometimes outperforms the cross-encoder. We hypothesize that for generation with shorter outputs (1–2 sentences), any capable NLI or cross-encoder model is sufficient. For longer generations, approaches that estimate similarity over chunks and then aggregate scores (e.g., AlignScore) may be more effective. Nevertheless, across all tasks, the cross-encoder achieves the best overall performance and, in scenarios where tuning is not feasible, serves as a strong default choice.
Alternative Formulations.
The next section of our ablation study focuses on alternative forms of combining model confidence $u(\mathbf{y}_{*}\mid\mathbf{x})$ and consistency $\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$; see Appendix C.2. First, we consider an additive form of combining them:
$\widehat{U}_{\text{AdditiveCoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})+\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x}).$ (15)
The results show that this additive formulation does not perform as well as the multiplicative one. The additive form tends to underemphasize the interaction between the two components, which is critical for capturing the nuanced relationships between confidence and consistency.
We also consider an alternative formulation of the consistency term $\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$, as the average of the full pairwise dissimilarity. In this formulation, $\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})$ represents the average inconsistency across all samples rather than focusing solely on the dissimilarity of the evaluated sequence with the other samples. Our experiments demonstrate that this formulation is not very strong. By distributing the consistency computation across all samples, it loses focus on the specific sequence being evaluated.
Lastly, in Appendix C.2, we also consider alternative formulations of the information-based metric that do not rely on logarithmic transformations. While we primarily use logarithms due to their numerical stability, we explore an alternative approach by converting these values back to probabilities and analyzing their impact on uncertainty quantification. Our findings indicate that both formulations exhibit consistent performance and yield similar results. This suggests that while logarithmic transformations enhance numerical stability, the choice between log-based and probability-based formulations does not affect much the overall performance.
6 Limitations
While our proposed CoCoA approach demonstrates robust empirical performance, several important considerations remain.
Task and Domain Dependency. CoCoA relies on an information-based confidence score and a semantic similarity function, whose effectiveness can vary across models, tasks, and domains. In open-ended tasks like creative generation, producing diverse outputs is expected; on the other hand, tasks requiring precise reasoning can be sensitive to subtle errors that generic similarity metrics may miss. Adapting these components to specific domains remains an important direction for future work.
Limited Sample Size. CoCoA estimates consistency by sampling multiple outputs. Generating many samples is computationally expensive, and thus our experiments, e.g., most sampling-based methods, use relatively small sets. While even a few samples can yield meaningful estimates, they may miss the full diversity of model outputs for complex prompts.
Quality Metric. Finally, the CoCoA’s performance assessment depends on quality metrics (e.g., COMET for machine translation, and Accuracy for QA) that may not capture every nuance of textual outputs. Further refining or extending quality metrics to account for deeper reasoning, factual faithfulness, and stylistic appropriateness would better align uncertainty scores with real-world perceptions of model correctness.
7 Conclusion and Future Work
We presented CoCoA, a unified approach that integrates Confidence and Consistency for uncertainty quantification in LLMs. By combining confidence scores with semantic similarity between multiple sampled outputs, CoCoA offers a more holistic view of uncertainty than either approach alone. In extensive evaluations on question answering, summarization, and translation, our approach outperformed existing baselines and state-of-the-art UQ methods. Moreover, CoCoA’s flexible design allows easy adaptation to a variety of tasks and settings.
Moving forward, several directions are open for further exploration. These include incorporating more adaptive sampling strategies that efficiently capture the model output space, refining the semantic similarity functions for domain-specific tasks, like code generation or commonsense reasoning.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction accurately describe the proposed family of Confidence and Consistency-based Approaches (CoCoA) to UQ. Theoretical contributions include new methods that merge information- and consistency-based measures for UQ in LLMs. Experimental validation of the new approaches is done across different NLP tasks: question answering, summarization, and translation. It aligns with the content presented in the paper (Section 35.2, Appendix B).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section 6 discusses limitations, including relying on an information-based confidence score and a semantic similarity function, having limited sample sets, and depending on a quality metric.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper presents theoretical formulations: uncertainty measures derived from Minimum Bayes Risk (MBR) and the definition of CoCoA and its approximated variant CoCoA Light (Sections 2.2 and 3). However, we do not make any formal theoretical statements in this work.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 5.1 outlines the experimental setup, datasets and evaluation metrics. The description of datasets used for each task includes TriviaQA, CoQA, MMLU, GSM8k for Question Answering, WMT14 French-English, WMT19 German-English for machine translation, and XSUM for summarization. The models used: LLaMA 3.1 8B, Mistral 7B and Falcon 3 7B. The similarity function is specified as RoBERTa-large fine-tuned on the Semantic Textual Similarity benchmark. The evaluation metric, Prediction Rejection Ratio (PRR), is clearly defined, and task-specific quality measures such as Accuracy, AlignScore, and COMET are used. Furthermore, the training setup for the CoCoA Light method, including the use of multilayer perceptron and middle layer embeddings of the base LLM, is described. The appendix provides additional details on ablation studies, similarity metrics, and decoding strategies (Appendix C and E).
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Code is open source and is provided as an extension to the LM-Polygraph library (Section 5.1).
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Experimental settings, including model architecture (base LLaMA 3.1 8B, Mistral 7B, Falcon 3 7B), embedding layer selection for auxiliary training (16th or 14th layer depending on model), and the auxiliary model structure (multilayer perceptron) are described in Section 5.1 and Appendix A.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: The paper does not provide errorbars as it operates in (almost) deterministic scenario: we operate with pre-trained models and there is no training involved whatsoever. The only randomization involved is the one related so sampling multiple sequences for consistency-based methods. The choice to not compute errorbars here is deliberate as our inference involves heavy computations and it is not possible to conduct main experiments multiple times due to that. However, we conducted experiments on 3 different LLMs and 7 datasets, which is equivalent to 21 independent run, demonstrating the robustness and reliability of our findings. For this reason, we do not report error bars, as the results are stable.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Appendix H discusses computational resources used to produce experimental results.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research does not involve human subjects or ethically sensitive applications. We believe it conforms the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Section 1 discusses the importance of Uncertainty Quantification, highlighting its positive societal impacts (helps to reduce overreliance on automated systems and encourages a more responsible LLM usage). We do not foresee any negative negative societal impact of this work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: N/A.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The paper properly cites the original sources of models (e.g., LLaMA 3.1 8B, Mistral 7B, Falcon 3 7B), datasets (CoQA, TriviaQA, MMLU, GSM8k, XSUM, WMT14FrEn, WMT19DeEn), and tools (e.g., RoBERTa, LM-Polygraph). However, it does not explicitly mention the specific licences and terms of use for these assets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [No]
  64. Justification: The paper introduces a new family of uncertainty quantification methods called CoCoA (Confidence and Consistency-based Approaches) and an approximated variant, CoCoA Light. However, it does not release any new dataset, model, or a software tool alongside the paper, and does not provide accompanying documentation or artifacts.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The research does not involve crowdsourcing experiments or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The research does not involve crowdsourcing experiments or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: The paper centers its contributions on the use of Large Language Models (LLMs) as the core subject of study. It introduces new methods for uncertainty quantification in LLMs, specifically evaluating base models such as LLaMA 3.1 8B, Mistral 7B, and Falcon 3 7B. Apart from that LLMs were used solely to aid in editing the final manuscript and writing some of the visualization code.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Decoding Strategy and Sample Selection
Modern LLMs are capable of producing output using a wide range of decoding strategies, and it is not readily apparent which one to use as a foundation for UQ experiments. On top of that, when sampling multiple outputs stochastically, one has to decide which sample to select for comparison with the target sequence and UQ purposes.
To facilitate the choice of decoding and sample selection strategies for our experiments, we conducted an evaluation of model performance with different approaches to both. Table 2 shows average values of corresponding quality metrics for all combinations of models and datasets. We considered 4 approaches for the selection of output that subsequently is used to calculate the quality of generation:
  • Greedy decoding produces single output by selecting top-1 candidate token at each generation step, thus no further selection of sample is needed.
  • Random sample corresponds to the case where random output is selected among the number of samples produced by repeatedly prompting the model with the same question. In practice we use first generated sample, highlighting model performance when stochastic decoding is done only once.
  • Most Probable sample selects the output with highest model-assigned probability among several sampled outputs.
  • MBR sample selects the output with highest average consistency with respect to other sampled outputs.
We note that selecting a random sample from the model outputs incurs a significant drop in the quality of results on several datasets, most prominently on GSM8k. Based on these observations, we evaluate the efficacy of UQ in three setups: greedy decoding, stochastic sampling with a focus on the most probable sample and MBR decoding.
Regardless of the way the main model response was obtained, responses that were used to quantify consistency in (8) were generated via repeated stochastic sampling. In all the experiments, where stochastic sampling is involved, it was performed with temperature $T=1.0$, top-k equal to 50, and top-p equal to $1.0$.
Sample Metric Greedy Random Sample Most Probable Sample MBR Sample
Falcon7b-Base
CoQA Align Score 0.793 0.706 0.783 0.793
GSM8k Accuracy 0.776 0.313 0.205 0.419
MMLU Accuracy 0.715 0.638 0.715 0.699
TriviaQA Align Score 0.557 0.473 0.568 0.559
WMT14FrEn Comet 0.867 0.833 0.857 0.855
WMT19DeEn Comet 0.846 0.807 0.826 0.828
XSUM Align Score 0.842 0.734 0.782 0.801
Llama8b-Base
CoQA Align Score 0.756 0.671 0.743 0.766
GSM8k Accuracy 0.548 0.234 0.261 0.346
MMLU Accuracy 0.570 0.368 0.577 0.469
TriviaQA Align Score 0.686 0.625 0.687 0.699
WMT14FrEn Comet 0.863 0.819 0.852 0.844
WMT19DeEn Comet 0.870 0.816 0.854 0.845
XSUM Align Score 0.848 0.608 0.825 0.703
Mistral7b-Base
CoQA Align Score 0.793 0.695 0.777 0.790
GSM8k Accuracy 0.382 0.169 0.190 0.259
MMLU Accuracy 0.632 0.552 0.632 0.612
TriviaQA Align Score 0.743 0.655 0.750 0.739
WMT14FrEn Comet 0.863 0.812 0.830 0.841
WMT19DeEn Comet 0.864 0.805 0.836 0.837
XSUM Align Score 0.803 0.578 0.775 0.665
Table 2: Base quality metrics for models for different evaluated sequence choice.
Appendix B Results Summary for Most Probable Sample and MBR Sample strategies
Tables 3 and 4 report the PRR scores under the Most Probable Sample and MBR Sample generation setups, as discussed in Section  5.2. We observe that CoCoA-family methods, and their Supervised versions, consistently improve base uncertainty estimators.
Metric Llama8b-Base Mistral7b-Base Falcon7b-Base
QA NMT SUM QA NMT SUM QA NMT SUM
MCSE 0.356 0.380 0.192 0.453 0.406 0.162 0.460 0.409 0.128
MCNSE 0.380 0.429 0.186 0.466 0.489 0.196 0.530 0.424 0.153
Semantic Entropy 0.396 0.411 0.194 0.482 0.438 0.164 0.479 0.440 0.134
DegMat 0.422 0.342 0.191 0.465 0.425 0.205 0.543 0.386 0.177
EigValLaplacian 0.388 0.274 0.190 0.426 0.366 0.197 0.498 0.336 0.174
SAR 0.478 0.506 0.159 0.542 0.576 0.175 0.590 0.488 0.193
P(True) -0.071 0.066 0.058 0.051 0.371 0.207 0.281 0.245 0.022
Consistency 0.535 0.536 0.030 0.572 0.689 0.071 0.626 0.571 0.282
SP 0.395 0.376 0.464 0.444 0.252 0.330 0.343 0.381 0.099
$\text{CoCoA}_{SP}$ 0.484 $\uparrow$ 0.607 $\uparrow$ 0.484 $\uparrow$ 0.526 $\uparrow$ 0.721 $\uparrow$ 0.366 $\uparrow$ 0.529 $\uparrow$ 0.631 $\uparrow$ 0.210 $\uparrow$
PPL 0.532 0.563 0.458 0.587 0.686 0.365 0.627 0.589 0.275
$\text{CoCoA}_{PPL}$ 0.571 $\uparrow$ 0.617 $\uparrow$ 0.450 0.613 $\uparrow$ 0.745 $\uparrow$ 0.372 $\uparrow$ 0.647 $\uparrow$ 0.648 $\uparrow$ 0.310 $\uparrow$
MTE 0.476 0.469 0.449 0.559 0.637 0.350 0.602 0.492 0.186
$\text{CoCoA}_{MTE}$ 0.547 $\uparrow$ 0.579 $\uparrow$ 0.451 $\uparrow$ 0.600 $\uparrow$ 0.720 $\uparrow$ 0.373 $\uparrow$ 0.641 $\uparrow$ 0.614 $\uparrow$ 0.289 $\uparrow$
Table 3: Results for Evaluated Sequence – Most Probable Sample: Mean PRR across datasets for each task. The best performing method is in bold, and the second-best is underscored. Arrows indicate improvement in CoCoA over the base version.
Metric Llama8b-Base Mistral7b-Base Falcon7b-Base
QA NMT SUM QA NMT SUM QA NMT SUM
MCSE 0.402 0.297 0.158 0.474 0.330 0.237 0.437 0.346 0.076
MCNSE 0.392 0.389 0.124 0.441 0.435 0.192 0.446 0.440 0.078
Semantic Entropy 0.443 0.316 0.159 0.507 0.352 0.236 0.458 0.373 0.077
DegMat 0.478 0.290 0.047 0.471 0.322 0.124 0.499 0.371 -0.018
EigValLaplacian 0.434 0.224 0.037 0.437 0.279 0.115 0.460 0.326 -0.022
SAR 0.509 0.456 0.187 0.532 0.448 0.266 0.531 0.478 0.083
P(True) -0.049 0.094 -0.014 -0.044 0.113 0.079 0.160 0.178 -0.124
Consistency 0.504 0.365 0.235 0.518 0.367 0.284 0.538 0.410 0.099
SP 0.415 0.404 0.266 0.447 0.433 0.312 0.405 0.376 0.170
$\text{CoCoA}_{SP}$ 0.527 $\uparrow$ 0.496 $\uparrow$ 0.293 $\uparrow$ 0.553 $\uparrow$ 0.515 $\uparrow$ 0.348 $\uparrow$ 0.500 $\uparrow$ 0.526 $\uparrow$ 0.154
PPL 0.408 0.421 0.164 0.432 0.452 0.273 0.445 0.438 0.100
$\text{CoCoA}_{PPL}$ 0.532 $\uparrow$ 0.504 $\uparrow$ 0.215 $\uparrow$ 0.554 $\uparrow$ 0.514 $\uparrow$ 0.328 $\uparrow$ 0.552 $\uparrow$ 0.537 $\uparrow$ 0.109 $\uparrow$
MTE 0.467 0.461 0.148 0.515 0.496 0.219 0.490 0.521 0.069
$\text{CoCoA}_{MTE}$ 0.542 $\uparrow$ 0.479 $\uparrow$ 0.201 $\uparrow$ 0.574 $\uparrow$ 0.486 0.290 $\uparrow$ 0.562 $\uparrow$ 0.527 $\uparrow$ 0.081 $\uparrow$
Table 4: Results for Evaluated Sequence - MBR Decoding: Mean PRR across datasets for each task. The best performing method is in bold, and the second-best is underscored. Arrows indicate improvement in CoCoA over the base version.
Appendix C Ablation
C.1 Choice of Similarity Function
For sample consistency estimation, one could come up with a variety of similarity functions $s(\mathbf{y},\mathbf{y}^{\prime})$. We perform a comparison of the effectiveness of CoCoA-family methods using several such functions. We consider the following functions:
  • AlignScore [zha2023alignscore] with AlignScore-large model;
  • RougeL [lin-2004-rouge];
  • NLI [he2021deberta] based on microsoft/deberta-large-mnli model;
  • CrossEncoder [liu2019robertarobustlyoptimizedbert] based on cross-encoder/stsb-roberta-large model.
Tables 5, 6 and 7 report these results. There exists a considerable variation of relative effectiveness of proposed methods with various similarity function choices, depending on a task at hand. We opt to report all results in other sections with CrossEncoder-based similarity as it by itself provides a significant improvement over baselines, and for consistency and ease of comparison reasons. However, when applying these methods to a particular task, we encourage users to select appropriate underlying similarity function for best results.
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
Mistral7b-Base
$\text{CoCoA}_{SP}$
AlignScore 0.334 0.293 0.445 0.354 0.655 0.466 0.550
RougeL 0.289 0.358 0.546 0.369 0.649 0.466 0.536
NLI 0.308 0.313 0.477 0.403 0.677 0.470 0.568
CrossEncoder 0.327 0.397 0.595 0.381 0.671 0.466 0.505
$\text{CoCoA}_{PPL}$
AlignScore 0.307 0.308 0.489 0.373 0.666 0.466 0.536
RougeL 0.226 0.369 0.531 0.352 0.653 0.466 0.466
NLI 0.233 0.316 0.501 0.376 0.682 0.470 0.480
CrossEncoder 0.281 0.371 0.565 0.365 0.674 0.466 0.465
$\text{CoCoA}_{MTE}$
AlignScore 0.302 0.299 0.477 0.366 0.664 0.450 0.555
RougeL 0.212 0.377 0.528 0.345 0.652 0.449 0.497
NLI 0.219 0.313 0.488 0.362 0.681 0.453 0.490
CrossEncoder 0.282 0.368 0.560 0.351 0.673 0.448 0.486
Llama8b-Base
$\text{CoCoA}_{SP}$
AlignScore 0.367 0.331 0.452 0.308 0.596 0.484 0.401
RougeL 0.336 0.393 0.545 0.321 0.563 0.474 0.375
NLI 0.344 0.352 0.467 0.364 0.606 0.478 0.419
CrossEncoder 0.375 0.454 0.583 0.350 0.598 0.480 0.367
$\text{CoCoA}_{PPL}$
AlignScore 0.422 0.346 0.450 0.337 0.596 0.453 0.446
RougeL 0.370 0.408 0.486 0.319 0.552 0.441 0.418
NLI 0.374 0.354 0.438 0.348 0.600 0.446 0.409
CrossEncoder 0.380 0.444 0.514 0.339 0.593 0.447 0.429
$\text{CoCoA}_{MTE}$
AlignScore 0.419 0.340 0.438 0.339 0.605 0.411 0.459
RougeL 0.362 0.417 0.481 0.319 0.560 0.390 0.440
NLI 0.366 0.342 0.428 0.340 0.612 0.396 0.420
CrossEncoder 0.374 0.441 0.511 0.337 0.601 0.394 0.444
Falcon7b-Base
$\text{CoCoA}_{SP}$
AlignScore 0.278 0.306 0.475 0.361 0.677 0.528 0.470
RougeL 0.205 0.394 0.499 0.378 0.678 0.527 0.417
NLI 0.236 0.361 0.511 0.407 0.684 0.532 0.532
CrossEncoder 0.253 0.436 0.577 0.396 0.685 0.529 0.428
$\text{CoCoA}_{PPL}$
AlignScore 0.252 0.340 0.523 0.410 0.678 0.528 0.521
RougeL 0.170 0.409 0.537 0.389 0.668 0.527 0.439
NLI 0.193 0.364 0.531 0.408 0.680 0.532 0.499
CrossEncoder 0.226 0.437 0.579 0.405 0.677 0.529 0.474
$\text{CoCoA}_{MTE}$
AlignScore 0.253 0.337 0.519 0.403 0.683 0.515 0.554
RougeL 0.170 0.426 0.540 0.382 0.673 0.514 0.472
NLI 0.190 0.364 0.525 0.398 0.687 0.521 0.514
CrossEncoder 0.223 0.438 0.575 0.395 0.685 0.517 0.505
Table 5: Comparison of PRRs of CoCoA-family methods with different choices of the similarity function. Main model response obtained by greedy decoding.
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
Mistral7b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.393 0.448 0.491 0.399 0.626 0.467 0.476
RougeL 0.344 0.602 0.597 0.420 0.622 0.466 0.538
NLI 0.340 0.615 0.604 0.445 0.651 0.470 0.456
CrossEncoder 0.366 0.712 0.730 0.430 0.644 0.466 0.562
$\text{CoCoA}_{PPL}$
AlignScore 0.474 0.619 0.657 0.408 0.638 0.467 0.910
RougeL 0.362 0.710 0.717 0.391 0.627 0.466 0.950
NLI 0.370 0.677 0.684 0.414 0.657 0.470 0.941
CrossEncoder 0.372 0.735 0.755 0.402 0.648 0.466 0.937
$\text{CoCoA}_{MTE}$
AlignScore 0.492 0.547 0.590 0.383 0.633 0.449 0.914
RougeL 0.355 0.695 0.684 0.366 0.624 0.448 0.959
NLI 0.364 0.656 0.658 0.387 0.656 0.453 0.918
CrossEncoder 0.373 0.708 0.732 0.373 0.645 0.447 0.935
Llama8b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.520 0.332 0.491 0.354 0.587 0.457 0.401
RougeL 0.471 0.470 0.588 0.362 0.551 0.446 0.499
NLI 0.466 0.442 0.577 0.386 0.597 0.446 0.470
CrossEncoder 0.484 0.529 0.685 0.384 0.587 0.452 0.513
$\text{CoCoA}_{PPL}$
AlignScore 0.546 0.406 0.561 0.376 0.577 0.429 0.875
RougeL 0.452 0.518 0.639 0.352 0.532 0.417 0.931
NLI 0.458 0.466 0.597 0.365 0.583 0.418 0.912
CrossEncoder 0.450 0.544 0.689 0.364 0.573 0.422 0.925
$\text{CoCoA}_{MTE}$
AlignScore 0.561 0.325 0.497 0.365 0.589 0.380 0.821
RougeL 0.448 0.496 0.598 0.336 0.539 0.361 0.921
NLI 0.449 0.446 0.565 0.344 0.598 0.359 0.881
CrossEncoder 0.451 0.520 0.638 0.346 0.582 0.363 0.900
Falcon7b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.181 0.378 0.473 0.410 0.654 0.528 0.239
RougeL 0.122 0.531 0.581 0.420 0.655 0.528 0.426
NLI 0.120 0.496 0.607 0.437 0.658 0.533 0.458
CrossEncoder 0.210 0.564 0.698 0.428 0.659 0.530 0.498
$\text{CoCoA}_{PPL}$
AlignScore 0.384 0.454 0.586 0.440 0.648 0.528 0.994
RougeL 0.280 0.565 0.668 0.410 0.637 0.528 1.000
NLI 0.283 0.515 0.671 0.424 0.647 0.533 0.998
CrossEncoder 0.310 0.579 0.717 0.415 0.644 0.530 1.000
$\text{CoCoA}_{MTE}$
AlignScore 0.292 0.386 0.498 0.435 0.648 0.515 0.972
RougeL 0.222 0.545 0.607 0.400 0.633 0.515 0.998
NLI 0.201 0.498 0.636 0.415 0.645 0.521 0.987
CrossEncoder 0.289 0.551 0.678 0.402 0.646 0.517 0.998
Table 6: Comparison of PRRs of CoCoA-family methods with different choices of similarity function. Main model response obtained by selecting the most probable sample.
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
Mistral7b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.287 0.326 0.460 0.312 0.575 0.515 0.747
RougeL 0.259 0.420 0.562 0.304 0.502 0.515 0.750
NLI 0.278 0.375 0.506 0.350 0.589 0.512 0.766
CrossEncoder 0.295 0.441 0.589 0.336 0.597 0.514 0.766
$\text{CoCoA}_{PPL}$
AlignScore 0.283 0.333 0.482 0.330 0.596 0.515 0.738
RougeL 0.219 0.427 0.561 0.300 0.517 0.515 0.744
NLI 0.219 0.390 0.508 0.334 0.602 0.512 0.723
CrossEncoder 0.258 0.451 0.577 0.322 0.612 0.514 0.768
$\text{CoCoA}_{MTE}$
AlignScore 0.269 0.314 0.447 0.357 0.663 0.497 0.746
RougeL 0.181 0.446 0.571 0.327 0.622 0.497 0.786
NLI 0.182 0.370 0.482 0.348 0.686 0.499 0.704
CrossEncoder 0.236 0.419 0.553 0.337 0.680 0.494 0.783
Llama8b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.221 0.309 0.462 0.261 0.467 0.622 0.717
RougeL 0.195 0.401 0.542 0.250 0.361 0.631 0.721
NLI 0.222 0.311 0.463 0.311 0.480 0.580 0.769
CrossEncoder 0.225 0.403 0.590 0.304 0.487 0.586 0.731
$\text{CoCoA}_{PPL}$
AlignScore 0.225 0.325 0.484 0.304 0.476 0.603 0.747
RougeL 0.169 0.416 0.537 0.274 0.370 0.612 0.758
NLI 0.180 0.334 0.460 0.319 0.488 0.558 0.761
CrossEncoder 0.198 0.428 0.580 0.320 0.496 0.562 0.751
$\text{CoCoA}_{MTE}$
AlignScore 0.209 0.293 0.460 0.320 0.563 0.556 0.749
RougeL 0.148 0.436 0.559 0.314 0.497 0.587 0.785
NLI 0.156 0.298 0.445 0.340 0.592 0.460 0.740
CrossEncoder 0.189 0.398 0.561 0.345 0.587 0.472 0.765
Falcon7b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.279 0.307 0.473 0.312 0.594 0.574 0.460
RougeL 0.202 0.424 0.487 0.307 0.566 0.574 0.447
NLI 0.231 0.362 0.508 0.355 0.618 0.577 0.531
CrossEncoder 0.248 0.461 0.590 0.336 0.617 0.575 0.472
$\text{CoCoA}_{PPL}$
AlignScore 0.278 0.320 0.534 0.350 0.607 0.574 0.617
RougeL 0.190 0.428 0.556 0.333 0.571 0.574 0.632
NLI 0.216 0.376 0.549 0.368 0.623 0.578 0.661
CrossEncoder 0.242 0.462 0.613 0.345 0.624 0.575 0.662
$\text{CoCoA}_{MTE}$
AlignScore 0.277 0.302 0.522 0.382 0.657 0.556 0.598
RougeL 0.175 0.472 0.573 0.364 0.642 0.556 0.617
NLI 0.196 0.359 0.530 0.400 0.671 0.562 0.639
CrossEncoder 0.237 0.454 0.600 0.373 0.666 0.557 0.650
Table 7: Comparison of PRRs of CoCoA-family methods with different choices of similarity function. Main model response obtained by MBR decoding.
C.2 Different Ways of Combining Confidence and Consistency
We justify the particular form of equation (11) by considering alternative ways to combine sample-focused confidence with consistency estimation. Results are presented in Tables 89 and 10. In particular, we investigate the performance of the additive approach (AdditiveCoCoA):
$U_{\text{AdditiveCoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})+\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x}),$ (16)
and the same multiplicative combination, replacing sample-focused dissimilarity from (8) with the average of the full pairwise dissimilarity matrix (24):
$U_{\text{FullSampleCoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})\cdot U_{\text{DegMat}}(\mathbf{x}).$ (17)
In addition, we present results where we compute $u(\mathbf{y}_{*}\mid\mathbf{x})$ using probability-based scores instead of log-likelihood, see equation (6). We refer to this variant as $U_{\text{ProbCoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})$. While the results are generally similar, the log-likelihood-based formulation offers greater numerical stability. Therefore, we adopt it as the default in the main paper.
It is evident that on average the multiplicative form proposed in equation (10) with both confidence and consistency terms focused on a single sample is the better performing variant.
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
Mistral7b-Base
$\text{AdditiveCoCoA}_{SP}$ 0.290 0.319 0.459 0.351 0.654 0.471 0.472
$\text{FullSampleCoCoA}_{SP}$ 0.319 0.385 0.590 0.357 0.668 0.467 0.505
$\text{ProbCoCoA}_{SP}$ 0.059 0.302 0.520 0.390 0.671 0.461 0.435
$\text{CoCoA}_{SP}$ 0.330 0.396 0.598 0.383 0.670 0.466 0.517
$\text{AdditiveCoCoA}_{PPL}$ 0.262 0.392 0.564 0.369 0.671 0.464 0.494
$\text{FullSampleCoCoA}_{PPL}$ 0.277 0.373 0.551 0.334 0.672 0.467 0.435
$\text{ProbCoCoA}_{PPL}$ 0.297 0.369 0.566 0.373 0.674 0.464 0.475
$\text{CoCoA}_{PPL}$ 0.286 0.375 0.568 0.369 0.674 0.466 0.467
$\text{AdditiveCoCoA}_{MTE}$ -0.279 -0.058 -0.072 0.098 0.312 0.079 0.187
$\text{FullSampleCoCoA}_{MTE}$ 0.274 0.368 0.543 0.309 0.668 0.442 0.456
$\text{CoCoA}_{MTE}$ 0.288 0.374 0.564 0.355 0.673 0.447 0.491
Llama8b-Base
$\text{AdditiveCoCoA}_{SP}$ 0.330 0.345 0.462 0.301 0.566 0.502 0.326
$\text{FullSampleCoCoA}_{SP}$ 0.358 0.434 0.564 0.333 0.589 0.488 0.354
$\text{ProbCoCoA}_{SP}$ 0.031 0.405 0.471 0.371 0.612 0.461 0.368
$\text{CoCoA}_{SP}$ 0.378 0.456 0.582 0.349 0.597 0.485 0.372
$\text{AdditiveCoCoA}_{PPL}$ 0.368 0.431 0.504 0.336 0.595 0.455 0.437
$\text{FullSampleCoCoA}_{PPL}$ 0.389 0.420 0.487 0.314 0.580 0.450 0.399
$\text{ProbCoCoA}_{PPL}$ 0.381 0.445 0.513 0.345 0.599 0.446 0.438
$\text{CoCoA}_{PPL}$ 0.387 0.448 0.514 0.338 0.593 0.452 0.433
$\text{AdditiveCoCoA}_{MTE}$ -0.331 -0.042 -0.122 0.089 0.321 -0.122 0.117
$\text{FullSampleCoCoA}_{MTE}$ 0.383 0.410 0.481 0.308 0.588 0.363 0.414
$\text{CoCoA}_{MTE}$ 0.380 0.446 0.511 0.337 0.601 0.402 0.447
Falcon7b-Base
$\text{AdditiveCoCoA}_{SP}$ 0.203 0.318 0.409 0.350 0.674 0.533 0.379
$\text{FullSampleCoCoA}_{SP}$ 0.225 0.423 0.571 0.388 0.678 0.533 0.404
$\text{ProbCoCoA}_{SP}$ 0.226 0.367 0.515 0.416 0.680 0.526 0.426
$\text{CoCoA}_{SP}$ 0.257 0.433 0.578 0.396 0.684 0.529 0.436
$\text{AdditiveCoCoA}_{PPL}$ 0.222 0.433 0.580 0.413 0.677 0.525 0.489
$\text{FullSampleCoCoA}_{PPL}$ 0.204 0.425 0.565 0.393 0.669 0.533 0.437
$\text{ProbCoCoA}_{PPL}$ 0.235 0.433 0.576 0.410 0.680 0.528 0.482
$\text{CoCoA}_{PPL}$ 0.229 0.436 0.580 0.406 0.677 0.529 0.478
$\text{AdditiveCoCoA}_{MTE}$ 0.001 -0.103 -0.106 0.114 0.041 0.138 0.221
$\text{FullSampleCoCoA}_{MTE}$ 0.201 0.425 0.557 0.377 0.675 0.519 0.470
$\text{CoCoA}_{MTE}$ 0.228 0.439 0.577 0.395 0.685 0.517 0.510
Table 8: Comparison of PRRs of CoCoA-family methods with alternative formulations. Main model response obtained via greedy decoding.
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
Mistral7b-Base
$\text{AdditiveCoCoA}_{SP}$ 0.333 0.239 0.310 0.406 0.631 0.472 0.311
$\text{FullSampleCoCoA}_{SP}$ 0.354 0.543 0.565 0.412 0.643 0.468 0.428
$\text{ProbCoCoA}_{SP}$ 0.076 0.684 0.721 0.428 0.643 0.464 0.846
$\text{CoCoA}_{SP}$ 0.366 0.712 0.730 0.430 0.644 0.466 0.562
$\text{AdditiveCoCoA}_{PPL}$ 0.368 0.737 0.751 0.406 0.644 0.465 0.939
$\text{FullSampleCoCoA}_{PPL}$ 0.383 0.714 0.723 0.379 0.649 0.468 0.933
$\text{ProbCoCoA}_{PPL}$ 0.369 0.738 0.756 0.401 0.649 0.467 0.935
$\text{CoCoA}_{PPL}$ 0.372 0.735 0.755 0.402 0.648 0.466 0.937
$\text{AdditiveCoCoA}_{MTE}$ 0.368 0.723 0.702 0.332 0.643 0.452 0.942
$\text{FullSampleCoCoA}_{MTE}$ 0.380 0.661 0.653 0.331 0.643 0.442 0.929
$\text{CoCoA}_{MTE}$ 0.373 0.708 0.732 0.373 0.645 0.447 0.935
Llama8b-Base
$\text{AdditiveCoCoA}_{SP}$ 0.466 0.349 0.425 0.333 0.555 0.473 0.285
$\text{FullSampleCoCoA}_{SP}$ 0.476 0.462 0.619 0.363 0.574 0.464 0.379
$\text{ProbCoCoA}_{SP}$ 0.035 0.491 0.617 0.398 0.598 0.433 0.795
$\text{CoCoA}_{SP}$ 0.484 0.529 0.685 0.384 0.587 0.452 0.513
$\text{AdditiveCoCoA}_{PPL}$ 0.454 0.536 0.673 0.358 0.575 0.425 0.923
$\text{FullSampleCoCoA}_{PPL}$ 0.459 0.525 0.649 0.343 0.556 0.430 0.914
$\text{ProbCoCoA}_{PPL}$ 0.438 0.547 0.689 0.364 0.574 0.419 0.923
$\text{CoCoA}_{PPL}$ 0.450 0.544 0.689 0.364 0.573 0.422 0.925
$\text{AdditiveCoCoA}_{MTE}$ 0.457 0.496 0.579 0.304 0.561 0.361 0.901
$\text{FullSampleCoCoA}_{MTE}$ 0.455 0.464 0.577 0.313 0.563 0.341 0.878
$\text{CoCoA}_{MTE}$ 0.451 0.520 0.638 0.346 0.582 0.363 0.900
Falcon7b-Base
$\text{AdditiveCoCoA}_{SP}$ 0.100 0.397 0.394 0.393 0.649 0.534 -0.156
$\text{FullSampleCoCoA}_{SP}$ 0.144 0.531 0.607 0.416 0.654 0.533 0.189
$\text{ProbCoCoA}_{SP}$ 0.282 0.522 0.670 0.434 0.658 0.529 0.978
$\text{CoCoA}_{SP}$ 0.210 0.564 0.698 0.428 0.659 0.530 0.498
$\text{AdditiveCoCoA}_{PPL}$ 0.297 0.582 0.706 0.417 0.643 0.526 1.000
$\text{FullSampleCoCoA}_{PPL}$ 0.297 0.560 0.670 0.405 0.641 0.533 1.000
$\text{ProbCoCoA}_{PPL}$ 0.311 0.587 0.718 0.414 0.648 0.531 1.000
$\text{CoCoA}_{PPL}$ 0.310 0.579 0.717 0.415 0.644 0.530 1.000
$\text{AdditiveCoCoA}_{MTE}$ 0.253 0.554 0.634 0.383 0.630 0.523 0.997
$\text{FullSampleCoCoA}_{MTE}$ 0.237 0.502 0.554 0.383 0.636 0.519 0.989
$\text{CoCoA}_{MTE}$ 0.289 0.551 0.678 0.402 0.646 0.517 0.998
Table 9: Comparison of PRRs of CoCoA-family methods with alternative formulations. Main model response obtained by selecting the most probable sample.
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
Mistral7b-Base
$\text{AdditiveCoCoA}_{SP}$ 0.259 0.382 0.488 0.231 0.388 0.512 0.711
$\text{FullSampleCoCoA}_{SP}$ 0.287 0.447 0.591 0.318 0.570 0.516 0.748
$\text{ProbCoCoA}_{SP}$ 0.258 0.374 0.473 0.149 0.099 -0.265 0.710
$\text{CoCoA}_{MTE}$ 0.295 0.441 0.589 0.336 0.597 0.514 0.766
$\text{AdditiveCoCoA}_{PPL}$ 0.234 0.430 0.560 0.255 0.517 0.507 0.721
$\text{FullSampleCoCoA}_{PPL}$ 0.251 0.458 0.571 0.305 0.587 0.516 0.745
$\text{ProbCoCoA}_{PPL}$ 0.094 -0.123 -0.143 -0.202 -0.404 -0.423 0.027
$\text{CoCoA}_{PPL}$ 0.258 0.451 0.577 0.322 0.612 0.514 0.768
$\text{AdditiveCoCoA}_{MTE}$ 0.201 0.468 0.585 0.296 0.669 0.501 0.760
$\text{FullSampleCoCoA}_{MTE}$ 0.232 0.426 0.544 0.318 0.673 0.496 0.772
$\text{CoCoA}_{MTE}$ 0.236 0.419 0.553 0.337 0.680 0.494 0.783
Llama8b-Base
$\text{AdditiveCoCoA}_{SP}$ 0.197 0.345 0.465 0.177 0.287 0.571 0.676
$\text{FullSampleCoCoA}_{SP}$ 0.214 0.409 0.589 0.289 0.449 0.589 0.724
$\text{ProbCoCoA}_{SP}$ 0.196 0.337 0.449 0.102 0.053 0.434 0.673
$\text{CoCoA}_{MTE}$ 0.225 0.403 0.590 0.304 0.487 0.586 0.731
$\text{AdditiveCoCoA}_{PPL}$ 0.182 0.406 0.500 0.229 0.401 0.540 0.715
$\text{FullSampleCoCoA}_{PPL}$ 0.186 0.433 0.563 0.304 0.461 0.565 0.746
$\text{ProbCoCoA}_{PPL}$ 0.101 -0.097 -0.152 -0.175 -0.367 0.099 0.053
$\text{CoCoA}_{PPL}$ 0.198 0.428 0.580 0.320 0.496 0.562 0.751
$\text{AdditiveCoCoA}_{MTE}$ 0.160 0.437 0.554 0.302 0.560 0.477 0.748
$\text{FullSampleCoCoA}_{MTE}$ 0.174 0.400 0.543 0.334 0.568 0.475 0.762
$\text{CoCoA}_{MTE}$ 0.189 0.398 0.561 0.345 0.587 0.472 0.765
Falcon7b-Base
$\text{AdditiveCoCoA}_{SP}$ 0.188 0.356 0.401 0.228 0.488 0.578 0.368
$\text{FullSampleCoCoA}_{SP}$ 0.237 0.468 0.586 0.325 0.593 0.576 0.453
$\text{ProbCoCoA}_{SP}$ 0.184 0.339 0.378 0.144 0.350 -0.345 0.364
$\text{CoCoA}_{MTE}$ 0.248 0.461 0.590 0.336 0.617 0.575 0.472
$\text{AdditiveCoCoA}_{PPL}$ 0.206 0.404 0.565 0.271 0.537 0.572 0.625
$\text{FullSampleCoCoA}_{PPL}$ 0.231 0.464 0.606 0.337 0.598 0.577 0.650
$\text{ProbCoCoA}_{PPL}$ 0.014 -0.167 -0.265 -0.187 -0.237 -0.416 -0.206
$\text{CoCoA}_{PPL}$ 0.242 0.462 0.613 0.345 0.624 0.575 0.662
$\text{AdditiveCoCoA}_{MTE}$ 0.191 0.488 0.612 0.353 0.644 0.564 0.617
$\text{FullSampleCoCoA}_{MTE}$ 0.219 0.452 0.591 0.368 0.654 0.556 0.629
$\text{CoCoA}_{MTE}$ 0.237 0.454 0.600 0.373 0.666 0.557 0.650
Table 10: Comparison of PRRs of CoCoA-family methods with alternative formulations. Main model response obtained by MBR decoding.
Appendix D Detailed Description of Uncertainty Quantification Methods
In this section, we provide a detailed description of the uncertainty quantification methods used in this study.
D.1 Information-Based Methods
Information-based methods are commonly used to estimate uncertainty by analyzing the probability distributions of tokens within a given output. These methods examine different levels of model generation, such as the model’s confidence in producing a specific sequence, its ability to predict individual tokens at each generation step, and the variability in the token-level predictions across the sequence.
Sequence Probability (SP) is one of the simplest and most direct methods for estimating uncertainty. It measures the probability of the most likely output sequence given a specific input. Thus, uncertainty is quantified by calculating the probability of the sequence with the highest likelihood, under the assumption that the model is most confident in this output. In its simplest form it is given by (6). An equivalent (in terms of ordering of predictions for different inputs) but more numerically stable formulation is logarithmic. It is defined as:
$U_{\text{SP}}(\mathbf{y}\mid\mathbf{x})=-\log p(\mathbf{y}\mid\mathbf{x}).$ (18)
Perplexity (PPL) is another widely used metric for estimating uncertainty in language models [fomicheva-etal-2020-unsupervised]. It measures the model’s confidence by evaluating the average likelihood of generating the sequence tokens:
$U_{\mathrm{PPL}}(\mathbf{y}\mid\mathbf{x})=-\frac{1}{L}\log p(\mathbf{y}\mid\mathbf{x}).$ (19)
Mean Token Entropy takes a broader view of uncertainty by considering the token-level predictions across the entire sequence [fomicheva-etal-2020-unsupervised]. Instead of evaluating the model’s confidence in a single output or individual token predictions, Mean Token Entropy calculates the average entropy of the token probability distributions for each token in the sequence:
$U_{\mathcal{H}_{T}}(\mathbf{y}\mid\mathbf{x})=\frac{1}{L}\sum_{l=1}^{L}\mathcal{H}(y_{l}\mid\mathbf{y}_{<l},\mathbf{x}),$ (20)
where $\mathcal{H}(y_{l}\mid\mathbf{y}_{<l},\mathbf{x})$ is an entropy of the token distribution $p(y_{l}\mid\mathbf{y}_{<l},\mathbf{x})$.
The TokenSAR method, introduced in [duan-etal-2024-shifting], generalizes length-normalized log probability by computing a weighted average of the negative log probabilities of generated tokens, where weights are based on token relevance to the overall text. Using a similarity function $s(\cdot,\cdot)$ and token relevance function $R_{T}(y_{k},\mathbf{y},\mathbf{x})=1-s(\mathbf{x}\cup\mathbf{y},\mathbf{x}\cup\mathbf{y}\setminus y_{k})$, the uncertainty estimate is calculated as:
$U_{\mathrm{TokenSAR}}(\mathbf{y}\mid\mathbf{x})=-\sum_{l=1}^{L}\tilde{\mathrm{R}}_{T}(y_{l},\mathbf{y},\mathbf{x})\log p(y_{l}\mid\mathbf{y}_{<l},\mathbf{x}),$ (21)
where
$\tilde{\mathrm{R}}_{T}(y_{k},\mathbf{y}\mid\mathbf{x})=\frac{\mathrm{R}_{T}(y_{k},\mathbf{y},\mathbf{x})}{\sum\nolimits_{l=1}^{L}\mathrm{R}_{T}(y_{l},\mathbf{y},\mathbf{x})}.$ (22)
This measure is central for computing SAR uncertainty measure.
D.2 Consistency-Based Methods
Consistency-based methods assess the uncertainty of a language model by evaluating the semantic consistency of its predictions across multiple outputs for the same prompt. The core idea is that semantically similar outputs indicate higher confidence, while diverse or conflicting outputs suggest greater uncertainty. Since language models can express the same meaning in different surface forms, these methods construct a semantic similarity matrix $S=(s_{ij})$, where each entry represents the degree of similarity between pairs of responses. By clustering responses into groups with equivalent meanings, these methods provide a semantic measure of the model’s consistency.
The work [lin2023generating] offers two similarity measures to evaluate the similarity of sequences. The first is the Jaccard similarity, which treats sequences as sets of words and calculates the proportion of shared words to the total number of unique words in both sequences: $s(\mathbf{y},\mathbf{y}^{\prime})=|\mathbf{y}\cap\mathbf{y}^{\prime}|/|\mathbf{y}\cup\mathbf{y}^{\prime}|$.
Natural Language Inference (NLI) provides another method for computing similarity between sequences. We use the DeBERTa-large NLI model [he2021deberta], following [kuhn2023semantic]. For each pair of sequences, an NLI model predicts two probabilities: ${p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$, indicating entailment, and ${p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$, indicating contradiction. Similarity is then defined as either $s_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})={p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$ or $s_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})=1-{p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$.
Among the simplest consistency-based approaches are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian [lin2023generating]. Number of Semantic Sets estimates how many distinct “meanings” the model produces by clustering its outputs with an NLI model. The number of semantic sets is initially equal to the total number of generated answers, $M$. Two sentences are grouped into the same cluster if the following conditions are satisfied: ${p}_{\text{entail}}\bigl(\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr)>{p}_{\text{contra}}\bigl(\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr)\quad\text{and}\quad{p}_{\text{entail}}\bigl(\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr)>{p}_{\text{contra}}\bigl(\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr)$. This computation is performed for all pairs of answers, and the final number of distinct clusters is denoted by $U_{\text{NumSemSets}}$.
Sum of Eigenvalues of the Graph Laplacian examines global diversity: it constructs a similarity matrix among the sampled outputs and computes a continuous uncertainty score from the eigenvalues of the Laplacian of that similarity graph. The work [lin2023generating] proposes computing an averaged similarity matrix as $s_{ij}=\bigl(s\bigl(\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr)+s\bigl(\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr)\bigr)/2$. The Laplacian for the matrix $S$ is defined as $L=I-D^{-\frac{1}{2}}SD^{-\frac{1}{2}}$, where $D$ is a (diagonal) degree matrix with elements $D_{ii}=\sum_{j=1}^{M}s_{ij}$. Consequently, the following formula is derived:
$U_{\text{EigV}}(\mathbf{x})=\sum_{i=1}^{M}\max(0,1-\lambda_{i}(\mathbf{x})).$ (23)
Both Number of Semantic Sets and Sum of Eigenvalues of the Graph Laplacian effectively capture overall variation in generated text but cannot produce an individual uncertainty score for each output. To address this, the work [lin2023generating] proposes to use the diagonal Degree Matrix $D(\mathbf{x})$ which represents the total similarity of each answer with all others. The corrected trace of $D(\mathbf{x})$ provides an average pairwise distance between answers, and uncertainty is computed as:
$U_{\mathrm{DegMat}}(\mathbf{x})=1-\mathrm{trace}(D(\mathbf{x}))/M^{2}.$ (24)
D.3 Information-Based Methods with Repeated Sampling
The natural idea is to somehow benefit from having multiple samples from the model while using important information contained in the output probabilities estimated by an LLM. Below, we examine several approaches that have sought to achieve this.
Averaging uncertainties.
We can compute the entropy on the sequence level $\mathbb{E}\bigl[-\log p(\mathbf{y}\mid\mathbf{x})\bigr]$, where the expectation is taken over the sequences $\mathbf{y}$ randomly generated from the distribution $p(\mathbf{y}\mid\mathbf{x})$. Unfortunately, while for token level, we have an exact way of computing the entropy, for the sequence level, we need to adhere to some approximations. In practice, we can use Monte-Carlo integration, i.e. generate several sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$ via random sampling and compute Monte Carlo Sequence Entropy [kuhn2023semantic]:
$U_{\mathcal{H}_{S}}(\mathbf{x})=-\frac{1}{M}\sum_{i=1}^{M}\log p(\mathbf{y}^{(i)}\mid\mathbf{x}).$ (25)
We can replace $p(\mathbf{y}^{(i)}\mid\mathbf{x})$ with its length-normalized version $\bar{p}(\mathbf{y}^{(i)}\mid\mathbf{x})$ leading to a more reliable uncertainty measure in some cases.
While simple averaging represents a natural way to aggregate uncertainties, it has certain issues related to the nature of LLMs. First of all, in the vast majority of applications, an LLM-based system should produce a single output $\mathbf{y}_{*}$ for an input query. When we consider $U_{\mathcal{H}_{S}}(\mathbf{x})$ or other similar measure, we essentially perform averaging of uncertainties of different sequences, thus somewhat assessing the uncertainty related to the entire generative distribution $p(\mathbf{y}\mid\mathbf{x})$ for the input $\mathbf{x}$, but not for a particular generated sequence $\mathbf{y}_{*}$. This averaged uncertainty might not be adequate for this particular sequence and, remarkably, often performs worse than the uncertainty $u_{*}=U(\mathbf{y}_{*}\mid\mathbf{x})$, which is related solely to the output $\mathbf{y}_{*}$. Moreover, although intuitive, this naïve aggregation method assumes that all outputs contribute equally to the final uncertainty estimate, regardless of their semantic relationships. This can lead to inconsistencies when semantically equivalent outputs have varying uncertainty scores or when outputs with low similarity are treated as equally important.
Semantically weighted averaging. Semantic Entropy [kuhn2023semantic] addresses the issue of generated sequences with similar meanings but differing probabilities according to the model, which can heavily influence the resulting entropy value (25). The method clusters generated sequences $\mathbf{y}^{(i)},\,i=1,\dots,M$ into semantically homogeneous groups $\mathcal{C}_{k},~k=1,\dots,K$ (where $K\leq M$) using a bi-directional entailment algorithm. Probabilities of sequences are averaged within each cluster. The entropy estimate is then defined as:
$U_{\mathrm{SE}}(\mathbf{x})=-\sum_{k=1}^{K}\frac{|\mathcal{C}_{k}|}{M}\log\hat{p}_{k}(\mathbf{x}),$ (26)
where $\hat{p}_{k}(\mathbf{x})=\sum_{\mathbf{y}\in\mathcal{C}_{k}}p(\mathbf{y}\mid\mathbf{x})$ represents the aggregated probability for cluster $\mathcal{C}_{k}$.
SentenceSAR [duan-etal-2024-shifting] enhances the probability of sentences that are more relevant. It uses a sentence relevance measure $s\bigl(\mathbf{y}^{(j)},\mathbf{y}^{(k)}\bigr)$ to evaluate the relevance of $\mathbf{y}^{(j)}$ with respect to $\mathbf{y}^{(k)}$. SentenceSAR is calculated as:
$U_{\mathrm{SentSAR}}(\mathbf{x})=-\frac{1}{M}\sum_{i=1}^{M}\log\Bigl(p(\mathbf{y}^{(i)}\mid\mathbf{x})+\frac{1}{t}\mathrm{R}_{S}(\mathbf{y}^{(i)},\mathbf{x})\Bigr),$ (27)
where $t$ is a temperature parameter used to control the scale of shifting to relevance, and
$\mathrm{R}_{S}(\mathbf{y}^{(j)},\mathbf{x})\!=\sum_{k\neq j}s\bigl(\mathbf{y}^{(j)},\mathbf{y}^{(k)}\bigr)p\bigl(\mathbf{y}^{(k)}\mid\mathbf{x}\bigr).$ (28)
The combination of SentenceSAR and TokenSAR results in a unified method called SAR [duan-etal-2024-shifting]. In this approach, the generative probability $p(\mathbf{y}\mid\mathbf{x})$ in the SentenceSAR formula is replaced with the token-shifted probability $p^{\prime}(\mathbf{y}\mid\mathbf{x})=\exp\bigl\{-\mathrm{TokenSAR}(\mathbf{y},\mathbf{x})\bigr\}$, creating a comprehensive measure that integrates both sentence- and token-level adjustments.
The aggregation approaches like Semantic Entropy [kuhn2023semantic] or SAR [duan-etal-2024-shifting] can be unified into a semantically-aware Generalized Monte Carlo uncertainty estimate, defined as
$U_{\text{GMCU}}(\mathbf{x})=\frac{1}{M}\sum_{i=1}^{M}h\Biggl(\sum_{j=1}^{M}s\bigl(\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr)p\bigl(\mathbf{y}^{(j)}\mid\mathbf{x}\bigr)\Biggr).$ (29)
Here, the inner summation aggregates sequence probabilities $p\bigl(\mathbf{y}^{(j)}\mid\mathbf{x}\bigr)$ weighted by their semantic similarity to the $i$-th output, and the outer summation averages these contributions across all samples. The function $h(\cdot)$ provides an additional layer of flexibility, transforming the reweighted uncertainty scores, making the method a generalized framework for uncertainty quantification.
Unfortunately, methods that fall under GMCU, while offering benefits, also inherit the aforementioned issues from both categories of methods. In particular, the outer summation in (29), similarly to the case of simple Monte Carlo averaging, often fails to outperform the uncertainty $U_{*}(\mathbf{x})=U(\mathbf{y}_{*}\mid\mathbf{x})$ of a single generated sequence $\mathbf{y}_{*}$.
D.4 Verbalized Uncertainty
Verbalized uncertainty methods refer to approaches that prompt a model to explicitly express its confidence. In our experiment, we utilize the P(True) method, implemented following the description by [kadavath2022language]. Specifically, the model is presented with the original question and its answer, and then prompted to indicate whether the answer is True or False. We use the negative log-probability of the token “True” as the uncertainty score.
Appendix E Detailed Experimental Results
Tables 3 and 1 presented PRR scores averaged over datasets corresponding to each task. Here we present expanded results for each dataset, see Tables 11, 12 and 13.
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
Mistral7b-Base
MCSE 0.007 0.257 0.350 0.247 0.496 0.337 0.475
MCNSE 0.009 0.342 0.478 0.238 0.540 0.356 0.401
Semantic Entropy 0.008 0.271 0.382 0.271 0.562 0.387 0.472
DegMat 0.137 0.229 0.382 0.336 0.646 0.410 0.299
EigValLaplacian 0.132 0.207 0.328 0.301 0.624 0.398 0.241
SAR 0.094 0.353 0.517 0.313 0.644 0.419 0.471
P(True) 0.179 0.118 0.033 -0.061 -0.128 0.068 0.005
Consistency Light -0.006 0.317 0.457 0.306 0.597 0.432 0.441
Consistency 0.051 0.285 0.500 0.379 0.647 0.423 0.435
SP 0.287 0.315 0.451 0.326 0.628 0.474 0.471
$\text{CoCoA}_{SP}$ 0.330 0.396 0.598 0.383 0.670 0.466 0.517
$\text{CoCoA}_{SP}$ Light 0.309 0.380 0.580 0.351 0.674 0.480 0.506
PPL 0.204 0.365 0.489 0.281 0.632 0.474 0.311
$\text{CoCoA}_{PPL}$ 0.286 0.375 0.568 0.369 0.674 0.466 0.467
$\text{CoCoA}_{PPL}$ Light 0.260 0.404 0.557 0.326 0.684 0.480 0.428
MTE 0.182 0.392 0.484 0.243 0.619 0.456 0.350
$\text{CoCoA}_{MTE}$ 0.288 0.374 0.564 0.355 0.673 0.447 0.491
$\text{CoCoA}_{MTE}$ Light 0.254 0.415 0.550 0.303 0.672 0.465 0.458
Llama8b-Base
MCSE 0.033 0.293 0.354 0.237 0.482 0.171 0.351
MCNSE 0.022 0.370 0.415 0.219 0.501 0.170 0.344
Semantic Entropy 0.033 0.297 0.389 0.272 0.549 0.229 0.375
DegMat 0.081 0.250 0.355 0.353 0.622 0.342 0.309
EigValLaplacian 0.079 0.198 0.278 0.332 0.604 0.292 0.273
SAR 0.077 0.427 0.483 0.311 0.595 0.352 0.398
P(True) 0.058 0.047 0.037 -0.037 -0.066 -0.180 0.026
Consistency Light -0.022 0.475 0.441 0.353 0.552 0.279 0.378
Consistency 0.024 0.389 0.453 0.375 0.614 0.391 0.368
SP 0.328 0.342 0.456 0.277 0.526 0.508 0.324
$\text{CoCoA}_{SP}$ 0.378 0.456 0.582 0.349 0.597 0.485 0.372
$\text{CoCoA}_{SP}$ Light 0.358 0.448 0.556 0.340 0.598 0.504 0.353
PPL 0.369 0.351 0.422 0.253 0.507 0.461 0.303
$\text{CoCoA}_{PPL}$ 0.387 0.448 0.514 0.338 0.593 0.452 0.433
$\text{CoCoA}_{PPL}$ Light 0.382 0.483 0.491 0.319 0.597 0.467 0.398
MTE 0.357 0.357 0.408 0.239 0.497 0.349 0.326
$\text{CoCoA}_{MTE}$ 0.380 0.446 0.511 0.337 0.601 0.401 0.447
$\text{CoCoA}_{MTE}$ Light 0.372 0.501 0.487 0.316 0.599 0.387 0.412
Falcon7b-Base
MCSE 0.159 0.297 0.337 0.258 0.549 0.420 0.427
MCNSE 0.108 0.371 0.474 0.293 0.586 0.442 0.299
Semantic Entropy 0.164 0.307 0.389 0.294 0.581 0.463 0.418
DegMat 0.201 0.274 0.431 0.407 0.651 0.480 0.395
EigValLaplacian 0.201 0.229 0.394 0.381 0.645 0.454 0.358
SAR 0.144 0.398 0.517 0.381 0.649 0.508 0.387
P(True) -0.159 0.175 0.135 0.036 0.275 0.027 0.133
Consistency Light 0.232 0.483 0.468 0.327 0.657 0.362 0.363
Consistency 0.226 0.337 0.496 0.408 0.656 0.485 0.426
SP 0.201 0.312 0.400 0.321 0.662 0.539 0.377
$\text{CoCoA}_{SP}$ 0.257 0.433 0.578 0.396 0.684 0.529 0.436
$\text{CoCoA}_{SP}$ Light 0.242 0.463 0.566 0.349 0.691 0.522 0.405
PPL 0.155 0.375 0.525 0.316 0.644 0.539 0.326
$\text{CoCoA}_{PPL}$ 0.229 0.436 0.580 0.406 0.677 0.529 0.478
$\text{CoCoA}_{PPL}$ Light 0.234 0.497 0.558 0.364 0.694 0.522 0.468
MTE 0.152 0.409 0.537 0.291 0.633 0.533 0.367
$\text{CoCoA}_{MTE}$ 0.228 0.439 0.577 0.395 0.685 0.517 0.510
$\text{CoCoA}_{MTE}$ Light 0.234 0.518 0.549 0.345 0.693 0.491 0.500
Table 11: Detailed experimental results. Main model response obtained by greedy decoding.
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
Mistral7b-Base
MCSE 0.162 0.406 0.407 0.289 0.492 0.339 0.693
MCNSE 0.196 0.471 0.507 0.277 0.529 0.358 0.700
Semantic Entropy 0.164 0.434 0.442 0.312 0.554 0.389 0.675
DegMat 0.205 0.439 0.410 0.376 0.618 0.410 0.454
EigValLaplacian 0.197 0.388 0.344 0.342 0.600 0.399 0.361
SAR 0.175 0.563 0.590 0.347 0.620 0.421 0.780
P(True) 0.207 0.472 0.269 -0.058 -0.084 0.068 0.278
Consistency 0.071 0.670 0.708 0.405 0.614 0.423 0.846
SP 0.330 0.212 0.291 0.388 0.607 0.476 0.307
$\text{CoCoA}_{SP}$ 0.366 0.712 0.730 0.430 0.644 0.466 0.562
PPL 0.365 0.695 0.676 0.327 0.615 0.476 0.931
$\text{CoCoA}_{PPL}$ 0.372 0.735 0.755 0.402 0.648 0.466 0.937
MTE 0.350 0.668 0.606 0.254 0.594 0.457 0.932
$\text{CoCoA}_{MTE}$ 0.373 0.708 0.732 0.373 0.645 0.447 0.935
Llama8b-Base
MCSE 0.192 0.366 0.395 0.258 0.465 0.158 0.545
MCNSE 0.186 0.377 0.480 0.239 0.484 0.165 0.631
Semantic Entropy 0.194 0.371 0.451 0.286 0.528 0.213 0.557
DegMat 0.191 0.274 0.409 0.366 0.606 0.320 0.396
EigValLaplacian 0.190 0.216 0.333 0.339 0.587 0.274 0.351
SAR 0.159 0.441 0.571 0.327 0.578 0.340 0.667
P(True) 0.058 0.075 0.056 -0.011 -0.071 -0.120 -0.084
Consistency 0.030 0.473 0.598 0.394 0.600 0.353 0.793
SP 0.464 0.339 0.413 0.304 0.514 0.483 0.280
$\text{CoCoA}_{SP}$ 0.484 0.529 0.685 0.384 0.587 0.452 0.513
PPL 0.458 0.504 0.622 0.293 0.483 0.441 0.911
$\text{CoCoA}_{PPL}$ 0.450 0.544 0.689 0.364 0.573 0.422 0.924
MTE 0.449 0.437 0.501 0.238 0.458 0.326 0.883
$\text{CoCoA}_{MTE}$ 0.451 0.520 0.638 0.345 0.582 0.363 0.899
Falcon7b-Base
MCSE 0.128 0.399 0.419 0.285 0.535 0.421 0.598
MCNSE 0.153 0.395 0.452 0.318 0.588 0.443 0.771
Semantic Entropy 0.134 0.420 0.460 0.319 0.566 0.463 0.567
DegMat 0.177 0.350 0.422 0.422 0.637 0.480 0.633
EigValLaplacian 0.174 0.289 0.382 0.393 0.622 0.454 0.522
SAR 0.193 0.455 0.521 0.385 0.642 0.509 0.826
P(True) 0.022 0.245 0.245 0.038 0.244 0.028 0.815
Consistency 0.282 0.491 0.651 0.416 0.627 0.484 0.979
SP 0.099 0.385 0.378 0.369 0.638 0.540 -0.175
$\text{CoCoA}_{SP}$ 0.210 0.564 0.698 0.428 0.659 0.530 0.498
PPL 0.275 0.541 0.637 0.353 0.614 0.540 1.000
$\text{CoCoA}_{PPL}$ 0.310 0.579 0.717 0.415 0.644 0.530 1.000
MTE 0.186 0.475 0.510 0.317 0.573 0.534 0.984
$\text{CoCoA}_{MTE}$ 0.289 0.551 0.678 0.402 0.646 0.517 0.998
Table 12: Detailed experimental results. Main model response obtained by selecting most probable candidate among stochastically sampled responses.
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
Mistral7b-Base
MCSE 0.123 0.297 0.363 0.245 0.515 0.385 0.751
MCNSE 0.105 0.385 0.485 0.256 0.552 0.406 0.548
Semantic Entropy 0.127 0.310 0.395 0.272 0.580 0.439 0.735
DegMat 0.244 0.280 0.365 0.358 0.673 0.461 0.392
EigValLaplacian 0.243 0.238 0.319 0.334 0.654 0.443 0.318
SAR 0.204 0.392 0.503 0.334 0.658 0.478 0.656
P(True) 0.079 0.142 0.084 -0.056 -0.084 0.047 -0.084
Consistency 0.211 0.297 0.437 0.355 0.659 0.458 0.603
SP 0.259 0.381 0.486 0.216 0.346 0.514 0.711
$\text{CoCoA}_{MTE}$ 0.295 0.441 0.589 0.336 0.597 0.514 0.766
PPL 0.205 0.391 0.513 0.186 0.358 0.514 0.672
$\text{CoCoA}_{PPL}$ 0.258 0.451 0.577 0.322 0.612 0.514 0.768
MTE 0.156 0.443 0.550 0.234 0.598 0.508 0.717
$\text{CoCoA}_{MTE}$ 0.236 0.419 0.553 0.337 0.680 0.494 0.783
Llama8b-Base
MCSE 0.089 0.240 0.354 0.220 0.444 0.297 0.647
MCNSE 0.055 0.334 0.443 0.247 0.470 0.306 0.545
Semantic Entropy 0.089 0.250 0.382 0.245 0.514 0.361 0.650
DegMat 0.217 0.218 0.361 0.368 0.611 0.391 0.539
EigValLaplacian 0.217 0.167 0.280 0.347 0.592 0.305 0.492
SAR 0.125 0.393 0.520 0.345 0.578 0.472 0.642
P(True) -0.014 0.098 0.091 -0.014 -0.054 -0.210 0.083
Consistency 0.146 0.290 0.440 0.387 0.603 0.428 0.599
SP 0.197 0.344 0.463 0.161 0.247 0.577 0.675
$\text{CoCoA}_{MTE}$ 0.225 0.403 0.590 0.304 0.487 0.586 0.731
PPL 0.168 0.382 0.460 0.161 0.253 0.549 0.668
$\text{CoCoA}_{PPL}$ 0.198 0.428 0.580 0.320 0.496 0.562 0.751
MTE 0.130 0.412 0.509 0.245 0.466 0.459 0.699
$\text{CoCoA}_{MTE}$ 0.189 0.398 0.561 0.345 0.587 0.472 0.765
Falcon7b-Base
MCSE 0.188 0.331 0.362 0.274 0.560 0.455 0.460
MCNSE 0.126 0.387 0.494 0.306 0.605 0.479 0.396
Semantic Entropy 0.192 0.332 0.414 0.304 0.594 0.497 0.438
DegMat 0.246 0.271 0.470 0.392 0.672 0.516 0.415
EigValLaplacian 0.238 0.220 0.432 0.365 0.660 0.476 0.339
SAR 0.184 0.412 0.543 0.370 0.664 0.553 0.536
P(True) -0.124 0.194 0.161 0.016 0.279 0.061 0.284
Consistency 0.235 0.323 0.498 0.376 0.665 0.523 0.588
SP 0.187 0.355 0.398 0.213 0.458 0.583 0.367
$\text{CoCoA}_{SP}$ 0.248 0.461 0.590 0.336 0.617 0.575 0.472
PPL 0.177 0.371 0.504 0.201 0.439 0.583 0.556
$\text{CoCoA}_{PPL}$ 0.242 0.462 0.613 0.345 0.624 0.575 0.662
MTE 0.159 0.462 0.580 0.290 0.588 0.565 0.517
$\text{CoCoA}_{MTE}$ 0.237 0.454 0.600 0.373 0.666 0.557 0.650
Table 13: Detailed experimental results. Main model response obtained by MBR decoding.
Appendix F Alternative Performance Metrics for PRR
The choice of PRR as a UQ quality metric of choice is dictated by its ability to handle both continuous performance metrics, like Comet and AlignScore without the need for selecting arbitrary thresholds, as well as relative robustness to class imbalance in case of binary performance metric. However, PRR scores are calculated for a particular choice of underlying performance metric. For a truly comprehensive evaluation, we perform the same evaluation as in our main experimental run, but with alternative choice of performance metrics.
Tables 14 and 15 report PRRs for these metrics on all models for NMT and QA tasks. MetricX [juraska-etal-2024-metricx] was used for NMT and GPT-as-a-judge (gpt-4o-2024-08-06) was used to score QA datasets. The following is the prompt used to facilitate GPT QA scoring:
You are a text evaluator. The model was asked the following question:
{question}
The ’Generated’ text is a model’s response. The ’Target’ is the correct answer.
If the generated answer correctly answers the question based on the target, return 1.
If it is wrong, return 0.
Respond ONLY with a single digit: 1 or 0.
Generated: {model output}
Target: {target sequence}
NMT (MetricX) QA (GPT-4o)
Method WMT14FrEn WMT19DeEn CoQA/Gpt TriviaQA/Gpt MMLU/Gpt GSM8k/Gpt
Mistral7b-Base
MCSE 0.127 0.212 0.226 0.478 0.338 0.453
MCNSE 0.267 0.417 0.246 0.535 0.357 0.388
Semantic Entropy 0.158 0.261 0.254 0.554 0.387 0.454
CEDegMat 0.270 0.402 0.336 0.659 0.402 0.356
SAR 0.302 0.451 0.333 0.660 0.419 0.456
Semantic Density 0.291 0.366 0.056 0.688 0.335 0.195
SP 0.173 0.287 0.287 0.640 0.473 0.485
$\text{CoCoA}_{SP}$ 0.314 0.462 0.367 0.681 0.465 0.537
PPL 0.316 0.448 0.220 0.645 0.473 0.345
$\text{CoCoA}_{PPL}$ 0.361 0.512 0.331 0.687 0.465 0.499
MTE 0.350 0.455 0.190 0.631 0.456 0.387
$\text{CoCoA}_{MTE}$ 0.366 0.512 0.323 0.684 0.447 0.524
Llama8b-Base
MCSE 0.145 0.181 0.200 0.486 0.158 0.332
MCNSE 0.319 0.361 0.183 0.530 0.170 0.324
Semantic Entropy 0.157 0.235 0.239 0.552 0.203 0.361
CEDegMat 0.287 0.367 0.350 0.621 0.299 0.300
SAR 0.354 0.420 0.302 0.610 0.319 0.382
Semantic Density 0.252 0.328 0.069 0.650 0.335 0.299
SP 0.176 0.274 0.228 0.564 0.469 0.313
$\text{CoCoA}_{SP}$ 0.329 0.451 0.332 0.625 0.449 0.365
PPL 0.327 0.388 0.217 0.549 0.462 0.309
$\text{CoCoA}_{PPL}$ 0.395 0.469 0.329 0.623 0.443 0.445
MTE 0.353 0.388 0.214 0.543 0.358 0.329
$\text{CoCoA}_{MTE}$ 0.394 0.472 0.333 0.634 0.393 0.455
Falcon7b-Base
MCSE 0.168 0.218 0.228 0.587 0.420 0.443
MCNSE 0.304 0.399 0.265 0.640 0.442 0.323
Semantic Entropy 0.201 0.276 0.262 0.623 0.463 0.448
CEDegMat 0.287 0.428 0.380 0.712 0.476 0.347
SAR 0.339 0.457 0.367 0.715 0.508 0.410
Semantic Density 0.323 0.422 0.076 0.746 0.396 0.349
SP 0.176 0.275 0.279 0.737 0.539 0.391
$\text{CoCoA}_{SP}$ 0.335 0.483 0.364 0.763 0.529 0.459
PPL 0.322 0.454 0.245 0.722 0.539 0.341
$\text{CoCoA}_{PPL}$ 0.394 0.531 0.345 0.754 0.529 0.513
MTE 0.370 0.455 0.228 0.707 0.533 0.385
$\text{CoCoA}_{MTE}$ 0.403 0.529 0.341 0.762 0.516 0.551
Table 14: PRRs for all models on QA and NMT tasks with alternative choice of performance metrics. Main model response obtained by greedy decoding.
NMT (MetricX) QA (GPT-4o)
Method WMT14FrEn WMT19DeEn CoQA/Gpt TriviaQA/Gpt MMLU/Gpt GSM8k/Gpt
Mistral7b-Base
MCSE 0.305 0.324 0.244 0.464 0.339 0.392
MCNSE 0.360 0.466 0.252 0.516 0.358 0.188
Semantic Entropy 0.336 0.366 0.272 0.536 0.389 0.420
CEDegMat 0.468 0.485 0.339 0.623 0.404 0.239
SAR 0.455 0.527 0.336 0.628 0.421 0.298
Semantic Density 0.565 0.554 0.043 0.646 0.306 0.421
SP 0.169 0.252 0.311 0.608 0.474 0.460
$\text{CoCoA}_{SP}$ 0.607 0.623 0.376 0.647 0.466 0.550
PPL 0.510 0.552 0.265 0.619 0.474 0.096
$\text{CoCoA}_{PPL}$ 0.620 0.667 0.346 0.655 0.466 0.091
MTE 0.487 0.499 0.211 0.604 0.457 0.167
$\text{CoCoA}_{MTE}$ 0.581 0.638 0.324 0.653 0.447 0.151
Llama8b-Base
MCSE 0.284 0.271 0.191 0.451 0.155 0.302
MCNSE 0.354 0.419 0.197 0.493 0.171 0.251
Semantic Entropy 0.299 0.333 0.230 0.515 0.201 0.341
CEDegMat 0.343 0.451 0.369 0.598 0.301 0.241
SAR 0.412 0.497 0.315 0.584 0.323 0.311
Semantic Density 0.403 0.452 0.065 0.638 0.286 0.396
SP 0.247 0.289 0.225 0.523 0.467 0.332
$\text{CoCoA}_{SP}$ 0.473 0.568 0.346 0.597 0.440 0.418
PPL 0.437 0.517 0.238 0.498 0.459 0.185
$\text{CoCoA}_{PPL}$ 0.510 0.614 0.335 0.588 0.433 0.199
MTE 0.387 0.439 0.201 0.485 0.345 0.226
$\text{CoCoA}_{MTE}$ 0.495 0.589 0.330 0.604 0.372 0.237
Falcon7b-Base
MCSE 0.269 0.341 0.261 0.553 0.421 0.159
MCNSE 0.325 0.403 0.312 0.620 0.443 0.064
Semantic Entropy 0.308 0.397 0.296 0.588 0.463 0.172
CEDegMat 0.381 0.478 0.398 0.686 0.477 0.083
SAR 0.389 0.483 0.398 0.685 0.509 0.152
Semantic Density 0.434 0.518 0.047 0.661 0.377 0.299
SP 0.269 0.311 0.340 0.683 0.540 0.104
$\text{CoCoA}_{SP}$ 0.465 0.621 0.407 0.705 0.530 0.300
PPL 0.430 0.553 0.308 0.663 0.540 0.156
$\text{CoCoA}_{PPL}$ 0.526 0.655 0.376 0.691 0.530 0.128
MTE 0.418 0.452 0.275 0.615 0.535 0.054
$\text{CoCoA}_{MTE}$ 0.512 0.620 0.370 0.692 0.517 0.050
Table 15: PRRs for all models on QA and NMT tasks with alternative choice of performance metrics. Main model response obtained by selecting the most probable candidate among stochastically sampled responses.
Appendix G AUROC
We strongly believe that due to the considerations presented in Appendix F, PRR is a superior metric to AUROC when comparing relative performance of UQ methods. However, we acknowledge that AUROC is widely used in modern UQ literature, so we opt to include AUROC results on QA datasets. The results are presented in Tables 16 and 17. All results here were obtained with GPT-as-a-judge correctness scoring.
CoQA/Gpt TriviaQA/Gpt MMLU/Gpt GSM8k/Gpt
Mistral7b-Base
MCSE 0.642 0.784 0.737 0.716
MCNSE 0.653 0.806 0.744 0.684
Semantic Entropy 0.653 0.817 0.756 0.716
CEDegMat 0.690 0.859 0.759 0.682
SAR 0.694 0.862 0.770 0.717
Semantic Density 0.531 0.871 0.713 0.576
SP 0.674 0.849 0.794 0.684
$\text{CoCoA}_{SP}$ 0.708 0.867 0.790 0.729
PPL 0.636 0.855 0.794 0.628
$\text{CoCoA}_{PPL}$ 0.690 0.871 0.790 0.714
MTE 0.620 0.849 0.787 0.640
$\text{CoCoA}_{MTE}$ 0.687 0.870 0.783 0.727
Llama8b-Base
MCSE 0.635 0.779 0.620 0.705
MCNSE 0.630 0.799 0.629 0.689
Semantic Entropy 0.655 0.815 0.652 0.713
CEDegMat 0.708 0.845 0.703 0.691
SAR 0.694 0.842 0.715 0.724
Semantic Density 0.540 0.855 0.680 0.655
SP 0.647 0.816 0.787 0.669
$\text{CoCoA}_{SP}$ 0.700 0.847 0.776 0.718
PPL 0.640 0.815 0.778 0.661
$\text{CoCoA}_{PPL}$ 0.699 0.848 0.770 0.745
MTE 0.633 0.817 0.726 0.667
$\text{CoCoA}_{MTE}$ 0.702 0.854 0.746 0.753
Falcon7b-Base
MCSE 0.648 0.807 0.770 0.766
MCNSE 0.660 0.831 0.779 0.709
Semantic Entropy 0.669 0.830 0.788 0.761
CEDegMat 0.718 0.870 0.784 0.722
SAR 0.713 0.870 0.805 0.757
Semantic Density 0.544 0.881 0.746 0.716
SP 0.681 0.866 0.820 0.731
$\text{CoCoA}_{SP}$ 0.719 0.884 0.815 0.774
PPL 0.645 0.860 0.820 0.700
$\text{CoCoA}_{PPL}$ 0.705 0.882 0.815 0.800
MTE 0.630 0.854 0.818 0.721
$\text{CoCoA}_{MTE}$ 0.704 0.886 0.810 0.814
Table 16: AUROC for all models on QA tasks. Main model response was obtained by greedy decoding.
CoQA/Gpt TriviaQA/Gpt MMLU/Gpt GSM8k/Gpt
Mistral7b-Base
MCSE 0.650 0.777 0.738 0.676
MCNSE 0.656 0.798 0.745 0.602
Semantic Entropy 0.661 0.808 0.757 0.692
CEDegMat 0.693 0.845 0.760 0.625
SAR 0.695 0.849 0.771 0.643
Semantic Density 0.523 0.854 0.696 0.633
SP 0.684 0.836 0.795 0.663
$\text{CoCoA}_{SP}$ 0.713 0.853 0.790 0.720
PPL 0.657 0.843 0.795 0.592
$\text{CoCoA}_{PPL}$ 0.697 0.858 0.790 0.615
MTE 0.627 0.838 0.788 0.606
$\text{CoCoA}_{MTE}$ 0.689 0.858 0.783 0.632
Llama8b-Base
MCSE 0.632 0.763 0.617 0.653
MCNSE 0.633 0.783 0.629 0.629
Semantic Entropy 0.653 0.800 0.650 0.669
CEDegMat 0.716 0.836 0.703 0.633
SAR 0.698 0.831 0.715 0.659
Semantic Density 0.534 0.849 0.655 0.666
SP 0.647 0.801 0.787 0.627
$\text{CoCoA}_{SP}$ 0.706 0.836 0.771 0.686
PPL 0.648 0.796 0.776 0.635
$\text{CoCoA}_{PPL}$ 0.702 0.835 0.765 0.658
MTE 0.624 0.794 0.719 0.644
$\text{CoCoA}_{MTE}$ 0.701 0.842 0.736 0.673
Falcon7b-Base
MCSE 0.662 0.797 0.770 0.589
MCNSE 0.682 0.826 0.779 0.571
Semantic Entropy 0.684 0.821 0.788 0.593
CEDegMat 0.726 0.863 0.785 0.580
SAR 0.727 0.862 0.806 0.611
Semantic Density 0.525 0.856 0.734 0.684
SP 0.706 0.850 0.820 0.555
$\text{CoCoA}_{SP}$ 0.738 0.869 0.815 0.667
PPL 0.677 0.844 0.820 0.657
$\text{CoCoA}_{PPL}$ 0.719 0.866 0.815 0.657
MTE 0.654 0.825 0.818 0.603
$\text{CoCoA}_{MTE}$ 0.717 0.867 0.810 0.628
Table 17: AUROC for all models on QA tasks. Main model response obtained by selecting the most probable candidate among stochastically sampled responses
Appendix H Computational Budget
Total available computational resources used to produce results in this paper amounted to 12 compute nodes each having 4xA100 40Gb GPUs. Total computational budget spent to produce results was around 400 GPU-days. Each individual combination of model and dataset amounted roughly to 16 GPU-days on average to obtain all model outputs and hidden states needed for computing the results.
Appendix I CoCoA Light Training Details
The model architecture consisted of a simple multilayer perceptron (MLP) with a single hidden layer:
  • Input dimension: equal to the embedding size of the corresponding base model (4096 for Mistral7b-Base and Llama8b-Base, 3072 for Falcon7b-Base).
  • Hidden dimension: 2048.
  • Output dimension: 4096.
  • Dropout: 0.1 applied between layers.
We trained the MLP on mean pooled hidden-layer embeddings extracted from the middle layer of an LLM. Table 18 details the size of a train set for each model and dataset.
Dataset LLaMA Falcon Mistral
CoQA 10,000 10,000 10,000
GSM8K 3,000 3,000 2,500
MMLU 1,461 1,461 1,461
TriviaQA 10,000 10,000 10,000
WMT14 Fr-En 6,000 6,000 6,000
WMT19 De-En 6,000 6,000 6,000
XSum 7,500 5,000 6,500
Table 18: Training set sizes (number of examples) for Llama8b-Base, Falcon7b-Base, and Mistral7b-Base across several datasets.
Training was conducted for 20 epochs with the following hyperparameters:
  • Batch size: 4 (per device) with gradient accumulation of 7 steps (effective batch size of 28).
  • Learning rate: $1\times 10^{-5}$ with AdamW optimizer.
  • Weight decay: 0.1.
  • Warmup ratio: 0.05.
  • Gradient clipping: 1.0.
Appendix J Larger Model Experiments
We additionally evaluate our method using the Gemma 3 12B-Base model to assess its performance on a larger-scale architecture. As shown in Table 19, CoCoA continues to achieve the best average performance among all methods.
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
MCSE -0.014 0.277 0.408 0.262 0.556 0.449 0.386
MCNSE -0.031 0.393 0.442 0.280 0.595 0.468 0.344
Semantic Entropy -0.011 0.273 0.430 0.287 0.621 0.531 0.407
DegMat 0.130 0.246 0.370 0.356 0.692 0.561 0.359
EigValLaplacian 0.130 0.195 0.307 0.323 0.667 0.526 0.312
SAR 0.087 0.429 0.494 0.335 0.688 0.582 0.418
P(True) -0.022 0.012 -0.020 -0.006 0.195 0.046 0.145
Consistency Light -0.006 0.328 0.483 0.345 0.709 0.408 0.421
Consistency -0.163 0.452 0.496 0.311 0.657 0.293 0.387
MSP 0.257 0.315 0.499 0.319 0.672 0.630 0.335
$\text{CoCoA}_{MSP}$ 0.308 0.424 0.641 0.372 0.723 0.617 0.391
$\text{CoCoA}_{MSP}$ Light 0.287 0.416 0.634 0.356 0.720 0.613 0.360
PPL 0.303 0.369 0.451 0.288 0.681 0.630 0.275
$\text{CoCoA}_{PPL}$ 0.339 0.427 0.539 0.362 0.728 0.617 0.439
$\text{CoCoA}_{PPL}$ Light 0.327 0.489 0.518 0.334 0.727 0.613 0.353
MTE 0.284 0.372 0.413 0.269 0.664 0.617 0.310
$\text{CoCoA}_{MTE}$ 0.326 0.422 0.527 0.344 0.728 0.592 0.468
$\text{CoCoA}_{MTE}$ Light 0.308 0.501 0.498 0.313 0.719 0.565 0.386
Table 19: Detailed experimental results for Gemma 3 12B-Base model. Model response obtained by greedy decoding
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
MCSE 0.051 0.346 0.447 0.273 0.531 0.447 0.495
MCNSE 0.088 0.441 0.456 0.319 0.580 0.465 0.533
Semantic Entropy 0.054 0.348 0.492 0.294 0.597 0.527 0.505
DegMat 0.165 0.294 0.372 0.366 0.659 0.558 0.457
EigValLaplacian 0.164 0.228 0.306 0.327 0.634 0.522 0.415
SAR 0.109 0.474 0.526 0.353 0.661 0.577 0.593
P(True) -0.029 0.037 -0.015 0.019 0.193 0.046 0.278
Consistency Light 0.076 0.478 0.631 0.373 0.671 0.443 0.744
MSP 0.188 0.353 0.466 0.361 0.646 0.624 0.223
$\text{CoCoA}_{MSP}$ 0.251 0.542 0.718 0.406 0.693 0.613 0.433
PPL 0.178 0.501 0.619 0.349 0.660 0.624 0.750
$\text{CoCoA}_{PPL}$ 0.211 0.576 0.692 0.400 0.700 0.613 0.805
MTE 0.168 0.431 0.487 0.295 0.643 0.612 0.703
$\text{CoCoA}_{MTE}$ 0.212 0.546 0.641 0.369 0.702 0.591 0.796
Table 20: Detailed experimental results for Gemma 3 12B-Base model. Model response obtained by most probable sample decoding
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
MCSE 0.120 0.290 0.361 0.271 0.544 0.487 0.493
MCNSE 0.126 0.366 0.439 0.308 0.597 0.497 0.464
Semantic Entropy 0.121 0.267 0.396 0.307 0.611 0.550 0.468
DegMat 0.195 0.222 0.318 0.399 0.684 0.571 0.373
EigValLaplacian 0.197 0.170 0.254 0.367 0.653 0.522 0.319
SAR 0.204 0.409 0.464 0.372 0.680 0.591 0.508
P(True) 0.030 0.043 0.018 0.014 0.204 0.079 0.313
Consistency Light 0.192 0.311 0.419 0.368 0.690 0.421 0.467
MSP 0.222 0.390 0.491 0.240 0.330 0.542 0.441
$\text{CoCoA}_{MSP}$ 0.282 0.452 0.598 0.347 0.624 0.574 0.498
PPL 0.187 0.400 0.526 0.183 0.344 0.542 0.468
$\text{CoCoA}_{PPL}$ 0.270 0.476 0.587 0.337 0.638 0.574 0.557
MTE 0.139 0.425 0.515 0.251 0.628 0.623 0.498
$\text{CoCoA}_{MTE}$ 0.230 0.451 0.528 0.361 0.712 0.610 0.575
Table 21: Detailed experimental results for Gemma 3 12B-Base model. Model response obtained by MBR decoding